### v1

# Existence for Stable Rotating Star-Planet Systems

Abstract
This paper investigates the existence and properties of stable, uniformly rotating star-planet systems, i.e. mass ratio is sufficiently small. It is modeled by the Euler-Poisson equations. Following the framework established by McCann for binary stars [46], we adopt a variational approach, and prove the existence of local energy minimizers with respect to the Wasserstein $L^{\infty}$ metric, under the assumed equation of state $P(\rho)=K\rho^{\gamma}$ and under the condition that the mass ratio $m$ is sufficiently small, corresponding to a star–planet system. Such minimizers correspond to solutions of the Euler–Poisson system. We consider two cases. For $\gamma>2$, we not only prove existence but also show, via scaling arguments, that the radii (to be precise, the bounds of the supports of the minimizers) tend to zero. For $\frac{3}{2}<\gamma\leq 2$, we estimate an upper bound for the (potential) expansion rates of the radii, and it turns out that the existence result remains valid in this case as well. Finally, we provide estimates for the distances between different connected components of supports of minimizers and propose a conjecture regarding the number of connected components.
Key words: Gaseous Star-Planet Systems, Euler-Poisson Equations, Calculus of Variations, Wasserstein $L^{\infty}$ Topology, Scaling Method, Bootstrap Method
1 Introduction
In astrophysical fluid dynamics, gas stars and gas planets can be modeled as isolated self-gravitating fluid masses. A fundamental hydrodynamic description for Newtonian stars is provided by the Euler–Poisson system:
$\begin{split}\partial_{t}\rho+\nabla\cdot\left({\rho v}\right)&=0,\\ \rho\partial_{t}v+\rho\left({v\cdot\nabla}\right)v+\nabla P(\rho)&=\rho\nabla V,\\ \Delta V&=-4\pi\rho.\end{split}$ (EP)
Here, $\rho(x,t)\geq 0$ denotes the fluid density at position $x\in\mathbb{R}^{3}$ and time $t\geq 0$, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and $V(x,t)\in\mathbb{R}$ represents the gravitational potential. The pressure $P(\rho)$ depends solely on density. In this work, we adopt the polytropic equation of state $P(\rho)=K\rho^{\gamma}$.
The motion of stars and planets has long been a classical subject in celestial mechanics, astrophysics, and mathematics. Historical models such as heliocentrism, proposed by Copernicus, and later studies by Newton and others on stellar rotation, form the foundation of rotating star or planet models. Early developments by Maclaurin, Jacobi, Poincaré, and Liapounov focused on incompressible, homogeneous or nearly homogeneous bodies like the Earth. For gaseous bodies such as Jupiter, or systems like galaxies, density inhomogeneity and compressibility need to be accounted for. The first attempts to construct axisymmetric rotating solutions to the compressible Euler–Poisson equations were made by astrophysicists Milne [47] and Zeipel [57] in the early 1920s, followed by Chandrasekhar [11] and Lichtenstein [39] in 1933. Further historical accounts can be found in [12, 33].
Significant progress has been made for the single rotating star problem, where the star rotates about its axis. Two main approaches have been employed to prove the existence of such rotating stars. The first is variational method, as used by Auchmuty and Beals [3, 4] to establish the existence of axisymmetric equilibria for compressible fluids; see also [2, 4, 9, 13, 38, 16]. The second is a perturbative method based on the implicit function theorem, applied near non-rotating Lane–Emden stars [29, 30, 28, 53, 54]. Nonlinear dynamical stability of rotating stars was shown in [43, 44] via variational techniques, while stability theory for non-rotating stars can be found in [18, 32, 49].
These methods are quite general and apply not only to single-body problems but also to $N$-body systems. Variational methods were used by McCann [46] for binary stars and are applied in this paper to star–planet systems. Perturbative approaches appear in the work of Alonso-Orán, Kepka, and Velázquez [1] on incompressible Euler–Poisson systems with an external particle, as well as in Lichtenstein’s construction of rotating binary stars [39].
Each method has its advantages and limitations. For example, the variational approach typically requires $\gamma>\frac{4}{3}$ when $P(\rho)=K\rho^{\gamma}$ to prevent gravitational collapse (see [16, Remark 2.11 and Proposition 2.16] [46, Section 2], or [4, Section 1, 6 and 8]). In contrast, perturbative methods can handle the range $\frac{6}{5}<\gamma<2$ [30], but they rely on prior knowledge of solutions in special non-rotating cases.
When considering two stellar objects, such as binary stars or a star–planet system, we begin with the simplest scenario where individual rotations are neglected, and only orbital revolution is considered. We still refer to such configurations as rotating systems. The construction of rotating binary stars dates back to Lichtenstein [39]. As discussed by Jang and Seok [31], in the $N$-body uniformly rotating case, the Euler–Poisson equations reduce to a single equation:
$\omega^{2}\tilde{\rho}\left(x\right)P\left(x\right)+\nabla P\left(\tilde{\rho}\left(x\right)\right)-\tilde{\rho}\left(x\right)\left(\nabla V_{\tilde{\rho}}\left(x\right)\right)=0$ (EP’)
[…]
A notable feature of McCann’s binary stars is that each star admits a variational characterization as a Hamiltonian or energy minimizer under certain constraints, while also being interpretable as a perturbation of simpler objects—such as non-rotating Lane–Emden stars and point-mass relative equilibria (see subsection 5.3 and Section 6 of this paper or [31, Section 3.2]). This characterization is crucial for showing that the support of an energy-constrained minimizer lies in the interior of balls of certain radii. Consequently, the minimizer $\rho$ is in fact a local minimizer in the Wasserstein $L^{\infty}$ distance sense. This ensures that $\nabla P(\rho)$ is well-defined throughout $\mathbb{R}^{3}$, and that (EP’) holds in the whole space $\mathbb{R}^{3}$ (see Theorem 3.1, [46, Theorem 2.1], [15, Section 2]).
Beyond existence, many questions remain for two-body system. For instance, Jang and Seok [31] study the asymptotic profiles, uniqueness, and orbital stability of McCann’s uniformly rotating binary stars. In addition, conjectures concerning the maximal possible number of connected components of the support of solutions are discussed in Section 7 of this paper.
Researchers have also considered systems with more than two stellar objects, such as galaxies or general $N$-star systems. For rotating galaxies, single rotating galaxies were constructed in [52, 53] in the spirit of Lichtenstein and Heilig [28, 39]. There is an extensive literature on non-rotating galaxies: see [6, 7, 24, 50] and references therein. The orbital stability of stationary solutions has seen considerable progress over the past two decades [26, 21, 22, 23, 20, 25, 34, 35, 36, 37, 50, 56], with recent work on linearly oscillating galaxies in [27]. In [10], Campos, Del Pino, and Dolbeault constructed $N$-body rotating galaxies by connecting to relative equilibria in $N$-body dynamics with small uniform angular velocity, perturbing radial equilibria, and solving for the gravitational potential rather than the density, using a finite-dimensional reduction. Although natural, Jang and Seok remark in [31] that the framework in [10] is not ideal for stability analysis. They adapt Rein’s reduction method [49, 50] to extend the discussion of binary star solutions to establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [31]. Furthermore, they note that for $N\geq 3$, uniformly rotating $N$-body stellar configurations lack a variational structure analogous to the binary case and are not generally expected to be stable [31].
This paper presents the core results extracted from the author’s Master’s thesis [14]. Building upon the preparatory work in two companion papers [15, 16] originating from the same thesis, it investigates the existence and qualitative properties of stable, uniformly rotating star–planet systems governed by the Euler–Poisson equations. This work thereby extends McCann’s framework for binary stars [46] to the setting of small mass ratios, as established in Theorems 2.12 and 2.13.
The paper is structured as follows. Section 2 introduces the variational formulation and states the main results. Section 3 collects preliminary results about two-body systems and non-rotating single star cases from [46, 4, 40, 15, 16]. Section 4, 5 and 6 are the core of this paper, where we will prove existence of star-planet systems. Sections 7 studies distances between (possible) different connected components of planets or stars, leading to the conjecture that minimizer may consist of at most two connected components, corresponding to exactly one star and exactly one planet respectively.
2 Variational formulation and Statement of Main Results
2.1 Variational formulation
We introduce some notations and variational formulation here, which based on McCann’s construction of binary star solutions, see [46] and [15].
Definition 2.1 (Notations).
We give the following definitions:
  1. (i)
    k{k} dimensional open ball: $B_{R}(a):=\left\{x\in\mathbb{R}^{k}\mid|x-a\mid<R\right\}$ for given $a\in\mathbb{R}^{k},R\in\mathbb{R}$.
  2. (ii)
    The projection operator of $x$ to the $x_{1}x_{2}$ plane: $P_{12}(x)=P_{12}\left(x_{1},x_{2},x_{3}\right):=\left(x_{1},x_{2},0\right)$.
  3. (iii)
    A bilinear form $\langle\cdot,\cdot\rangle_{2}:\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}:\forall x,y\in\mathbb{R}^{3},\langle x,y\rangle_{2}=P_{12}(x)\cdot P_{12}(y)=x_{1}y_{1}+x_{2}y_{2}$.
  4. (iv)
    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=x_{1}^{2}+x_{2}^{2}$.
  5. (v)
    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
  6. (vi)
    Positive part $\lambda]_{+}:=\max\{\lambda,0\}$.
  7. (vii)
    $W^{\infty}$ metric: Wasserstein $L^{\infty}$ metric. (See Definition 2.8)
  8. (viii)
    Given $\delta>0$ the $\delta$-neighbourhood of $\Omega\subset\mathbb{R}^{3}$ to be the set Ω+B_δ(0):=⋃_y ∈Ω{x ∈R^3∣— x-y —¡δ}
  9. (ix)
    $\mu$: Lebesgue measure.
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we need to first consider the equation of state for pressure $P(\rho)$. In this paper, we take the polytropic law, i.e.
$P(\rho)=K\rho^{\gamma}$ (1)
We also define $A(s)$ as the following:
$A(s):=s\int_{0}^{s}P(\tau)\tau^{-2}d\tau=\frac{K}{\gamma-1}s^{\gamma}$ (2)
We first give a “potential admissible class” for $\rho$ and $v$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=1\right\}$ (3)
$V\left(\mathbb{R}^{3}\right)$ $:=\left\{v:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}\mid v\text{ is measurable.}\right\}$ (4)
Then given $\rho$ and $v$ in such sets, the energy $E(\rho,v)$ consists of three terms:
$E(\rho,v):=U(\rho)-\frac{G(\rho,\rho)}{2}+T(\rho,v)$ (5)
$U(\rho):=\int A(\rho(x))d^{3}x$ (6)
$G(\sigma,\rho):=\int V_{\sigma}\rho=\iint\frac{\rho(x)\sigma(y)}{|x-y|}d^{3}yd^{3}x$ (7)
$T(\rho,v):=\frac{1}{2}\int|v|^{2}\rho$ (8)
Here $A(\rho)$ is a convex function given in (2), and $U(\rho)$ is internal energy.
We can choose units so that the total mass of fluid is one and the gravitational constant $G=1$, then $V_{\rho}$ represents the gravitational potential of the mass density $\rho(x)$
$V_{\rho}(x):=\int\frac{\rho(y)}{|y-x|}d^{3}y$ (9)
Hence $G(\rho,\rho)$ is gravitational interaction energy.
We can also see $T(\rho,v)$ is kinetic energy.
Remark 2.2.
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [15, Proposition 2.10]. Hence $E(\rho,v)$ is well-defined.
We can choose a frame of reference in which the center of mass
$\bar{x}(\rho):=\frac{\int x\rho(x)d^{3}x}{\int\rho(x)d^{3}x}$ (10)
is at rest. We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$. The fluid angular momentum $\boldsymbol{J}(\rho,v)$ is given by:
$\boldsymbol{J}(\rho,v):=\int(x-\bar{x}(\rho))\times v\rho(x)d^{3}x$ (11)
We denote by $J_{z}$ the z-component of $\boldsymbol{J}$, that is, $J_{z}(\rho,v):=\hat{e}_{z}\cdot\boldsymbol{J}(\rho,v)$, where $\hat{e}_{z}=(0,0,1)^{T}$. For simplicity of notation, we will sometimes use $J$ to represent $J_{z}$ and call $J$ the angular momentum of the system when no confusion arises.
Since the z-component of the angular momentum is specified, the moment of inertia $I(\rho)$ of $\rho$ in the direction of $\hat{e}_{z}$ will be relevant. That is, we define $I(\rho)$ to be moment of inertia of $\rho$ in the direction of $\hat{e}_{z}$:
$I(\rho):=\int\rho r^{2}(x-\bar{x}(\rho))$ (12)
where $r$ is given in Definition 2.1.
Remark 2.3.
When $\rho$ has positive mass, we have $I(\rho)>0$. See [15, Remark 2.9].
Since the star and planet are separated, it is convenient to describe the relations between the total moment of inertia $I\left(\rho_{m}+\rho_{1-m}\right)$ and $I\left(\rho_{m}\right),I\left(\rho_{1-m}\right)$.
Lemma 2.4 (Expansion of Moment of Inertia [15, Lemma 2.15]).
Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int\rho=m_{1}<\infty,\int\sigma=m_{2}<\infty$. $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
  1. (1)
    If $m_{1}+m_{2}=0$, then $I(\rho+\sigma)=0$.
  2. (2)
    If $m_{1}+m_{2}>0$, then we have the moment of inertia of $\rho+\sigma$ satisfies
    $I(\rho+\sigma)=I(\rho)+I(\sigma)+\frac{m_{1}m_{2}}{m_{1}+m_{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$ (13)
Let support of $\rho$ be the smallest closed set carrying the full mass of $\rho$ (denoted by spt $\rho$). Intuitively we hope spt $\rho$ to be compact as this aligns with the case of a star-planet system and ensures a finite moment of inertia. Therefore, we also introduce a subset $R_{a}\left(\mathbb{R}^{3}\right)$ of $R\left(\mathbb{R}^{3}\right)$:
${R}_{a}\left(\mathbb{R}^{3}\right):=\left\{\rho\in{R}\left(\mathbb{R}^{3}\right)\mid\bar{x}(\rho)=a;\text{ spt }\rho\text{ is bounded.}\right\}$ (14)
Since the energy is translation-invariant, we may just consider looking for a energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, which means $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$.
Thanks to Theorem 3.1 below, it turns out the problem of minimizing $E(\rho,v)$ locally in $R_{0}(\mathbb{R}^{3})\times V(\mathbb{R}^{3})$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ locally in $R_{0}(\mathbb{R}^{3})$. Here $E_{J}(\rho)$ corresponds to uniform rotation with the angular momentum $\boldsymbol{J}=J\hat{e}_{z}=(0,0,J)^{T}$ specified a priori, and $E_{J}(\rho)$ is defined as the following:
$E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$ (15)
Here, $T_{J}$ is given by
$T_{J}(\rho):=\frac{J^{2}}{2I(\rho)}$ (16)
Remark 2.5.
Given $\rho\in R(\mathbb{R}^{3})$, if we further know $\rho\in L^{\infty}(\mathbb{R}^{3})$, then we know $U(\rho)$ is finite by interpolation inequality (Proposition A.2). This observation, together with Remark 2.2 and Remark 2.3, tells us $E_{J}(\rho)$ is well-defined and finite. For more general pressure assumptions, see [15, Lemma 5.2].
We intend to solve (EP’) by minimizing $E_{J}(\rho)$ and using results from calculus of variations, which involves the variational derivative of energy. We therefore now proceed to define it. We define the perturbation set $P_{\infty}(\rho)$, which depends on energy minimizer $\rho$, as the following:
$P_{\infty}(\rho):=\bigcup_{R<\infty}P_{R}(\rho)$ (17)
Here, $P_{R}(\rho)$ is defined as:
[…]
One can see $P_{\infty}(\rho)$ is a convex cone. And we have the following result:
Lemma 2.6 (Differentiability of Energy $E_{J}(\rho)$ [15, Lemma 5.3, Remark 5.4]).
Given $\rho\in mR(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, i.e. it is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$. Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int E_{J}^{\prime}(\rho)\sigma$  [^1] To remain consistent with the notation in [4, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion.. The function $E_{J}^{\prime}(\rho)$ on the right‑hand side is given by
$E_{J}^{\prime}(\rho)(x):=A^{\prime}(\rho(x))-V_{\rho}(x)-\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))$ (18)
A simple case is the non-rotating problem, i.e. $\boldsymbol{J}=0$, with energy
$E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}$ (19)
Consider the non-rotating minimizer $\sigma_{m}$ of $E_{0}(\rho)$ among configurations of mass $m\geq 0$, the corresponding minimum energy turns out to be finite due to Theorem 3.6. For the sake of convenience, we denote them by
$e_{0}(m):=E_{0}\left(\sigma_{m}\right)=\inf_{\rho\in{R}\left(\mathbb{R}^{3}\right)}E_{0}(m\rho)<\infty.$ (20)
We also denote $e_{0}(1)$ by $e_{0}$. Properties of non-rotating minimizers are introduced in [16, 46, 4, 40]. We will also discuss them in Theorem 3.6.
In general, the angular momentum $\mathbf{J}$ is not 0. But under the constraint $J_{z}(\rho,v)=J$, where $J_{z}$ denotes the $z$-component of the angular momentum $\mathbf{J}$, if we know $\widetilde{\rho}$ is a local (w.r.t the topology induced by Wasserstein $L^{\infty}$ distance) minimizer over ${R}_{0}\left(\mathbb{R}^{3}\right)$ of $E_{J}(\rho)$, then $\widetilde{\rho}$ is a solution to (EP’). Note the internal energy and gravitational potential energy are rotation‑invariant, and physically we expect celestial bodies to be rotating rather than stationary. Therefore, we can further define
$({\rho}(t,x),v(t,x)):=\left(\widetilde{\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$
where $\omega=\frac{J}{I\left(\rho\right)}$ and $R_{\theta}$ is a rotation map about $x_{3}$ axis given in Definition 2.1, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating star-planet system. Here “rotating” means the orbital revolution around each other, while the rotations of objects around their own axes are not considered. Moreover, $E(\rho,v)$ is local minimum, and $({\rho}(t,x),v(t,x))$ is solution to (EP). We will discuss these results in details in Theorem 3.1.
For prescribed angular momentum, we also know the energy $E_{J}(\rho)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$. However, as in Morgan [48], McCann [46, Section 3] demonstrates that this bound — although approached — will not be attained, see also [45]. Therefore, instead of searching for a global energy minimizer in star-planet system, we adopt McCann’s strategy [46]: first establishing the existence of a constrained minimizer of $E_{J}(\rho)$, and then proving that it is also a local minimizer under the appropriate topology induced by Wasserstein $L^{\infty}$ distance.
Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu^{2}}$, where $\mu=m(1-m)$ is their reduced mass. Note that the same symbol $\mu$ is also used as for Lebesgue measure, provided no ambiguity arises. For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ is minimized subject to the constraint
$W_{m}:=\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int\rho_{m}=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (21)
where $\Omega_{m}$ and $\Omega_{1-m}$ are subsets of $\mathbb{R}^{3}$, which are given in the following. Note $W_{m}$, $\Omega_{m}$ and $\Omega_{1-m}$ are actually related to both $m$ and $J$.
Fix two points $y_{m}$ and $y_{1-m}$ in $\mathbb{R}^{3}$ from the plane $z=0$, which are separated by
$\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{m^{2}(1-m)^{2}}$ (22)
i.e. $\eta=\left|y_{m}-y_{1-m}\right|$. The $\Omega_{m}$ and $\Omega_{1-m}$ are defined as closed balls in $\mathbb{R}^{3}$ centered at $y_{m}$ and $y_{1-m}$, whose size and separation scale with $\eta$ as the following:
$\Omega_{m}$ $=\left\{x\in\mathbb{R}^{3}\mid|x-y_{m}|\,\leq\frac{\eta}{4}\right\}$ (23)
$\Omega_{1-m}$ $=\left\{x\in\mathbb{R}^{3}\mid|x-y_{1-m}|\,\leq\frac{\eta}{4}\right\}$
The distance separating $\Omega_{m}$ and $\Omega_{1-m}$, and the diameter of their union is given by:
${dist}\left(\Omega_{m},\Omega_{1-m}\right)$ $=\frac{\eta}{2}$ (24)
${diam}\left(\Omega_{m},\Omega_{1-m}\right)$ $=\frac{3\eta}{2}$ (25)
Remark 2.7.
The reason we set the separation $\eta=\frac{J^{2}}{\mu^{2}}$ in the definitions above is inspired by the Kepler problem. Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass. Assume their separation is $d$, the gravitational energy plus kinetic energy is $\frac{\mu}{d}+\frac{J^{2}}{2\mu d^{2}}$, which reaches its minimum at separation $d=\eta$.
In the star-planet model, when $\eta$ is large, the distance between two objects is large, the gravitational interaction becomes weak. The system thus approximates the ideal case of two isolated, non-rotating, and non-interacting bodies. Therefore, it can be proven that the supports of the objects remain bounded, which sizes are hence negligible compared to the distance between them. It implies the distance between their centers of mass asymptotically approaches that of the point mass model. We will discuss these facts in more details in Section 5 and Section 6, assuming $m$ is sufficiently small, and then establish their connection to the existence of local energy minimizers under topology induced by $W^{\infty}$ (Wasserstein $L^{\infty}$) distance. For the case $J$ is sufficiently large, one can also check McCann’s paper [46, Section 6].
We note the choice of topology for ${R}\left(\mathbb{R}^{3}\right)$ is quite delicate: for $J>0$, it turns out local energy minimizer for $E_{J}(\rho)$ will not exist if the topology of $R(\mathbb{R}^{3})$ is inherited from a topological vector space ([46, Remark 3.7] and [15, Proposition 5.15]). Roughly speaking, the reason is that for any reasonable candidate local minimizer $\rho$, one can always split off a portion of it with very small mass and shift the portion arbitrarily far away, thereby decreasing the energy. However, the resulting configuration still lies in a small neighborhood of $\rho$ with respect to such a topology, which implies that $\rho$ cannot be a local minimizer. See also Remark 3.10.
Hence we need another topology, i.e. the topology induced by the $W^{\infty}$ distance.
Definition 2.8 (Wasserstein $L^{\infty}$ distance).
Let $(X,d)$ be a metric space and $(S,\Sigma,v)$ be a finite positive measure space, given $\rho,\kappa\in\mathcal{P}(X)$, the Wasserstein $L^{\infty}$ distance between $\rho$ and $\kappa$ is defined as
$W^{\infty}(\rho,\kappa):=\inf\left\{\begin{array}[]{l}\|d(f(x),g(x))\|_{L^{\infty},v}\mid f:S\rightarrow X\text{ Borel },\\ f_{\#}v=\rho\text{ and }g:S\rightarrow X\text{ Borel, }g_{\#}v=\kappa\end{array}\right\}$ (26)
Here $\|h\|_{L^{\infty},v}$ denotes the supremum of $|h|$ over $S$, discarding sets of $v$-measure zero. $M(S)$ and $M(X)$ are measure spaces. $f_{\#}:M(S)\rightarrow M(X)$ ($g_{\#}$ is similar) is the corresponding push forward operator defined by
$f_{\#}\tilde{\mu}(B):=\tilde{\mu}\left(f^{-1}(B)\right)\quad\text{for any Borel set }B\text{ in }X$
We call $f,g$ the transport maps. Thanks to Strassen’s Theorem, one can check $W^{\infty}$ is truly a metric, as explained in Givens and Shortt [19]. We gives two properties of $W^{\infty}$ distance. For other properties, one can see [46, Lemma 5.1] and [15, Lemma 4.7].
Lemma 2.9 (Simple Properties of the Wasserstein $L^{\infty}$ Metric [46, Lemma 5.1]).
Let $\rho,\kappa$ in $t{R}\left(\mathbb{R}^{3}\right)$, then
  1. (ii)
    $W^{\infty}(\rho,\kappa)$ does not exceed the diameter of $\operatorname{spt}(\rho-\kappa)$
  2. (ii)
    if $W^{\infty}(\rho,\kappa)<\delta$, each connected component of the $\delta$-neighbourhood of spt $\rho$ has the same mass for $\kappa$ as for $\rho$
Remark 2.10.
The advantage of the topology induced by Wasserstein $L^{\infty}$ metric is that a neighborhood of $\rho$ does not include such “portion-shifted-to-far-away” configurations mention above due to Lemma 2.9 (ii), while thanks to Lemma 2.9 (i), the neighborhood still contains all perturbations of the form $\rho+t\sigma$ for all $t$, where $\sigma$ is supported on a set which diameter is small enough, with $\int\sigma=0$. Hence this setting still allows one to legitimately consider variational derivatives (18) and consequently Euler-Lagrange equations (EL) (see also Remark 3.4.).
Remark 2.11.
According to Lemma 2.9, to prove $\rho(m)$ is a $W^{\infty}$ local minimizer (and hence solution to (EP’) due to Theorem 3.1), it suffices to show $\rho(m)$ is a constraint minimizer on $W_{m}$ such that its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. In fact, once we know $\rho(m)$’s support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$, one can also show $\rho(m)$ is a solution to (EP’), as well as most results in Theorem 3.1 (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) without specifying $\rho(m)$ is a $W^{\infty}$ local minimizer, see [15, Remark 3.10].
It motivates our choice to define $\Omega_{m}$ and $\Omega_{1-m}$ with their radii that increase as $\eta$ increases, and then show the size of stars or planets will not expand too much as $\eta$ increases. We will show those results for small $m$ later, inspired by McCann’s arguments for large $J$ [46, Section 6].
2.2 Statement of Main Results
We are ready to describe our main results:
Theorem 2.12 (Existence of Star-Planet Systems for $\gamma>2$).
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>2$, fix $J>0$, there is a $\delta>0$, such that for all $m\in(0,\delta)$, a constrained energy minimizer $\rho(m)=\rho_{m}+\rho_{1-m}$ on $W_{m}$ exists, which is actually also a Wasserstein $L^{\infty}$ local energy minimizer of $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$. Moreover, $\rho(m)$ satisfies the following properties:
  • $(i)$
    $\rho(m)\in R_{0}(\mathbb{R}^{3})$, and $(\rho(m),v)$ minimizes $E(\rho,v)$ locally on $R(\mathbb{R}^{3})\times V(\mathbb{R}^{3})$ subject to the constraint $J_{z}(\rho,v)\coloneq\hat{e}_{z}\cdot\boldsymbol{J}(\rho,v)=J$ or $\mathbf{J}(\rho,~v)=J{\hat{e}}_{z}$. Here $v(x):=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\rho(m))}$.
  • $(ii)$
    $\rho(m)$ is symmetric about the plane $z=0$ and a decreasing function of $|z|$.
  • $(iii)$
    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int{\frac{\rho(m)(y)}{\left|{y-x}\right|}d^{3}y}}$. Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
  • $(iv)$
    For the planet density, we have ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$. For the star density, $\left\|\rho_{1-m}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly for all small $m$ , that is $m$ is located in a small interval $(0,\delta)$.
  • $(v)$
    The support for the planet density spt $\rho_{m}$ is contained in a ball of a radius which goes to 0 when $m$ goes to 0. The support for the star density spt $\rho_{1-m}$ is contained in a ball of radius $R(J)$ for all small $m$.
Actually, the results can be true in a larger range of index $\gamma$ if we don’t require the support of the planet to shrink as the mass goes to 0.
Theorem 2.13 (Existence of Star-Planet Systems for $\gamma>\frac{3}{2}$).
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$, fix $J>0$, there is a $\delta>0$, such that for all $m\in(0,\delta)$, a constrained energy minimizer $\rho(m)=\rho_{m}+\rho_{1-m}$ on $W_{m}$ exists, which is actually also a Wasserstein $L^{\infty}$ local energy minimizer of $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$. Moreover, $\rho(m)$ satisfies the following properties:
  • $(i)$
    $\rho(m)\in R_{0}(\mathbb{R}^{3})$, and $(\rho(m),v)$ minimizes $E(\rho,v)$ locally on $R(\mathbb{R}^{3})\times V(\mathbb{R}^{3})$ subject to the constraint $J_{z}(\rho,v)\coloneq\hat{e}_{z}\cdot\boldsymbol{J}(\rho,v)=J$ or $\mathbf{J}(\rho,~v)=J{\hat{e}}_{z}$. Here $v(x):=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\rho(m))}$.
  • $(ii)$
    $\rho(m)$ is symmetric about the plane $z=0$ and a decreasing function of $|z|$.
  • $(iii)$
    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int{\frac{\rho(m)(y)}{\left|{y-x}\right|}d^{3}y}}$. Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
  • $(iv)$
    For the planet density, we have ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$. For the star density, $\left\|\rho_{1-m}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly for all small $m$ , that is $m$ is located in a small interval $(0,\delta)$.
  • $(v)$
    The support for the star density spt $\rho_{1-m}$ is contained in a ball of radius $R(J)$ for all small $m$.
We provide the proof of Theorem 2.12 and Theorem 2.13 from Section 4 to Section 6. The general outline is as follows: Given Theorem 3.1, the remaining task is to demonstrate the existence of a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer. To achieve this, we first prove the existence of constrained minimizers $\rho(m)$ of $E_{J}(\rho)$ on $W_{m}$ when the mass ratio is sufficiently small in Section 4. In Section 5, we prove convergence of scaling densities in $L^{p}$, uniform $L^{\infty}$ bound, and then derives quantitative bounds on the supports of minimizers, . In Section 6 we estimate the center of mass separation and show that $\rho(m)$ can be situated within the interior of the considered domains we consider, that is, $\text{dist }\left(\text{spt}~\rho(m),~\mathbb{R}^{3}\backslash\left(\Omega_{m}\cup\Omega_{1-m}\right)\right)>0$, and then we verify $\rho(m)$ satisfies results in Theorem 2.12 and Theorem 2.13.
3 Preliminary Results
Recall that in the calculus of variations, if an energy functional admits a minimizer, then this minimizer satisfies Euler–Lagrange equation. When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, Theorem A.1 or Theorem 3.1. In this section, we review some variational‑based results and their extensions. In subsection 3.1, we discuss McCann’s work on Wasserstein $L^{\infty}$ local minimizers and existence of binary star system results. In subsection 3.2, we review some results for single-star system.
3.1 Results for 2-body systems
In this subsection, we review McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution with slight modification. We also make some remarks on them.
The following are McCann’s results on $W^{\infty}$ local minimizers.
Theorem 3.1 (Properties of $W^{\infty}$-Local Energy Minimizers [46, Section 2][15, Section 2]).
Let $J>0$. If $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$, where $J_{z}$ denotes the $z$-component of the angular momentum $\mathbf{J}$, then:
  1. (i)
    the z-axis is a principal axis of inertia for $\rho$, with a moment of inertia $I(\rho)$ which is maximal and non-degenerate;
  2. (ii)
    $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ if and only if $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and $v(x)=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\rho(m))}$, which implies $E(\rho,v)=E_{J}(\rho)$. It describes uniformly rotating fluid, in the sense that after defining $({\widetilde{\rho}}(t,x),v(t,x))=\left({\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$. Notice that $E(\rho,v)=E(\widetilde{\rho},v)=E_{J}(\widetilde{\rho})=E_{J}(\rho)$.
  3. (iii)
    $\rho$ is continuous on $\mathbb{R}^{3}$;
  4. (iv)
    on each connected component $\Omega_{i}$ of $\{\rho>0\},\rho$ satisfies Euler-Lagrange equation:
    $A^{\prime}(\rho(x))=\left[\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+V_{\rho}(x)+\lambda_{i}\right]_{+}$ (EL)
    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component;
  5. (v)
    the equations (EL) continue to hold on a $\delta$-neighbourhood of the $\Omega_{i}$;
  6. (vi)
    $\rho\in C^{1}(\{\rho>0\})$;
  7. (vii)
    $\rho$ satisfies the reduced Euler-Poisson equations (EP’) on $\mathbb{R}^{3}$, where the pressure $P(\rho)$ is given in (1). Moreover, $({\widetilde{\rho}}(t,x),v(t,x))$ solves Euler-Poisson equations (EP) with $V(t,x)=V_{{\rho}}\left(R_{-\omega t}x\right)$.
  8. (viii)
    this solution is stable with respect to $L^{\infty}$-small perturbations of the Lagrangian fluid variables.
Remark 3.2.
In Theorem 3.1 we only mention the case where $P(\rho)$ satisfies polytropic law (1), but the results hold true for a more general form of $P(\rho)$, see for instance [15, Theorem 2.25].
Remark 3.3.
Notice in Theorem 3.1, since the local minimizer $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\bar{x}(\rho)=(0,0,0)^{T}$ in (EL). However, one can generalize the case where the center of mass is not at the origin and $\rho$ is constrained minimizer instead of $W^{\infty}$ local minimizer, hence we retain the term $\bar{x}(\rho)$ in (EL) in Theorem 3.1 (iv) and (EL) holds on $\Omega_{i}$ instead of $\delta$-neighborhood of $\Omega_{i}$. This is consistent with the form in McCann’s paper [46, Section 6].
Remark 3.4.
McCann initially introduced and proved Theorem 3.1 in [46]. We later complemented this work in [15] by showing why (EP) also holds on the boundary of support of minimizer. Additionally, in [15] we verified the local minimizer $\rho$ has finite internal energy $U(\rho)<\infty$, hence $E_{J}$ is indeed differentiable at $\rho$ due to Lemma 2.6. Further discussion—including the reason why the local minimum energy is attained under uniform rotation (Theorem 3.1 (ii)), the interpretation of the stability result (Theorem 3.1 (viii)), the generalization of the fixed angular momentum from $J_{z}$ to $\mathbf{J}$, and the equivalence between (EP’) and (EL)—can be found in [46, 15, 31].
Thanks to Theorem 3.1, together with the arguments of McCann [46], we know a constrained energy minimizer on $W_{m}$ exists and its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$ when $J$ is sufficiently large. Therefore, this constrained minimizer is a $W^{\infty}$ local minimizer. And we know the following theorem:
Theorem 3.5 (Existence of Binary Stars [46, Theorem 6.1, Corollary 6.2]).
[…]
What’s more, after another translation the center of mass of $\widetilde{\rho}$ can be 0 and is a local minimizer of $E_{J}(\rho)$. Let $v(x):=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\widetilde{\rho})}$, then the pair $(\widetilde{\rho},v)$ minimizes $E(\rho,v)$ locally on ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ (thus on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right))$ subject to the constraint $J_{z}(\rho,v)=J$ or $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$. $\widetilde{\rho}$ satisfies reduced Euler-Poisson equations (EP’). Moreover, the uniformly rotating fluid $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$, here $({\rho}(t,x),v(t,x))=\left(\widetilde{\rho}\left(R_{-\omega t}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
3.2 Results for Non-rotating Bodies
In this subsection, we introduce the existence and uniqueness of non-rotating solutions of Euler-Poisson equations, corresponding the single-star case. We also indicate some useful properties about the solutions.
Theorem 3.6 (Non-rotating Stars [4, 40, 46]).
For $E_{0}(\rho)$ from (19), $e_{0}(m)$ from (20) and $m\in[0,\infty)$:
  1. (i)
    $E_{0}(\rho)$ attains its minimum $e_{0}(m)$ among $\rho$ such that $\rho\in m{R}\left(\mathbb{R}^{3}\right)$, and the minimizer $\sigma_{m}$ is unique up to translation;
  2. (ii)
    $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave. In particular, for $m>0$, $e_{0}(m)<0$, $e_{0}^{\prime}\left(m^{+}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{+}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ and $e_{0}^{\prime}\left(m^{-}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{-}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ exists, and $e_{0}^{\prime}(m^{+})<e_{0}^{\prime}(m^{-})<0$;
  3. (iii)
    $\sigma_{m}$ is spherically symmetric and radially decreasing after translation;
  4. (iv)
    $\left\|\sigma_{m}\right\|_{L^{\infty}}\leq C(m)$;
  5. (v)
    spt $\sigma_{m}$ is contained in a ball of radius $R_{0}(m)$;
  6. (vi)
    $\sigma_{m}\in C^{0}(\mathbb{R}^{3})\cap C^{1}(\{\sigma_{m}>0\})$;
  7. (vii)
    $\sigma_{m}$ satisfies
    $A^{\prime}(\sigma_{m}(x))=\left[V_{\sigma_{m}}(x)+\lambda_{m}\right]_{+}$ (EL0)
    on all of $\mathbb{R}^{3}$ and a single Lagrange multiplier $\lambda_{m}=-(5\gamma-6)m^{\frac{2\gamma-2}{3\gamma-4}}U(\sigma)\left\{\begin{array}[]{l}<0,m>0\\ =0,m=0\end{array}\right$., where $\sigma=\sigma_{1}$ is the minimizer of $E_{0}(\rho)$ with mass $1$;
  8. (viii)
    the derivatives of $e_{0}(m)$ exists and satisfies: $e_{0}^{\prime}(m)=\lambda_{m}$;
  9. (ix)
    $\sigma_{m}$ satisfies the reduced Euler-Poisson equations (EP’).
Remark 3.7.
Theorem 3.6 (iii), which is based on the rearrangement inequality, establishes that the support of $\rho$ is simply connected. This configuration corresponds to a single-star system and aligns with our theoretical expectations. In contrast, although McCann’s binary star model [46] demonstrates that the support of $\rho$ is contained in two disjoint balls, it does not establish that the support consists of exactly two connected components. Whether the support actually comprises exactly two connected components—thus truly representing a binary star system—remains an open question. In Section 7, we estimate the distance between possibly multiple connected components and propose Conjecture 7.6, which asserts that there can be at most two components.
Remark 3.8.
Since $\sigma$ has mass 1 and satisfies Theorem 3.6 (iii), $\sigma\in L^{1}\cap L^{\infty}$, then $U(\sigma)$ is finite by interpolation inequality Proposition A.2. Hence the equation for $\lambda$ in Theorem 3.6 (vii) makes sense.
Remark 3.9.
In Theorem 3.6, we only mention the case where $P(\rho)$ satisfies polytropic law (1), but some results hold true for a more general form of $P(\rho)$, see for instance [4, 40, 46, 16]. Auchmuty and Beals initially concluded that (EP’) is satisfied in the region where $\rho>0$ [4, Theorem A]. We later improved the result in [16, Theorem 2.7] by showing the equation holds over the entire $\mathbb{R}^{3}$ space. Theorem 3.6 (ii) was stated without proof by McCann [46, Section 3], who mentioned it could be demonstrated using a method analogous to that in [40]—which is in the quantum mechanical framework. A proof for the polytropic case was later given in [16].
Note that without polytropic law assumption, the uniqueness of minimizer, the existence of $e^{\prime}_{0}(m)$ and the equality relations between $\lambda_{m}$, $U(\sigma$) and $e^{\prime}_{0}(m)$ in (vii) and (viii) are not guaranteed. Instead, we have $e_{0}^{\prime}\left(m^{+}\right)\leq\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$. Note $e_{0}^{\prime}\left(m^{-}\right)<0$ if $m>0$ due to Theorem 3.6 (ii). Although uniqueness result in general case is unknown, uniqueness can still be established under assumptions slightly weaker than the polytropic law; see [46, 16].
Remark 3.10.
Intuitively, when the distance between two celestial bodies is sufficiently large, the kinetic energy and the graviational interaction energy becomes negligible. Consequently, $E_{J}$ will approach $e_{0}$ without equaling it. This intuition is indeed correct [16, Remark 2.43 and Remark 2.44]. In Section 5, we will also compare $e_{0}$ and the energy $E_{0}(\widetilde{\rho}_{m})$ for scaling density $\widetilde{\rho_{m}}$. This comparison allows us to apply concentration-compactness lemma in [42] and will also provide a clearer mathematical understanding of the relation between $e_{0}$ and $E_{J}$.
It turns out under polytropic law (1) assumption, minimizers with different masses satisfy scaling relation, which helps to strengthen Theorem 3.6 (ii), as shown in the following results:
Theorem 3.11 (Relations between Minimizers and Minimal Energies with Different Mass [16, Theorem 3.2]).
Let $\sigma$ be the minimizer with mass 1 and the corresponding minimal energy is $e_{0}=E_{0}(\sigma)$, then any other minimizer with mass $m$ can be represented as $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$, where $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, and the corresponding minimal energy is $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$.
Sketch of Proof.
Recall (19) that $E_{0}(\sigma)=U(\sigma)-\frac{G(\sigma,\sigma)}{2}=\int A(\sigma)-\frac{1}{2}\iint\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y$, where $A(\sigma)=\frac{K}{\gamma-1}\sigma^{\gamma}$. Then let $\sigma_{\widetilde{A},B}=\widetilde{A}\sigma(Bx)$, we can compute
$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int A\left(\sigma_{\widetilde{A},B}\right)$ (27)
$=\int\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}$
$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}d^{3}(Bx)$
$=\widetilde{A}^{\gamma}B^{-3}U(\sigma)$
$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}d^{3}xd^{3}y$ (28)
$=\widetilde{A}^{2}B^{-5}\int\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}d^{3}(Bx)d^{3}(By)$
$=\widetilde{A}^{2}B^{-5}G(\sigma,\sigma)$
$\int\sigma_{\widetilde{A},B}$ $=\widetilde{A}B^{-3}\int\sigma(Bx)d^{3}(Bx)$ (29)
$=\widetilde{A}B^{-3}\int\sigma$
From the scaling relations (27) (28) (29), together with the uniqueness result Theorem 3.6 (i), if we take $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, then we conclude the proof. For a complete discussion, one can see [16, Theorem 3.2]. ∎
Remark 3.12.
Thanks to Theorem 3.6, we know $\|\sigma\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{1}$ and spt $\sigma$ is contained in a ball of radius $R_{1}$, therefore, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C_{m}=\frac{C_{1}}{A}$ and spt $\sigma_{m}$ is contained in a ball of radius $R_{m}=BR_{1}$. Since $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, if we further assume $\gamma>2$, we know $\lim\limits_{m\rightarrow 0}A=+\infty$ and $\lim\limits_{m\rightarrow 0}B=0$, thus $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ and the size of $\sigma_{m}$’s support will go to 0 when $m\rightarrow 0$, with rates $m^{\frac{2}{3\gamma-4}}$ and $m^{\frac{\gamma-2}{3\gamma-4}}$ respectively. When $\gamma<2$, we still have $\lim\limits_{m\rightarrow 0}A=+\infty$, but $\lim\limits_{m\rightarrow 0}B=\infty$, and the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread). It coincides with a result mentioned in Lieb and Yau’s paper [40, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$. In their paper quantum mechanics (fermions case) is discussed.
Remark 3.13.
Now we know the minimizer satisfies (EL). Actually the converse also holds true: all radial solutions of (EL) are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$. Furthermore, the minimizer (or radial solution to (EL)) can be uniquely parameterized by its central density, with each central density corresponding to a specific mass, see [16, Proposition 2.48]. Thanks to Theorem 3.6, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus Theorem 3.11 shows as the central density decreases, the mass also decreases. Moreover, it gives the decay rate. For more general pressure $P(\rho)$, this decreasing relationship is also established in [16, Corollary 2.41].
Let $\sigma_{\widetilde{A},B}=\widetilde{A}\sigma(Bx)$, with $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$ as above, actually we also have the following scaling relations:
$A^{\prime}\left(\sigma_{\widetilde{A},B}\right)(x)=A^{\prime}\left(\widetilde{A}\sigma\right)\left(Bx\right)=\widetilde{A}^{\gamma-1}A^{\prime}(\sigma)\left({B}x\right)=m^{\frac{2-2\gamma}{3\gamma-4}}A^{\prime}(\sigma)\left({B}x\right)$ (30)
$V_{\left(\sigma_{\widetilde{A},B}\right)}(x)=\frac{\widetilde{A}}{B^{2}}V_{\sigma}\left(Bx\right)=\widetilde{A}^{\gamma-1}V_{\sigma}\left({B}x\right)=m^{\frac{2-2\gamma}{3\gamma-4}}V_{\sigma}\left({B}x\right)$ (31)
$E_{0}^{\prime}\left(\sigma_{\widetilde{A},B}\right)(x)=\widetilde{A}^{\gamma-1}E_{0}^{\prime}(\sigma)\left(Bx\right)=m^{\frac{2-2\gamma}{3\gamma-4}}E_{0}^{\prime}(\sigma)\left({B}x\right)$ (32)
Therefore, we can further discuss the relationship of variational derivatives (see Lemma 2.6) between densities with different mass (not necessary to be minimizers).
Proposition 3.14 (Relations of Variational Derivatives Between Densities with Different Mass [16, Proposition 3.4]).
Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies $E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}$ $E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$, where $\rho(x):=\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, $\rho$ has mass 1.
4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
Similar to the proof of Theorem 3.6 (see e.g. [16]), instead of applying direct method to admissible class $W_{m}$ (21), we first construct a “double constrained” admissible class $W_{m,R}$ in which we can show the lower semi-continuity of the energy. We follow the definition of $W_{m,R}$ introduced in [46].
Definition 4.1 (“Double Constrained” Admissible Class).
Let $W_{m}$ be as defined in (21), we then define $W_{m,R}$ as the following:
$W_{m,R}:=\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in W_{m}\mid\rho(m)\leq R\text{ a.e.}\right\}$
Remark 4.2.
$W_{m,R}\subset L^{1}\cap L^{\infty}$ and it is not empty when $R$ is larger than $\frac{384}{\pi\eta^{3}}$, since we can pick $\rho_{m}(x)=\left\{\begin{array}[]{ll}\frac{384m}{\pi\eta^{3}},&\left|x-y_{m}\right|<\frac{\eta}{8}\\ 0,&\left|x-y_{m}\right|\geq\frac{\eta}{8}\end{array}\right$. and $\rho_{1-m}$ similarly, then check $\rho(m)=$ $\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$.
We want to prove the existence of “double constrained” minimizers via direct method [17, Theorem 1.15]. Here, “double constrained” means $\rho(m)\in W_{m}$ and $\rho(m)\leq R$ a.e.. To do that we first show the lower semicontinuity of $E_{J}$ by reformulating arguments in [4] where the authors only consider non-rotating one body case.
With an abuse of notation, we denote sequences $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and their limit $\rho_{R}$. Note their total mass $\int\rho_{n,R}$} is always one.
Lemma 4.3 (Lower Semicontinuity of $E_{J}$ in $W_{m,R}$ w.r.t the Weak Topology on $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$).
Fix $J>0,m>0$, if $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, then $\rho_{R}\in W_{m,R}$ and $E_{J}\left(\rho_{R}\right)\leq\liminf\limits_{n\rightarrow\infty}E_{J}\left(\rho_{n,R}\right)$.
Proof.
If $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$, then for almost every $x\in\mathbb{R}^{3},0\leq\rho_{n,R}\leq$ $R$ and up to subsequence $\rho_{n,R}\rightarrow\rho_{R}$. Then $0\leq\rho_{R}\leq R$ almost everywhere. Since $\Omega:=$ $\Omega_{m}\cup\Omega_{1-m}$ is bounded, by Hölder’s inequality we know $\left\|\rho_{n,R}-\rho_{R}\right\|_{L^{1}}\leq C\|\rho_{n,R}$- $\rho_{R}\|_{L^{\frac{4}{3}}}\rightarrow 0$. It implies $\rho_{R}\in W_{m,R}$ thus $W_{m,R}$ is closed. Easy to check it is also a convex subset of $L^{\frac{4}{3}}$. Therefore, $W_{m,R}$ is weakly closed.
To proof the lower semicontinuity, we recall $E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)=\int A(\rho)$- $\frac{\int V_{\rho}\cdot\rho}{2}+T_{J}(\rho)=\int\frac{K(\rho(x))^{\gamma}}{\gamma-1}d^{3}x-\frac{1}{2}\iint\frac{\rho(x)\cdot\rho(y)}{|x-y|}d^{3}xd^{3}y+\frac{J^{2}}{2(I(\rho))}$. We first notice $A(\rho)$ is a convex function, so $\int A(\rho)$ is a convex functional. Then for each constant $k,\left\{\rho_{R}\in\right$. $\left.W_{m,R}\mid\int A\left(\rho_{R}\right)\leq k\right\}$ is convex. $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$ implies almost everywhere convergence up to a subsequence, and by construction we know $\rho_{n,R}\leq R\cdot 1_{\Omega}$, here $1_{\Omega}$ is the indicator function of compact set $\Omega$. By (2) we know $A(0)=0$, $A(\cdot)$ is non-decreasing, and thus $A(\rho_{n,R})\leq A(R)\cdot 1_{\Omega}\in L^{1}$. Therefore, if $\int A\left(\rho_{n,R}\right)\leq k$ then $\int A\left(\rho_{R}\right)\leq k$ by dominated convergence theorem or Fatou’s lemma, which means $\left\{\rho_{R}\in W_{m,R}\mid\int A\left(\rho_{R}\right)\leq k\right\}$ is not only convex but also closed. Thus, it is also weakly closed and then $\int A(\rho)$ is a weakly lower semicontinuous functional. Next, thanks to Proposition A.5, we know $\mathcal{V}:\rho\rightarrow V_{\rho}$ is compact from $L^{\frac{4}{3}}(\Omega)$ to $L^{r}(\Omega)$ for any $1\leq r<12$, thus $\mathcal{V}$ is completely continuous. Take $r=11$, then $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}(\Omega)$ implies $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{11}(\Omega)$. Since $\Omega$ is bounded, by Holder’s inequality we know $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$. Since $\rho_{R}$ and $\rho_{n,R}$ vanish outside $\Omega$, we have
$\left|G\left(\rho_{R},\rho_{R}\right)-G\left(\rho_{n,R},\rho_{n,R}\right)\right|=$ $\left|\int_{\mathbb{R}^{3}}V_{\rho_{R}}\cdot\rho_{R}-\int_{\mathbb{R}^{3}}V_{\rho_{n,R}}\cdot\rho_{n,R}\right|$
$=$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|+\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}-\int_{\Omega}V_{\rho_{n,R}}\cdot\rho_{n,R}\right|$
$\leq$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|+\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}$
The last inequality comes from the Hölder’s inequality. Note $V_{\rho_{R}}\in L^{11}(\Omega)\subset L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$, together with $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ we know $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|\rightarrow 0$. Note $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ also implies $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq C$, thus $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}\rightarrow 0$ due to $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$. Hence we have $G\left(\rho_{n,R},\rho_{n,R}\right)\rightarrow G\left(\rho_{R},\rho_{R}\right)$. Finally, if $\rho\in W_{m,R}$, then $\bar{x}(\rho)=\frac{\int_{\mathbb{R}^{3}}\rho(x)x}{\int_{\mathbb{R}^{3}}\rho}=\int_{\Omega}\rho(x)x$, and $\rho_{n,R}\rightharpoonup\rho_{R}$ implies $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ since $f(x)=x\in L^{\infty}\left(\Omega;\mathbb{R}^{3}\right)\subset\left(L^{\frac{4}{3}}\left(\Omega;\mathbb{R}^{3}\right)\right)^{*}$. Moreover,
$\qquad\left|I\left(\rho_{n,R}\right)-I\left(\rho_{R}\right)\right|$
$=\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
$\leq\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|+\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
$\leq\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right\|_{L^{4}(\Omega)}+\left|\int_{\Omega}\left(\rho_{n,R}(x)-\rho_{R}(x)\right)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
By an argument similar to the above, we know $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)$ and $r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ are in $L^{\infty}(\Omega)\subset$ $L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$. $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ means $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\rightarrow r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ everywhere, thus by dominated convergence theorem we know $\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\|_{L^{4}(\Omega)}\rightarrow 0$. $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies weak convergence in $L^{\frac{4}{3}}(\Omega)$, $\left\|\rho_{n,R}\right\|_{L^{3}(\Omega)}\leq C$, then we have $I\left(\rho_{n,R}\right)\rightarrow I\left(\rho_{R}\right),\frac{J^{2}}{2\left(I\left(\rho_{n,R}\right)\right)}\rightarrow\frac{J^{2}}{2\left(I\left(\rho_{R}\right)\right)}$. Collect the results above we obtain $E_{J}\left(\rho_{R}\right)\leq\liminf\limits_{n\rightarrow\infty}E_{J}\left(\rho_{n,R}\right)$. ∎
Remark 4.4.
Several results can be established using similar arguments:
  1. (i)
    The functional $E_{J}$ is lower semicontinuous with respect to the weak topology on $L^{p}(\mathbb{R}^{3})$ for $\frac{6}{5}<p\leq\infty$.
  2. (ii)
    $E_{J}$ is continuous in $W_{m,R}$ with respect to the norm topology on $L^{p}(\mathbb{R}^{3})$, where $\frac{6}{5}<p\leq\infty$.
  3. (iii)
    The non‑rotating energy $E_{0}$ is lower semicontinuous in $W_{m}$ (not only in $W_{m,R}$) with respect to the norm topology on $L^{\frac{4}{3}}(\mathbb{R}^{3})$. The proof follows the same line of reasoning as in the proof of Theorem 5.9 below.
  4. (iv)
    We note that a different energy functional $E$ is studied in [4] for a single‑star model.
We then have the existence of “double constrained” minimizers, and results about their variational derivatives $E_{J}^{\prime}$ (18).
Theorem 4.5 (Existence of “Double Constrained” Minimizers).
Fix $J>0$, give $m>0,R>\frac{384}{\pi\eta^{3}}$, there is a $\rho_{R}\in W_{m,R}$ which minimizes $E_{J}(\rho)$ on $W_{m,R}$. Any such $\rho_{R}$ is continuous on $\Omega=\Omega_{m}\cup\Omega_{1-m}$. Moreover, there are some constants $\lambda_{m,R}$ and $\lambda_{1-m,R}$ such that
[…]
In particular,
[…]
Proof.
Given a minimizing sequence $\left\{\rho_{n,R}\right\}$ with $\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$. By Interpolation Inequality Proposition A.2 we know $W_{m,R}$ is bounded in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. Since $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ is reflexive, $\exists\rho_{R}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ such that $\rho_{n,R}\rightharpoonup\rho_{R}$ up to subsequence [8, Theorem 3.18]. By Lemma 4.3 $W_{m,R}$ is weakly compact in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right),\rho_{R}\in W_{m,R}$, and $\inf\limits_{\rho\in W_{m,R}}E(\rho)\leq E\left(\rho_{R}\right)\leq\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$, which means $\rho_{R}$ is a minimizer. A slight modification of the proof of [4, Lemma 2], [15, Proposition 5.8] and [16, Lemma 2.21], which means we need to consider the perturbations have compact supports in $\Omega$, shows that (33) and (34) hold for almost all points which are mentioned in (33) (34) with suitable $\lambda_{m,R}$ or $\lambda_{1-m,R}$. The continuity of $\rho_{R}$ on $\Omega$ then follows as in the proof of Theorem 3.6 (see e.g. [15, Theorem 2.25] and [16, Remark 2.15]), where the idea is making use of (33) (34) (35) and improving the regularity of potential $V_{\rho_{R}}$ and density $\rho_{R}$ in turns (bootstrap). Then we know (33) and (34) hold indeed for all points we considered. ∎
Remark 4.6.
Another type of bootstrap, commonly used in fluid dynamics and general relativity, refers to bootstrapping with respect to time. This differs from the regularity bootstrap discussed in our paper.
We modify the arguments in [4, Section 6] and apply again bootstrap method to show that there is a $\widetilde{R}$ such that the “double constrained” (“DC” in short) minimizer $\rho_{R}$ remains in $W_{m,\widetilde{R}}$ for all $R\geq\widetilde{R}$. Then we argue that $\rho_{R}$ minimizes $E_{J}(\rho)$ on $W_{m}$ to prove the existence of the constrained minimizers.
Lemma 4.7 (Uniform Bound of DC Minimizers in $L^{\frac{4}{3}}$ Space).
There is a constant $k_{0}$ such that $\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}}\leq k_{0}$, all $R\geq R_{0}=\frac{384}{\pi\eta^{3}}$.
Proof.
We have $\forall R\in\left[R_{0},\infty\right),\rho_{R_{0}}\in W_{m,R}$, then $E\left(\rho_{R}\right)\leq E\left(\rho_{R_{0}}\right)<\infty$. Then, one can adapt the arguments from the non-rotating case—see, for instance, [16, Lemma 2.21] or [4, Section 6]—to show that the $L^{\frac{4}{3}}$ norm of $\rho_{R}$ is actually bounded by a constant depends only on $E\left(\rho_{R_{0}}\right)$. ∎
Deducing the uniform boundedness in $L^{\infty}$ from that in $L^{\frac{4}{3}}$ can be accomplished in several ways. For instance, one may refer to the methods of Li in [38, Proposition 1.4], as McCann mentioned in [46, Section 6]. Li’s method [38] can be applied to obtain uniform boundedness in $L^{\infty}$ for $\gamma>\frac{4}{3}$ by choosing an appropriate test function, although it does not first verify whether Lagrange multiplier is negative — Li verified this after the $L^{\infty}$ bound was obtained in [38].
Here, we present an alternative approach, which is inspired by the method in [4]. Both methods rely on the result that velocity $v^{2}(x)=\frac{J^{2}r^{2}(x-\bar{x}(\rho))}{I^{2}(\rho)}$ is sufficiently small. Under assumption $\gamma>\frac{3}{2}$, we first prove that Lagrange multipliers are negative, using the idea from [4, Lemma 5] and [38, Proposition 2.5].
Lemma 4.8 (Negavie Lagrange-multipliers for DC Minimizers).
$\exists m_{0}>0,\forall 0<m<m_{0},\exists R_{1}>0$, such that $\forall R>R_{1},\lambda_{m,R}<0$, $\lambda_{1-m,R}<0$
Proof.
For $\lambda_{m,R}$, since the volume (measure) of $\Omega_{m}$ is ${Vol}:=\frac{4}{3}\pi\left(\frac{\eta}{4}\right)^{3}=$ $\frac{\pi J^{6}}{48m^{6}(1-m)^{6}}$, while the mass in $\Omega_{m}$ is $\int_{\Omega_{m}}\rho_{R}=m$. Claim: $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$. In fact, if not, then $\rho_{R}(x)-\frac{m}{Vol}>0$ almost everywhere in $\Omega_{m}$, while $\int_{\Omega_{m}}\left(\rho_{R}(x)-\frac{m}{Vol}\right)=0$ implies $\rho_{R}(x)-\frac{m}{Vol}=0$ almost everywhere in $\Omega_{m}$ [51, Section 4.3]. Since the measure of $\Omega_{m}$ is not 0, it makes a contradiction. Now we have $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$, which means there is at least a point $x_{0}$ in $\Omega_{m}$ such that $\rho_{R}\left(x_{0}\right)\leq\frac{m}{Vol}$ and (33) holds true. Notice ${Vol}\sim m^{-6}$ as $m\ll 1$, i.e. $\exists C_{1}>C_{2}>0$, such that if $m$ is sufficiently small, then $C_{2}m^{-6}\leq{Vol}\leq C_{1}m^{-6}$. Therefore, $\frac{m}{Vol}\sim$ $m^{7},A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\left(x_{0}\right)\stackrel{{\scriptstyle\gamma>\frac{3}{2}}}{{=}}O\left(m^{\frac{7}{3}}\right)$. By construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\forall y\in\Omega_{1-m},\left|x_{0}-y\right|\leq\operatorname{diam}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}},V_{\rho_{R}}\left(x_{0}\right)\geq$ $\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{\left|x_{0}-y\right|}d^{3}y\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)$ when $0<m<m_{1}$ for some $m_{1}$. Let $R>\frac{m_{1}}{{Vol}\left(m_{1}\right)}$, then $\rho_{R}(x_{0})$ that satisfies the above relation exists, and (33) holds true thus
$\lambda_{m,R}$ $\leq E_{J}^{\prime}\left(\rho_{R}\right)\left(x_{0}\right)$
$=A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)-V_{\rho_{R}}\left(x_{0}\right)-\frac{J^{2}r^{2}\left(x_{0}-\bar{x}\left(\rho_{R}\right)\right)}{I^{2}(\rho)}$
$\leq A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)-V_{\rho_{R}}\left(x_{0}\right)<0$
[…]
Let $R>\frac{m_{2}}{{Vol}\left(m_{2}\right)}$, we have $\lambda_{1-m,R}<0$. Set $m_{0}=\min\left\{m_{1},m_{2}\right\},R_{1}=\max$ $\left\{\frac{m_{1}}{{Vol}\left(m_{1}\right)},\frac{m_{2}}{{Vol}\left(m_{2}\right)}\right\}$, then we get the conclusion. ∎
Lemma 4.9 (Moment of Inertia Increasing Rate).
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall 0<m<M$ and $\forall R>0$, $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $I\left(\rho_{1-m}+\rho_{m}\right)>{Cm}^{-3}$.
Proof.
$\rho(m)=\rho_{1-m}+\rho_{m}\in{W}_{m,R}$. The centers of mass $\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m},\bar{x}\left(\rho_{m}\right)\in\Omega_{m}$, as well as their projections onto $z=0$, are separated by at least $\frac{\eta}{2},\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$. Due to Lemma 2.4, we have
$I(\rho(m))=\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+I\left(\rho_{1-m}\right)+I\left(\rho_{m}\right)\geq\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)\geq\frac{J^{4}}{4\mu^{3}}$
Since $\mu=m(1-m)$, we know $\exists C>0$, such that $I(\rho(m))>Cm^{-3}$ when $m\rightarrow 0$. ∎
Corollary 4.10 (Velocity Decreasing Rate).
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $0<m<M$, $R>0$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $v(x,m)=\frac{Jr(x-\bar{x}(\rho(m)))}{I(\rho(m))}<Cm$.
Proof.
By construction we know $\bar{x}(\rho(m))\in\text{conv}(\Omega_{m}\cup\Omega_{1-m})$, where conv menas the convex hull. Moreover, $r(x-\bar{x}(\rho(m)))\leq$ $\frac{3}{2}\eta$ for all $x\in\Omega_{m}\cup\Omega_{1-m}$. Thus, when $m$ small enough, by Lemma 4.9 we have $v(x,m)=\frac{Jr(x-\bar{x}(\rho(m)))}{I(\rho(m))}<Cm$ for some $C>0$. ∎
Theorem 4.11 (Uniform Bound for DC Minimizer in $L^{\infty}\left(\mathbb{R}^{3}\right)$).
Fix $J>0,\exists M>0$, such that $\forall 0<m<M,\exists k_{1}>0,\forall R>0$ and $\rho_{R}$ is a “double constrained” minimizer in $W_{m,R}$, we have $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}<k_{1}$. In particular, $E_{J}(\rho_{R})=E_{J}(\rho_{k_{1}})$ for $R\geq k_{1}$.
Remark 4.12.
Here we assume that the inequality $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}<k_{1}$ holds automatically if $W_{m,R}$ is empty. But due to Remark 4.2, $W_{m,R}$ is not empty for large $R$.
Proof of Theorem 4.11.
We want to apply bootstrap method. Thanks to Corollary 4.10, when $m$ is sufficiently small, $\exists C_{1}>0$, such that $\forall x\in\Omega=\Omega_{m}\cup\Omega_{1-m},v^{2}(x)\leq C_{1}m^{2}$. Moreover, similarly as above if $x\in\Omega$, we have $\forall y\in\Omega_{1-m},|x-y|\leq\operatorname{diam}\left(\Omega_{m}\cup\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}},V_{\rho_{R}}(x)\geq$ $\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{|x-y|}d^{3}y\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>C_{2}m^{2}$ for some $C_{2}>0$ when $m$ is small enough. Therefore, we know $\exists\widetilde{C}>0$, such that $\forall x\in\Omega$, $v^{2}(x)\leq\widetilde{C}V_{\rho_{R}}(x)$. By (34), and Lemma 4.8 (Lagrange-multipliers are less then 0), when $R>R_{1}$, we have if $\rho_{R}(x)>0$, then $A^{\prime}\left(\rho_{R}(x)\right)\leq V_{\rho}(x)+v^{2}(x)\leq CV_{\rho}(x)$. If $\rho_{R}(x)=0$, the $0=$ $A^{\prime}\left(\rho_{R}(x)\right)\leq CV_{\rho_{R}}(x)$ is obvious.
Thanks to Lemma 4.7, we know if $R>R_{0}+R_{1},\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}}\leq k_{0}$. By HardyLittlewood-Sobolev Inequality (Proposition A.3), with an abuse of notation $C$, we have $\left\|V_{\rho_{R}}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}\leq$ $C\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\leq C$. $A^{\prime}\left(\rho_{R}\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\leq CV_{\rho}$ implies $\left\|\rho_{R}\right\|_{L^{12(\gamma-1)}}=$ $\left(\int\rho_{R}{}^{12(\gamma-1)}(x)d^{3}x\right)^{\frac{1}{12(\gamma-1)}}\leq C\left(\int V_{\rho_{R}}^{12}(x)d^{3}x\right)^{\frac{1}{12(\gamma-1)}}=C\left\|V_{\rho_{R}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C$. Since we know $\gamma>\frac{3}{2}>\frac{4}{3}$ and $\Omega$ is bounded, by interpolation inequality (Proposition A.2) we have $\left\|\rho_{R}\right\|_{L^{6}}\leq C$. Since we also know $\left\|\rho_{R}\right\|_{L^{1}\left(\mathbb{R}^{3}\right)}=1$ are bounded, thanks to Proposition A.4, we have $V_{\rho_{R}}$ is continuous and $\left\|V_{\rho_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is bounded uniformly. Again by $\rho_{R}{}^{\gamma-1}\leq CV_{\rho_{R}}$ we know $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq\hat{C}$ for some $\hat{C}>0$ when $R>R_{1}+R_{0}$. When $R\leq R_{1}+R_{0}$, by construction of $W_{m,R}$ we know $\left\|\rho_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq R_{1}+R_{0}$. Pick $k_{1}=\max\left\{\hat{C},R_{1}+R_{0}\right\}+1$ and then we get the result. For $R\geq k_{1}$, $E_{J}(\rho_{R})=E_{J}(\rho_{k_{1}})$ comes from the fact that they both minimize $E_{J}$ on $W_{m,k_{1}}$. ∎
Remark 4.13.
One can observe in $A^{\prime}\left(\rho_{R}(x)\right)\leq CV_{\rho_{R}}(x)$, such $C$ is actually not only independent of $R$ but also a uniform constant when $0<m<m_{0}$ for some $m_{0}$. This inequality also holds for constrained minimizer $\rho$, not only for a double constrained minimizer. The existence of a constrained minimizer follows from Theorem 4.14 below.
Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, the $L^{\infty}$ norm of $\sigma$ may be unbounded, hence in order to show the “double constrained minimizer” $\rho_{R}$ is indeed a constrained minimizer for $R$ large enough, we need to prove that energy of $\sigma$ still exceeds energy of $\rho_{R}$, which was not addressed in McCann’s work [46]. We provides a supplementary proof for it by adapting the arguments in the single-star case [4, Section 6].
Theorem 4.14 (Existence of the Constrained Minimizers).
Fix $J>0,\exists m_{0}>0,\forall 0<m<m_{0}$, there is a $\rho\in W_{m}$ which minimizes $E_{J}(\rho)$ on $W_{m}$. Moreover, there is a constant $R(m)$ such that for any constrained minimizer $\rho$, $\rho$ is continuous on $\Omega=\Omega_{m}\cup\Omega_{1-m}$ and $\rho<R(m)$, that is $\rho\in W_{m,R(m)}$.
Proof.
The facts that $\rho$ is continuous on $\Omega$ and $\rho<R(m)$ can be handled similarly to the case for “double constrained” minimizer, see Theorem 4.5, Theorem 4.11, and Remark 4.13.
To prove the existence of constrained minimizer, let $k_{1}$ is given in Theorem 4.11, we have $E_{J}\left(\rho_{R}\right)=E_{J}\left(\rho_{k_{1}}\right)$ for $R\geq k_{1}$ due to Theorem 4.11. We will prove $\rho_{k_{1}}$ is a constrained minimizer. First we note that $\rho_{k_{1}}\in W_{m,k_{1}}\subset L^{1}(\mathbb{R}^{3})\cap L^{\infty}(\mathbb{R}^{3})$, then $E_{J}(\rho_{k})<\infty$ due to Remark 2.5. Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, due to Remark 2.2 we know $G(\sigma,\sigma)<\infty$. For $\sigma$ with $U(\sigma)=\int A(\sigma(x))d^{3}x=\infty$, we know $E_{J}(\rho_{k_{1}})<E_{J}(\sigma)=\infty$. For $\sigma$ with $U(\sigma)<\infty$, let
[…]
When $R$ is large enough, we can find a $\tau\in\widetilde{P_{R}}(\sigma)$, with $\tau\geq 0$ and $\int_{\Omega_{m}}\tau=\int_{\Omega_{1-m}}\tau=1$. (See e.g. [15, Remark 4.10].) Let $\sigma_{s}^{[1]}(x),\sigma_{s}^{[2]}(x)$ and $\sigma_{s}$ be
$\sigma_{s}^{[1]}(x)=\left\{\begin{array}[]{ll}\sigma(x),&\text{ if }x\in\Omega\text{ and }\sigma(x)<\frac{1}{2}s\\ 0,&\text{ otherwise }\end{array}\right$.
$\sigma_{s}^{[2]}(x)=\left\{\begin{array}[]{ll}\left(m-\int_{\Omega_{m}}\sigma_{s}^{[1]}\right)\tau,&\text{ if }x\in\Omega_{m}\\ \left((1-m)-\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\right)\tau,&\text{ if }x\in\Omega_{1-m}\\ 0,&\text{ otherwise }\end{array}\right$.
$\sigma_{s}=\sigma_{s}^{1}+\sigma_{s}^{2}$
By construction we have $\int_{\Omega_{m}}\sigma_{s}=m$ and $\int_{\Omega_{1-m}}\sigma_{s}=1-m$, then $\sigma_{s}\in W_{m}$. Similar again to [15, Remark 4.10], we have $\int_{\Omega_{m}}\sigma_{s}^{[1]}\rightarrow\int_{\Omega_{m}}\sigma=m$, $\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\rightarrow\int_{\Omega_{1-m}}\sigma=1-m$, therefore $\sigma_{s}^{[2]}\rightarrow 0$ almost everywhere in $\mathbb{R}^{3}$ as $s\rightarrow\infty$ due to $\tau\in L^{\infty}$. $\sigma\in W_{m}$ implies $\sigma$ is finite almost everywhere, thus $\sigma_{s}^{[1]}$ increases to $\sigma$ almost everywhere as $s\rightarrow\infty$. Then we know when $s\rightarrow\infty$, $\sigma_{s}\rightarrow\sigma$ a.e. in $\mathbb{R}^{3}$. Moreover, for almost every point $x\in\left\{x\in\mathbb{R}^{3}\mid\tau(x)=0\right\}$, $A\left(\sigma_{s}\right)=A\left(\sigma_{s}^{[1]}\right)$ increases to $A(\sigma)$ (except at points where $\tau=\infty$). For $\Omega_{>0}\coloneq\{x\in\mathbb{R}^{3}\mid\tau(x)>0\}$ (a subset of $\Omega$ thus bounded), $\sigma_{s}\rightarrow\sigma$ implies $A\left(\sigma_{s}\right)\rightarrow A(\sigma)$ a.e. as $s\rightarrow\infty$, and we also know $\left|\sigma_{s}\right|\leq\left|\sigma_{s}^{[1]}\right|+\left|\sigma_{s}^{[2]}\right|\leq\sigma+\|\tau\|_{L^{\infty}}$ a.e., which implies $A(\sigma_{s})\leq A(\sigma+\|\tau\|_{L^{\infty}})$ a.e. We claim $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$. In fact, since $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ is convex, we know
$A(\sigma+\|\tau\|_{L^{\infty}})=A(\frac{2\sigma+2\|\tau\|_{L^{\infty}}}{2})\leq\frac{A(2\sigma)+A(2\|\tau\|_{L^{\infty}})}{2}$
Note $U(\sigma)=\int A(\sigma(x))d^{3}x<\infty$ implies $A(2\sigma)\in L^{1}(\mathbb{R}^{3})$, and $\Omega_{>0}$ is bounded implies $A(2\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$. Therefore, we know $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$.
We can then split $\mathbb{R}^{3}$ into $\mathbb{R}^{3}\setminus\Omega_{>0}$ and $\Omega_{>0}$, and show $\int_{\mathbb{R}^{3}}A\left(\sigma_{s}\right)\rightarrow\int_{\mathbb{R}^{3}}A(\sigma)$ as $s\rightarrow\infty$ by the monotone convergence theorem and dominated convergence theorem. Since $\Omega$ is bounded, by Hölder’s inequality we know $\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq C\|\tau\|_{L^{\infty}(\Omega)}$, then $\left\|\sigma_{s}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\left\|\sigma_{s}^{[1]}\right\|_{L^{\frac{4}{3}}(\Omega)}+\left\|\sigma_{s}^{[2]}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq$ $\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+C\|\tau\|_{L^{\infty}(\Omega)}$. It means $\left\{\sigma_{s}\right\}$ is a bounded set in $L^{\frac{4}{3}}(\Omega)$. Since $L^{\frac{4}{3}}(\Omega)$ is reflexive, thanks to [8, Theorem 3.18], there is a sequence $\left\{\sigma_{s_{n}}\right\}$ which converges weakly to a weak limit $\widetilde{\sigma}$. Since we know $\sigma_{s_{n}}\rightarrow\sigma$ a.e., by [55], we have $\widetilde{\sigma}=\sigma$. That is, $\sigma_{s_{n}}\rightharpoonup\sigma$ in $L^{\frac{4}{3}}(\Omega)$, therefore, we can apply what we did in the proof of Lemma 4.3, and obtain $G\left(\sigma_{s_{n}},\sigma_{s_{n}}\right)\rightarrow G(\sigma,\sigma),T_{J}\left(\sigma_{s_{n}}\right)\rightarrow T_{J}(\sigma)$. Collect the results above, we have $E_{J}(\sigma)=\lim\limits_{n\rightarrow\infty}E_{J}\left(\sigma_{s_{n}}\right)$. By the same arguments, we know for any sequence of $\left\{\sigma_{s_{m}}\right\}$, we can always pick a subsequence $\left\{\sigma_{s_{m_{k}}}\right\}$ which also satisfies $E_{J}(\sigma)=\lim\limits_{k\rightarrow\infty}E_{J}\left(\sigma_{s_{m_{k}}}\right)$. Thus, we get, $E_{J}(\sigma)=\lim\limits_{s\rightarrow\infty}E_{J}\left(\sigma_{s}\right)$ (one can show it by contradiction argument).
Since for almost every $x\in\mathbb{R}^{3}$, $\left|\sigma_{s}\right|\leq\left|\sigma_{s}^{[1]}\right|+\left|\sigma_{s}^{[2]}\right|<\frac{1}{2}s+\|\sigma\|_{L^{\infty}}$. For $s>\max\{2\|\sigma\|_{L^{\infty}},k_{1}\}$, we have $\left|\sigma_{s}\right|<s$ a.e., thus $\sigma_{s}\in W_{m,s}$, and then $E_{J}\left(\sigma_{s}\right)\geq E_{J}\left(\rho_{s}\right)=E_{J}\left(\rho_{k_{1}}\right)$ due to Theorem 4.11. Let $s\rightarrow 0$, then we know $E_{J}(\sigma)\geq E_{J}\left(\rho_{k_{1}}\right)=E\left(\rho_{R}\right)$ for $R\geq k_{1}$ by Theorem 4.11. Since we choose $\sigma\in W_{m}$ arbitrarily, we know $\rho_{R}$ is a constrained minimizer (i.e. $\rho_{R}$ minimizes $E_{J}$ on $W_{m}$) for $R\geq k_{1}$. In particular, we find that $\rho_{k_{1}}$ is a constrained minimizer. This gives the result. ∎
Remark 4.15.
In the proof of Theorem 4.14, the verification that $A(\sigma+|\tau|{L^{\infty}})\in L^{1}(\Omega{>0})$ relies only on the convexity of $A$, not on its specific form. Consequently, the same argument is applicable to more general pressure. For the polytropic case studied in this paper, one may also use the following estimate: $A(\sigma+\|\tau\|_{L^{\infty}})=\frac{K}{\gamma-1}(\sigma+\|\tau\|_{L^{\infty}})^{\gamma}\leq\frac{K}{\gamma-1}(2\max\{\sigma,\|\tau\|_{L^{\infty}}\})^{\gamma}=\frac{K}{\gamma-1}(2^{\gamma}\max\{\sigma^{\gamma},\|\tau\|_{L^{\infty}}^{\gamma}\})\leq\frac{K2^{\gamma}}{\gamma-1}(\sigma^{\gamma}+\|\tau\|_{L^{\infty}}^{\gamma})=2^{\gamma}(A(\sigma)+A(\|\tau\|_{L^{\infty}}))$.
Remark 4.16.
Following the statement in Section 5, one can show that the constant $R(m)$ is actually a uniform constant for all $m\in(0,m_{0})$ for some small $m_{0}$.
Remark 4.17.
Note that the continuity of $\rho$ has so far been established only on $\Omega$, not on the entire $\mathbb{R}^{3}$. This would suggest that there might be a sense of discontinuous density transition around the boundary for stars or planets. However, by additionally showing that the support lies in the interior, we will rule out this scenario and further extend the result to continuity on $\mathbb{R}^{3}$. Moreover, we have established the existence of a constrained minimizer, one can verify that such a minimizer satisfies the Euler–Lagrange equations (EL) on $\Omega$. However, to conclude that the reduced Euler–Poisson equations (EP’) hold in $\mathbb{R}^{3}$, we need to show the existence of $\nabla P(\rho)$ in $\mathbb{R}^{3}$, particularly on $\partial\{\rho>0\}$. It turns out $W^{\infty}$ local energy minimizers can do this job as we mention in Remark 3.4, and we have Theorem 3.1. Thanks to Remark 2.11, it suffices to show that the minimizer can be placed inside the interior of $\Omega$, which will be discussed in the following sections.
5 Bound of the Supports of Density Functions
In this section we estimate a uniform bound for the supports of the density functions, under small mass ratio assumption which is desirable physically in star-planet systems. The analysis focuses on the planet’s density. For stars densities, the estimates are similar or even simpler, and can be adapted directly from the arguments in [46] and does not require the scaling method in subsection 5.1 and subsection 5.2. With these uniform bounds we will show in Section 6 that the minimizers can be located in the interior of the domains we define.
5.1 Convergence of scaling densities in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$
As the mass ratio $m$ goes to 0, the energy of planet $\rho_{m}$ goes to 0 (see Remark 5.12), hence the concentration-compactness lemma in [42] cannot be directly applied. Therefore, we introduce scaling method. Given the constrained minimizers $\rho(m)=\rho_{m}+\rho_{1-m}$ of $E_{J}(\rho)$, we show that, after suitable scaling and translation, the scaling planet densities $\widetilde{\rho_{m}}$, which have mass $1$, will converge to some limit function as $m\rightarrow 0$ by showing the energies will converge. Intuitively, when $m\rightarrow 0$, the distances between star and planet become large. As a result, the gravitational influence they exert on each other becomes relatively weak, even after scaling. Therefore, the limit function turns out to be minimizer of $E_{0}(\rho)$ with respect to mass $1$, as we will prove later.
Definition 5.1 (Scaling Density for Planet).
For any given planet’s mass $m>0$, we define scaling density $\widetilde{\rho_{m}}$ as following
$\widetilde{\rho_{m}}(x)=A\rho_{m}\left({Bx}\right)$ (36)
where $\rho_{m}$ is the planet’s density of the constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ in Section 4, AA,$B$ are the same coefficients introduced in Theorem 3.11, that is, $A(m)=m^{-\frac{2}{3\gamma-4}}$, $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$.
Remark 5.2.
[…]
Moreover, we have
$\quad E_{J}\left({\rho_{1-m}+\rho_{m}}\right)$ (37)
$={\int{\frac{K{\left({\rho_{1-m}(x)+\rho_{m}(x)}\right)}^{\gamma}}{\gamma-1}d^{3}x}}-\frac{1}{2}{\iint{\frac{\rho(m)(x)\cdot\rho(m)(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}+\frac{J^{2}}{2\left({I\left({\rho_{1-m}+\rho_{m}}\right)}\right)}$
$=E_{0}\left(\rho_{1-m}\right)+E_{0}\left(\rho_{m}\right)-G\left({\rho_{1-m},\rho_{m}}\right)+T_{J}\left({\rho_{1-m}+\rho_{m}}\right)$
$=E_{0}\left(\rho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)-G\left({\rho_{1-m},\rho_{m}}\right)+T_{J}\left({\rho_{1-m}+\rho_{m}}\right)$
$>E_{0}\left(\rho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)-G\left({\rho_{1-m},\rho_{m}}\right)$
Notice since $\rho_{1-m}\in\Omega_{1-m}$, $\rho_{m}\in\Omega_{m}$, which distance is at least $\frac{\eta}{2}=\frac{J^{2}}{2m^{2}(1-m)^{2}}$ (24), we obtain $G\left({\rho_{1-m},\rho_{m}}\right)\leq\frac{2m\left({1-m}\right)}{\eta}<Cm^{3}$ for some $C>0$.
[…]
Lemma 5.3 (Moment of Inertia Increasing Rate as Mass Decreases).
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $I\left(\rho_{1-m}+\rho_{m}\right)>Cm^{-3}$.
Proof.
The proof is similar as in the proof of Lemma 4.9. ∎
Corollary 5.4 (Kinetic Energy Decreasing Rate as Mass Decreases).
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $T_{J}(\rho(m))<Cm^{3}$.
Proof.
The result can be obtained from Lemma 5.3 and the definition of $T_{J}(\rho(m))=\frac{J^{2}}{2I(\rho(m))}$. ∎
Not only the kinetic energy, we can also estimate velocity $v(x,m)=\frac{Jr(x-\overline{x}({\rho(m)}))}{I(\rho(m))}$, which will be useful later.
Corollary 5.5 (Velocity Decreasing Rate as Mass Decreases).
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $v(x,m)<Cm$.
Proof.
The proof is similar as in the proof of Lemma 4.10. ∎
Remark 5.6.
One can show the Lemma 5.3, Corollary 5.4 and Corollary 5.5 actually hold true for any $\rho\in W_{m}$.
[…]
$\varrho(m)(x)=\varrho_{1-m}(x)+\varrho_{m}(x)=\sigma_{1-m}\left({x-y_{1-m}}\right)+\sigma_{m}\left(x-y_{m}\right)$ (38)
$y_{1-m}$ and $y_{m}$ are the centers of $\Omega_{1-m}$ and $\Omega_{m}$. $\sigma_{1-m}$ and $\sigma_{m}$ are the non-rotating minimizers in Theorem 3.6 with mass (1m)(1-m) and $m$. In particular, $E_{0}(\varrho_{m})=E_{0}(\sigma_{m})\leq E_{0}(\rho_{m})$, since translations do not change the non-rotating energy. Similarly, $E_{0}(\varrho_{1-m})=E_{0}(\sigma_{1-m})\leq E_{0}(\rho_{1-m})$.
Remark 5.7.
By the scaling method and uniqueness of the minimizer, we know $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$ (see Theorem 3.11), where $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$ are the same as before, $\sigma$ is the non-rotating minimizer with mass 1. Similarly, $\sigma_{1-m}=\frac{1}{\widetilde{A}}\sigma\left(\frac{1}{\widetilde{B}}x\right)$, where $\widetilde{A}=(1-m)^{-\frac{2}{3\gamma-4}}$, $\widetilde{B}=(1-m)^{\frac{\gamma-2}{3\gamma-4}}$. Moreover, $\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\varrho_{m}\right)$, $\frac{{\widetilde{B}}^{3}}{{\widetilde{A}}^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{1-m}\right)=E_{0}\left(\varrho_{1-m}\right)$.
Proposition 5.8 (Energy Converges to Non-Rotating Minimum).
Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have
$\begin{split}e_{0}\leq E_{0}\left(\widetilde{\rho_{m}}\right)<e_{0}+Cm^{\frac{4\gamma-6}{3\gamma-4}}\end{split}$ (39)
where $e_{0}$ is the infimum of $E_{0}(\rho)$ with mass 1, that is, $e_{0}=E_{0}(\sigma)$. Since $\gamma>\frac{3}{2}$, in particular we have ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{m}}\right)}}=e_{0}$.
Proof.
[…]
$E_{J}\left({\varrho_{m}+\varrho_{1-m}}\right)-E_{0}\left(\varrho_{1-m}\right)-E_{0}\left(\varrho_{m}\right)=-G\left({\varrho_{1-m},\varrho_{m}}\right)+T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)$ (40)
Again, by Theorem 3.6, $\varrho_{1-m}$ and $\varrho_{m}$ have some nice properties, in particular they are spherically symmetry and radially decreasing after translations. Therefore, by comparison with the point masses, we get $G\left({\varrho_{1-m},\varrho_{m}}\right)=\mu\eta^{-1}=\frac{\mu^{3}}{J^{2}}$ (one can verify this by using Legendre polynomials or refer to [16, Corollary 2.31]), and $T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)\leq\frac{\mu^{3}}{2J^{2}}$ similarly as in Corollary 5.4. Thus, the right side of (40) is less than 0. This observation, together with Remark 5.7 gives us that:
$E_{0}\left(\varrho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)$ $=E_{0}\left(\varrho_{1-m}\right)+E_{0}\left(\varrho_{m}\right)$
$>E_{J}\left({\varrho_{m}+\varrho_{1-m}}\right)$
$\geq E_{J}\left({\rho_{m}+\rho_{1-m}}\right)$
$>E_{0}\left(\rho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)-G\left({\rho_{1-m},\rho_{m}}\right)$
$\geq E_{0}\left(\varrho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)-G\left({\rho_{1-m},\rho_{m}}\right)$
Therefore, we get $\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)<G\left({\rho_{1-m},\rho_{m}}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)$, or $E_{0}\left(\widetilde{\rho_{m}}\right)<\frac{A^{\gamma}}{B^{3}}G\left({\rho_{1-m},\rho_{m}}\right)+E_{0}(\sigma)$. Since $\frac{A^{\gamma}}{B^{3}}=m^{\frac{6-5\gamma}{3\gamma-4}}$, $G\left({\rho_{1-m},\rho_{m}}\right)\leq Cm^{3}$ due to Remark 5.2, we have
$E_{0}\left(\widetilde{\rho_{m}}\right)<Cm^{\frac{4\gamma-6}{3\gamma-4}}+E_{0}(\sigma)$
as we expect. ∎
Finally, we can show the convergence of $\left\{\widetilde{\rho_{m}}\right\}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$.
Theorem 5.9 (Convergence of Scaling Densities of Planets in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$).
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A\rho_{m}\left({Bx}\right)$ as in Definition 5.1, we have, up to translation, $\{\widetilde{\rho_{m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$. Here $\sigma$ is the non-rotating minimizer with mass 1.
Proof.
Due to Proposition 5.8, we know ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{m}}\right)}}=e_{0}<0$. Moreover, we can apply the concentration-compactness lemma of Lions [42, Theorem II.2 and Corollary II.1] to get that $\{\widetilde{\rho_{m}}\}$ is relatively compact up to translations in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$. Thus, there is a subsequence $\{\widetilde{\rho_{m_{k}}}\}$ that converges, up to translations, to some function $f$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, $L^{1}\left(\mathbb{R}^{3}\right)$ and almost everywhere in $\mathbb{R}^{3}$. We denote again the subsequence after translations by $\left\{\widetilde{\rho_{m_{k}}}\right\}$. In particular, $\left\|f\right\|_{L^{1}}=1$.
And we claim that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}$. In fact, we recall $E_{0}(\rho)={\int{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}-\frac{1}{2}{\iint{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}$. For the inertial energy $\int{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}$, by $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$. a.e. and Fatou’s Lemma we have ${\int{\frac{K\left({f(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}\leq\lim\limits_{k\rightarrow\infty}\inf{\int{\frac{K\left({\widetilde{\rho_{m_{k}}}(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}$. For the potential energy $\frac{1}{2}{\iint{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}=\frac{1}{2}{\int{\rho(x)V_{\rho}(x)d^{3}x}}$, we have
$\begin{split}&\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &=\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}+{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &\leq\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}}\right|+\left|{{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &\leq\left\|{\widetilde{\rho_{m_{k}}}-f}\right\|_{L^{\frac{4}{3}}}\cdot\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}+\left\|f\right\|_{L^{\frac{12}{11}}}\cdot\left\|{V_{f}-V_{\widetilde{\rho_{m_{k}}}}}\right\|_{L^{12}}\end{split}$
The last inequality above comes from the Hölder’s inequality. Thanks to Hardy-Littlewood-Sobolev Inequality Proposition A.3, we have $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$, and $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$. in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies $\left.V_{\widetilde{\rho_{m_{k}}}}\rightarrow V_{f}(x)\right$. in $L^{12}\left(\mathbb{R}^{3}\right)$. Thanks to the Interpolation Inequality, we have $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}\leq\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{1}}\cdot\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{4}{3}}}$, $\left\|f\right\|_{L^{\frac{12}{11}}}\leq\left\|f\right\|_{L^{1}}\cdot\left\|f\right\|_{L^{\frac{4}{3}}}$. Thus $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$ and $\left\|f\right\|_{L^{\frac{12}{11}}}$ are bounded, which—together with the inequality $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$—implies the boundedness of $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}$. Then
$\left.\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\rightarrow 0\right$.
Thus, we know the claim above holds that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}=e_{0}=E_{0}(\sigma)$.
In fact, the convergence up to a subsequence can be strengthened to convergence of the whole sequence. Since the non-rotating minimizer is unique up to translation due to Theorem 3.6, we get $\left\{\widetilde{\rho_{m_{k}}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ up to translations. By the same arguments, we know for any subsequence $\left\{\widetilde{\rho_{m_{k}}}\right\}$, we can always pick a subsequence $\left\{\widetilde{\rho_{m_{k_{l}}}}\right\}$ of $\left\{\widetilde{\rho_{m_{k}}}\right\}$ which converges strongly to $\sigma$ up to translations. Thus, by an argument of contradiction, we get, up to translations, $\left\{\widetilde{\rho_{m}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$. ∎
Actually, we can also show the convergence of $\{\widetilde{\rho_{1-m}}\}$ and $\{{\rho_{1-m}}\}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$.
Corollary 5.10 (Convergence of Scaling Densities of Stars in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$).
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{1-m}}=A(1-m)\rho_{1-m}\left({B(1-m)x}\right)$ as in Definition 5.1, we have, up to translations, $\{\widetilde{\rho_{1-m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
Proof.
We can apply the same arguments as above and obtain:
$E_{0}\left(\widetilde{\rho_{1-m}}\right)<C(1-m)^{\frac{6-5\gamma}{3\gamma-4}}\cdot m^{3}+E_{0}(\sigma)$
Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow the similar proof as in Theorem 5.9 above to get the result. ∎
Corollary 5.11 (Convergence of Densities of Stars in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$).
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, we have, up to translations, $\{\rho_{1-m}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
Proof.
Similarly, we can still compare $E_{0}(\sigma_{1-m})$ and $E_{0}(\rho_{1-m})$ without scaling the densities and obtain:
$E_{0}\left(\rho_{1-m}\right)<Cm^{3}+(1-m)^{\frac{5\gamma-6}{3\gamma-4}}E_{0}(\sigma)$
Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left({\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow the similar proof as in Theorem 5.9 above to get the result. ∎
Remark 5.12.
For $\rho_{m}$, we can similarly obtain:
$E_{0}\left(\rho_{m}\right)<Cm^{3}+m^{\frac{5\gamma-6}{3\gamma-4}}\cdot E_{0}(\sigma)$
Then we have ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\rho_{m}\right)}}=0$, which violates the condition of concentration-compactness lemma [42, Theorem II.2 and Corollary II.1]. In this case, we can actually show that $\rho_{m}$ converges to 0: from the scaling relation (36) and the fact $\lim\limits_{m\rightarrow 0}{A(m)}$$\infty$ and ${\lim\limits_{m\rightarrow 0}{B(m)}}=0$, we can show $\left.\left\|\rho_{m}\right\|_{L^{\frac{4}{3}}}=\left({A(m)}\right)^{-\frac{4}{3}}\left({B(m)}\right)^{3}\left\|\widetilde{\rho_{m}}\right\|_{L^{\frac{4}{3}}}\rightarrow 0\right$. as $m\rightarrow 0$. However, this vanishing limit result is not enough. We seek a stronger estimate that will allow us to estimate the Lagrange multiplier and then the support size of $\rho_{m}$ using the arguments in the following. This is the reason we introduce the scaling density.
Remark 5.13.
Note the convergence holds true for the whole sequence instead of up to subsequence.
5.2 Uniform bound of scaling densities in $L^{\infty}\left(\mathbb{R}^{3}\right)$
Thanks to the convergence results in subsection 5.1, now we can estimate the uniform $L^{\infty}$ bound of scaling densities $\left\{\widetilde{\rho_{m}}\right\}$ when $m$ is sufficiently small. Since $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ (Theorem 5.9), thus when $m$ is sufficiently small, $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$. The next steps are the following: (1) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{12}\left(\mathbb{R}^{3}\right)$; (2) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{6}\left(\mathbb{R}^{3}\right)$; (3) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$; (4) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$. Since we consider $\left\{\widetilde{\rho_{m}}\right\}$ and $\left\{V_{\widetilde{\rho_{m}}}\right\}$ alternatively, we call such method the bootstrap method.
Theorem 5.14 (Uniform bound of Scaling Densities of Planets and Potentials in $L^{\infty}\left(\mathbb{R}^{3}\right)$).
Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14. Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A(m)\rho_{m}\left({B(m)x}\right)$ as in Definition 5.1, then $\exists C>0$, $\delta>0$, $\forall 0<m<\delta$, we have $\widetilde{\rho_{m}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, $V_{\widetilde{\rho_{m}}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, where $V_{\widetilde{\rho_{m}}}(x)={\int{\frac{\widetilde{\rho_{m}}(y)}{\left|{x-y}\right|}d^{3}y}}$ is the potential from $\widetilde{\rho_{m}}$. Moreover, $V_{\widetilde{\rho_{m}}}(x)$ is continuous.
Proof.
[…]
$\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}=\left({\int{{\widetilde{\rho_{m}}}^{12{({\gamma-1})}}(x)d^{3}x}}\right)^{\frac{1}{12{({\gamma-1})}}}\leq{\widetilde{C}\left({\int{V_{\widetilde{\rho_{m}}}^{12}(x)d^{3}x}}\right)}^{\frac{1}{12{({\gamma-1})}}}=\widetilde{C}\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C_{2}$
By assumption $\gamma>\frac{3}{2}>\frac{4}{3}$, together with Interpolation Inequality Proposition A.2 we have $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}}\leq\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}^{\alpha}\cdot\left\|\widetilde{\rho_{m}}\right\|_{L^{1}}^{1-\alpha}\leq C$ for some $C>0$. Since we already know $\left\|\widetilde{\rho_{m}}\right\|_{L^{1}{(\mathbb{R}^{3})}}=1$ and $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}{(\mathbb{R}^{3})}}$ are bounded uniformly, thanks to Proposition A.4, we have $V_{\widetilde{\rho_{m}}}$ is continuous and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}{(\mathbb{R}^{3})}}$ is bounded uniformly. Again by ${\widetilde{\rho_{m}}}^{\gamma-1}\leq CV_{\widetilde{\rho_{m}}}$ we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is also bounded uniformly. ∎
Remark 5.15.
The uniform bound in $L^{\infty}(\mathbb{R}^{3})$ can be comparable with Li’s result [38, Proposition 1.4], where the rotating stars (corresponding to single-star case) are considered and the bound $C(m)$ however depends on $m$. Although one might conjecture $C(m)$ is continuous when $m>0$, they may still need to consider if $\lim\limits_{m\rightarrow 0}{\sup{C(m)}}$$\infty$ in order to obtain a result analogous to Theorem 5.14.
Remark 5.16.
By the similar arguments, we can obtain that when $m$ is sufficient small, $\left\{\widetilde{\rho_{1-m}}\right\}$ and $\left\{\rho_{1-m}\right\}$ are also bounded uniformly in $L^{\infty}(\mathbb{R}^{3})$.
Remark 5.17.
Similarly as Remark 3.12, once we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly, by definition we have $\rho_{m}=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)$, $A=m^{-\frac{2}{3\gamma-4}}$, thus ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$.
5.3 Estimate for the supports of densities
In this subsection, we finish the estimate for the supports of scaling densities as well as the original densities. We first seek an upper bound for the Lagrange multiplier. To avoid confusion with the Lagrange multiplier $\lambda_{m}$ corresponding to star planet system, we will use $\kappa_{m}$ to represent the corresponding non-rotating Lagrange multiplier with mass $m$ in this subsection.
Since both the function $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ in (2) and the scaling coefficient $A$ in (36) will appear frequently in this subsection, to avoid confusion, we will use $\mathcal{A}(\rho)$ to denote the function $A(\rho)$, while keeping $A$ to represent the coefficient $A$ in this subsection.
Similar to the arguments in the proof of Theorem 4.5 (see also [4, Lemma 2] [15, Proposition 5.8] [16, Lemma 2.21]), it can be shown that given a constrained minimizer $\rho(m)$, then $\rho(m)$ satisfies Euler-Lagrange equation in the following sense:
$\mathcal{A}^{\prime}(\rho(m)(x))=\left[\frac{J^{2}}{2I^{2}(\rho(m))}r^{2}(x-\bar{x}(\rho(m)))+V_{\rho(m)}(x)+\lambda_{\alpha}\right]_{+}\text{on }\Omega_{\alpha}$ (ELm)
where $\alpha\in\{m,1-m\}$.
Similar to Proposition 4.8 we can show $\lambda_{\alpha}<0$ for $m$ small enough. Now we give a stronger result in the following lemma.
Lemma 5.18 (Bound for the Lagrange Multiplier).
Fix $J>0,\forall 0<\epsilon<\kappa_{1},\exists\delta>0,\forall 0<m<\delta$, if $\rho(m)=\rho_{m}+\rho_{1-m}$ minimizes $E_{J}(\rho)$ on $W_{m}$, then the Lagrange multiplier $\lambda_{m}$ in the Euler-Lagrange equations (ELm) satisfies $A^{\gamma-1}(m)\cdot\lambda_{m}\leq\kappa_{1}+\epsilon<0$. Here $\kappa_{1}$ denotes the non-rotating Lagrange multiplier from Theorem 3.6 with respect to mass $1$.
Proof of Lemma 5.18.
Let $m$ be smaller than $m_{0}$ given in Theorem 4.14, then minimizers exist. From Euler-Lagrange equations (ELm) we have
$\lambda_{m}\leq\mathcal{A}^{\prime}\left(\rho_{m}\right)-V_{\rho_{m}}\text{ a.e. on }\Omega_{m}$
where $\mathcal{A}(\rho)(x)=\frac{K\left(\rho^{\gamma}(x)\right)}{\gamma-1},\mathcal{A}^{\prime}(\rho)(x)=\frac{K\gamma\left(\rho^{\gamma-1}(x)\right)}{\gamma-1}$. Due to (30) (31), we have $\mathcal{A}^{\prime}\left(\rho_{m}\right)(x)=\mathcal{A}^{\prime}\left(\frac{1}{A}\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)$, $V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$. Therefore, we have
$A^{\gamma-1}\lambda_{m}\leq\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)-V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)\text{ a.e. on }\Omega_{m}$
Then $\mu\left(\left\{x\in\Omega_{m}\mid A^{\gamma-1}\lambda_{m}>\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)-V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)\right\}\right)=0$
[…]
$\frac{\left|\mathcal{A}^{\prime}(\widetilde{\rho_{m}})-\mathcal{A}^{\prime}(\sigma)\right|}{\left|\widetilde{\rho_{m}}-\sigma\right|}=\frac{K_{\gamma}}{\gamma-1}\cdot\frac{\left|\widetilde{\rho_{m}}^{\gamma-1}-\sigma^{\gamma-1}\right|}{\left|\widetilde{\rho_{m}}-\sigma\right|}\leq K\gamma\xi^{\gamma-2}\leq K\gamma C^{\gamma-2}$
Thus $\left|\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)-\mathcal{A}^{\prime}(\sigma)\right|\leq C_{1}\left|\widetilde{\rho_{m}}-\sigma\right|$ with $C_{1}=K\gamma C^{\gamma-2}>0$. If $\left|\widetilde{\rho_{m}}(x)-\sigma(x)\right|=0$, then $\widetilde{\rho_{m}}(x)=\sigma(x)$, thus $\left|\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)-\mathcal{A}^{\prime}(\sigma)\right|\leq C_{1}\left|\widetilde{\rho_{m}}-\sigma\right|$ also holds true. Since $\widetilde{\rho_{m}}$ converges to $\sigma$ in measure, we know when $m\rightarrow 0$,
$\mu\left(\left\{x\in\mathbb{R}^{3}\mid\left|\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)-\mathcal{A}^{\prime}(\sigma)\right|(x)\geq\epsilon\right\}\right)\leq\mu\left(\left\{x\in\mathbb{R}^{3}\mid\left|\widetilde{\rho_{m}}(x)-\sigma(x)\right|\geq\frac{\epsilon}{c_{1}}\right\}\right)\rightarrow 0.$
[…]
By Theorem 3.6 (v) and the fact $\sigma$ has mass 1, we know $0<\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma(x)>0\right\}\right)<\infty$. Since $0\leq\sigma<C$ a.e. on $\mathbb{R}^{3}$, let $E_{n}=\left\{x\in\mathbb{R}^{3}\mid\frac{C}{n}\leq\sigma(x)<C\right\}$, then $\mathbb{R}^{3}=\cup_{n=1}^{\infty}E_{n}$. And then $1=\int_{\mathbb{R}^{3}}\sigma=\sum_{n=1}^{\infty}\int_{E_{n}}\sigma$. Then $\exists N>0$, such that $\frac{1}{2}<\sum_{n=1}^{N}\int_{E_{n}}\sigma<\sum_{n=1}^{N}\int_{E_{n}}C=C\mu(\{\{x\in\mathbb{R}^{3}\mid\frac{C}{N}\leq\sigma(x)<C\})$. Therefore, let $M:=\mu\left(\left\{x\in\mathbb{R}^{3}\mid\frac{C}{N}<\sigma(x)<C\right\}\right)$, then $0<\frac{1}{2C}<M<\infty$. Thus $\exists\delta>0,\forall 0<m<\delta$, due to the claim we have
$\mu\left(\left\{x\in\mathbb{R}^{3}\mid\left|\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{x}{B}\right)-\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)\right|\geq\frac{\epsilon}{3}\right\}\right)$
$=B^{3}\mu\left(\left\{y\in\mathbb{R}^{3}\mid\left|\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)(y)-\mathcal{A}^{\prime}(\sigma)(y)\right|\geq\frac{\epsilon}{3}\right\}\right)$
$<B^{3}\frac{M}{4}$
and similarly,
$\mu\left(\left\{x\in\mathbb{R}^{3}\mid\left|V_{\widetilde{\rho_{m}}}\left(\frac{x}{B}\right)-V_{\sigma}\left(\frac{x}{B}\right)\right|\geq\frac{\epsilon}{3}\right\}\right)<B^{3}\frac{M}{4}$
$\mu\left(\left\{x\in\mathbb{R}^{3}\mid\left|\widetilde{\rho_{m}}\left(\frac{x}{B}\right)-\sigma\left(\frac{x}{B}\right)\right|\geq\frac{C}{3N}\right\}\right)<B^{3}\frac{M}{4}$
Let $E=\left\{x\in\mathbb{R}^{3}\mid\frac{C}{N}<\sigma\left(\frac{x}{B}\right)<C\right\}$, $\mu(E)=B^{3}M$. Let
$E_{0}$ $=\left\{x\in\mathbb{R}^{3}\mid\left|\mathcal{A}^{\prime}(\widetilde{\rho_{m}})\left(\frac{x}{B}\right)-\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)\right|\geq\frac{\epsilon}{3}\right\}$
$\quad\cup\left\{x\in\mathbb{R}^{3}\mid\left|V_{\widetilde{\rho_{m}}}\left(\frac{x}{B}\right)-V_{\sigma}\left(\frac{x}{B}\right)\right|\geq\frac{\epsilon}{3}\right\}$
$\quad\cup\left\{x\in\mathbb{R}^{3}\mid\left|\widetilde{\rho_{m}}\left(\frac{x}{B}\right)-\sigma\left(\frac{x}{B}\right)\right|\geq\frac{C}{3N}\right\}$
$\quad\cup\left\{x\in\Omega_{m}\mid A^{\gamma-1}\lambda_{m}>\mathcal{A}^{\prime}(\widetilde{\rho_{m}})\left(\frac{1}{B}x\right)-V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)\right\}$
[…]
$A^{\gamma-1}\lambda_{m}$ $\leq\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)-V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$ (41)
$=\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)-\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)+\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)-V_{\sigma}\left(\frac{x}{B}\right)+V_{\sigma}\left(\frac{x}{B}\right)-V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$
$\leq\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)-V_{\sigma}\left(\frac{x}{B}\right)+\frac{2\epsilon}{3}$
$<\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)-V_{\sigma}\left(\frac{x}{B}\right)+\epsilon$
$=\kappa_{1}+\epsilon$
The last inequality comes from the facts that $x\in E$ implies $\frac{x}{B}\in\{\sigma>0\}$ and then $\mathcal{A}^{\prime}(\sigma)(\frac{x}{B})-V_{\sigma}(\frac{x}{B})=\kappa_{1}$ by Theorem 3.6 (vii).
Remark 5.19.
Note that the above proof relies on convergence in measure. In McCann’s paper [46, Lemma 6.5], only almost everywhere convergence up to subsequence is used. Nonetheless, since in this paper we assume $m$ is very small, relying solely on almost everywhere convergence up to subsequence will lead to the following two issues. One is that it is up to a subsequence, i.e. $\widetilde{\rho_{m_{k}}}(x)\rightarrow\sigma(x)$ a.e. on $\mathbb{R}^{3}$ instead of $\left\{\widetilde{\rho_{m}}\right\}$ itself, while we hope the bound for the Lagrange multiplier should be valid for all sufficient small mass. Another one is when $m\rightarrow 0$, not only $\widetilde{\rho_{m}}(x)\rightarrow\sigma(x)$ a.e. on $\mathbb{R}^{3}$, but the measures of relevant sets above go to 0. For example, when $m\rightarrow 0$, $\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma\left(\frac{x}{B(m)}\right)>0\right\}\right)=B^{3}(m)\mu\left(\left\{x\in\mathbb{R}^{3}\mid\sigma(x)>0\right\}\right)\rightarrow 0$. Therefore, we need to rule out the case that $\widetilde{E}$ is empty; otherwise, it is not clear whether there exists an $x$ that satisfies the last inequality in (41). Therefore, we make more use of $L^{p}$ convergence—namely, that it implies convergence in measure, to solve the issues.
Before we finally estimate the sizes of supports of scaling densities, we supplement the proof of a proposition about the $L^{\infty}$ bound of potential mentioned in [46, Proposition 6.6], which can be viewed as a complement to [4, Proposition 5], in which the indices $\frac{2}{3}$ and $\frac{1}{3}$ can be approximated but cannot be reached.
Proposition 5.20 (Bound of Potential in $L^{\infty}\left(\mathbb{R}^{3}\right)$ [46, Section 6]).
$\exists k>0,\forall\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, one has $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$, where $V_{\rho}(x)=\int\frac{\rho(y)}{|x-y|}d^{3}y$ is the potential from $\rho$.
Proof.
$V_{\rho}(x)$ $=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}d^{3}y$
[…]
$\leq CR^{2}\|\rho\|_{L^{\infty}}+\frac{1}{R}\|\rho\|_{L^{1}}$
Consider function $f(R)=CR^{2}\|\rho\|_{L^{\infty}}+\frac{1}{R}\|\rho\|_{L^{1}},R\geq 0$, it has the maximal value at $R=\left(\frac{\|\rho\|_{L^{1}}}{2C\|\rho\|_{L^{\infty}}}\right)^{\frac{1}{3}}$ and $f\left(\left(\frac{\|\rho\|_{L^{1}}}{2C\|\rho\|_{L^{\infty}}}\right)^{\frac{1}{3}}\right)=k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$ for some $k>0$. Thus, we have
$\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$
Theorem 5.21 (Bound on the sizes of supports of scaling densities $\widetilde{\rho_{m}}$).
There exists a radius $R(J)$ and $\delta>0$ independent of $m$, $\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\widetilde{\rho_{m}}$ is contained in a ball of radius $R(J)$, where $\widetilde{\rho_{m}}$ is defined in Definition 5.1.
Proof.
Take $m$ and $\epsilon$ small enough so that constrained minimizers $\rho(m)$ exist by Theorem 4.14 and $A(m)^{\gamma-1}\cdot\lambda_{m}\leq\kappa_{1}+\epsilon<$ 0 by Lemma 5.18, and the velocity $v(x,m)$ satisfies $v(x,m)^{2}\leq Cm^{2}$ by Corollary 5.5. By the construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\operatorname{dist}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{\eta}{2}$, $\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$, thus $\exists C>0,\forall x\in\Omega_{m}$, $V_{\rho_{1-m}}(x)<Cm^{2}$. In the Euler-Lagrange equation (ELm), we know for almost every $x\in\Omega_{m}$,
$\mathcal{A}^{\prime}\left(\rho_{m}\right)(x)=\frac{K\gamma}{\gamma-1}\rho_{m}^{\gamma-1}(x)=\left[v(x,m)^{2}+V_{\rho_{m}}(x)+V_{\rho_{1-m}}(x)+\lambda_{m}\right]_{+}$
Since $\rho_{m}(x)=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{1}{B}x\right)$, without loss of generality we assume $\frac{K\gamma}{\gamma-1}=1$, then we have
${\widetilde{\rho_{m}}}^{\gamma-1}\left(\frac{x}{B}\right)$ $=A^{\gamma-1}\left[v(x,m)^{2}+\frac{1}{A^{\gamma-1}}V_{\widetilde{\rho_{m}}}\left(\frac{x}{B}\right)+V_{\rho_{1-m}}(x)+\lambda_{m}\right]_{+}$
$\leq\left[CA^{\gamma-1}m^{2}+V_{\widetilde{\rho_{m}}}\left(\frac{x}{B}\right)+\kappa_{1}+\epsilon\right]_{+}$
Recall $A=m^{-\frac{2}{3\gamma-4}}$, $A^{\gamma-1}m^{2}=m^{\frac{4\gamma-6}{3\gamma-4}}$. Thus when $m$ is small enough, for almost every $y\in\Omega_{m}^{B}:=\left\{y\in\mathbb{R}^{3}\mid{B}y\in\Omega_{m}\right\}$, we have ${\widetilde{\rho_{m}}}^{\gamma-1}(y)\leq\left[V_{\widetilde{\rho_{m}}}(y)+\frac{\kappa_{1}+\epsilon}{2}\right]_{+}$. Now, let $R_{0}$ from Theorem 3.6 (v) bound the support radius of non-rotating minimizer $\sigma$, and choose $\tilde{R}(J)\geq R_{0}$ large enough so that $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$. Using Theorem 5.9, we know up to translation, $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{1}\left(\mathbb{R}^{3}\right)$. In particular, $\forall\delta>0$, when $m$ is small enough, all but mass $\delta$ of $\widetilde{\rho_{m}}$ is forced into a ball of radius $R_{0}$, which is, without loss of generality, centered on 0. Then $\widetilde{\rho_{m}}=\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$, $V_{\widetilde{\rho_{m}}}=V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$. Given $y$ outside the larger ball of radius $\tilde{R}(J)$ centered on 0, $V_{\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}}(y)\leq-\frac{\kappa_{1}+\epsilon}{6}$ due to $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$. For $V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}(y)$, by Theorem 5.14 we can show $\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$ with mass $\delta$. And we pick a $\delta$ satisfying $k\delta^{\frac{2}{3}}\tilde{C}^{\frac{1}{3}}<-\frac{\kappa_{1}+\epsilon}{6}$, where $\tilde{C}$ is the uniform bound of $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}}$ in Theorem 5.14. Thus, when $m$ is small enough, by Proposition 5.20, we have (note $V_{\widetilde{\rho_{m}}}$ is continuous due to Theorem 5.14)
$V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\{|x|\geq R_{0}\}}(y)$ $\leq\bigl\|V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\{|x|\geq R_{0}\}}\bigr\|_{L^{\infty}}$
$\leq k\bigl\|\widetilde{\rho_{m}}\cdot\mathbf{1}_{\{|x|\geq R_{0}\}}\bigr\|_{L^{1}}^{2/3}\bigl\|\widetilde{\rho_{m}}\cdot\mathbf{1}_{\{|x|\geq R_{0}\}}\bigr\|_{L^{\infty}}^{1/3}$
$\leq k\delta^{2/3}\tilde{C}^{1/3}$
$<-\frac{\kappa_{1}+\epsilon}{6}.$
By construction of $\Omega_{m}$ we know $\widetilde{\rho_{m}}=0$ outside $\Omega_{m}^{B}$. On $\Omega_{m}^{B}$ but outside the ball $B_{\tilde{R}(J)}(0)$, we know $0\leq\widetilde{\rho_{m}}{}^{\gamma-1}(y)\leq\left[V_{\widetilde{\rho_{m}}}(y)+\frac{\kappa_{1}+\epsilon}{2}\right]_{+}\leq\left[\frac{\kappa_{1}+\epsilon}{6}\right]_{+}=0$ a.e. In Theorem 4.14 we know $\rho_{m}$ is continuous on $\Omega_{m}$, hence $\widetilde{\rho_{m}}(x)=A\rho_{m}(Bx)$ is continuous on $\Omega_{m}^{B}$, then we can replace “a.e.” by “everywhere”. Therefore, up to translation, spt $\widetilde{\rho_{m}}$ is contained in $B_{2\tilde{R}(J)}(0)$. ∎
Remark 5.22.
There are various ways to bound $\left\|V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}\right\|_{L^{\infty}}$. For example, instead of applying Proposition 5.20, one can apply [4, Proposition 5] directly.
Remark 5.23.
In Lemma 5.18 and Theorem 5.21 we consider the the polytropic law (1) $P(\rho)=K\rho^{\gamma}$. McCann actually considered more general form of pressure $P(\rho)$ [46, Lemma 6.5 and Theorem 6.6]. In [46], $\kappa_{1}$ is replaced by $e_{0}^{\prime}(1^{-})$, where $e_{0}^{\prime}(1^{-})$ is negative and is a upper bound of Lagrange multipliers corresponding to non-rotating minimizers with fixed mass $1$, even in cases where multiple non-rotating minimizers exist (see Theorem 3.6 and Remark 3.9). Nevertheless, since we are dealing with scaling density here, we focus on this polytropic law.
Corollary 5.24 (Bound on the sizes of supports of original densities $\rho_{m}$).
There exists a radius $R(J)$ and $\delta>0$ independent of $m,\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\rho_{m}$ is contained in a ball of radius $B\cdot R(J)$, where $B$ is defined in Definition 5.1. In particular, spt $\rho_{m}$ is contained in a ball of a radius which goes to 0 with rate $B$ since $\lim\limits_{m\rightarrow 0}B=0$ if $\gamma>2$.
Proof.
[…]
We have shown the convergence rates of the bound of the supports. In particular, when planets’ mass goes to zero, their density supports will be contained in some balls with same radius $R$. Same results can be shown for stars’ scaling densities and original densities.
Corollary 5.25 (Bound on the sizes of supports of scaling densities $\widetilde{\rho_{1-m}}$).
There exists a radius $R(J)$ and $\delta>0$ independent of $m,\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\widetilde{\rho_{1-m}}$ is contained in a ball of radius $R(J)$, where $\widetilde{\rho_{1-m}}$ is defined in Definition 5.1.
Proof.
The proof is essentially the same as the proof above. ∎
Corollary 5.26 (Bound on the sizes of supports of original densities $\rho_{1-m}$).
There exists a radius $R(J)$ and $\delta>0$ independent of $m,\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\rho_{1-m}$ is contained in a ball of radius $R(J)$.
Proof.
The proof is essentially the same as the proof above, even without scaling stars’ densities. ∎
6 Existence for Star-Planet Systems
So far, we have dealt with the uniform bound of the supports of density functions with different mass. Our next step is to show the centers of mass can be chosen to be not too far away from the centers of $\Omega_{1-m}$ and $\Omega_{m}$. Thanks to those results, we then can show the constrained minimizers are actually $W^{\infty}$ local energy minimizers, which represent solutions to the reduced Euler-Poisson system (EP’) up to translation due to Theorem 3.1.
6.1 Estimate for the Center of Mass Separation
In this subsection, we estimate for the centers of mass separation by modifying the arguments in [46, Lemma 6.7 and Proposition 6.8] such that they can be applied in our case.
Lemma 6.1.
For $\epsilon>0$, define $g_{\epsilon}(z)=-\frac{1}{z-2\epsilon}+\frac{1}{2\left(z^{2}+\epsilon^{\frac{3}{2}}R^{\frac{1}{2}}J^{-1}\right)}+\frac{1}{1+2\epsilon}-\frac{1}{2}$. For $\epsilon$ is small enough, $g_{\epsilon}$ converges uniformly to $g_{0}(z)=-\frac{1}{z}+\frac{1}{2z^{2}}+\frac{1}{2}$ as $\epsilon\rightarrow 0$ on $\left\{z\in\mathbb{C}\mid|z|\geq\frac{1}{2}\right\}$.
Proof.
One can observe when $\epsilon$ is small enough, the function $g_{\epsilon}$ and $g_{\epsilon}^{\prime}$ are analytic and uniformly bounded on $\left\{z\in\mathbb{C}\mid|z|>\frac{1}{4}\right\}$. The result can be check by direct computation and estimation. ∎
Proposition 6.2 (Estimate for the Center of Mass Separation).
Fix $J>0$, let $0<\zeta<\frac{1}{2},\exists\delta>0$, such that $\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then $1-\zeta<\frac{\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|}{\left|y_{m}-y_{1-m}\right|}<1+\zeta$, where $y_{m}$ and $y_{1-m}$ are the centers of $\Omega_{m}$ and $\Omega_{1-m},\left|y_{m}-y_{1-m}\right|=\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
Proof.
Take $m$ small enough such that Corollary 5.24 and Corollary 5.26 provide bound $R(J)$ for the supports of $\rho_{m}$ and $\rho_{1-m}$, and $R:=2R(J)<\frac{\eta}{4}$. Due to Corollary 5.24, we know spt $\rho_{m}\subset B_{R(J)}\left(z_{m}\right)$ for some $z_{m}\in\mathbb{R}^{3}$, then $\bar{x}\left(\rho_{m}\right)\in B_{R(J)}\left(z_{m}\right)$. If $x$ is outside $B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$, then $\left|x-z_{m}\right|\geq\left|x-\bar{x}\left(\rho_{m}\right)\right|-\mid z_{m}$- $\bar{x}\left(\rho_{m}\right)\mid>R(J),\rho_{m}(x)=0$. Therefore, spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$. Similarly, spt $\rho_{1-m}\subset$ $B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$. Now we translate $\rho_{m}$ and $\rho_{1-m}$ separately so that their centers of mass are positioned at $y_{m}$ and $y_{1-m}$ respectively, and denote the shifted functions as $\kappa_{m}$ and $\kappa_{1-m}$. Therefore, $\kappa(m)\coloneq\kappa_{m}+\kappa_{1-m}\in W_{m}$. Due to the proof of Lemma 6.3 below we know $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=c$. Hence the distance between them equals the distance between their projections onto $x_{1}x_{2}$ plane. That is, $d:=\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|=r\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)$, where $r$ is given in Definition 2.1. Since spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$, we have (recall $\mu=m(1-m)$)
$\frac{\mu}{d+2R}\leq G\left(\rho_{m},\rho_{1-m}\right)=\iint\frac{\rho_{m}(x)\cdot\rho_{1-m}(y)}{|x-y|}d^{3}xd^{3}y\leq\frac{\mu}{d-2R}$
By Lemma 2.4, we know
$I(\rho(m))=\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+\int r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{1-m}(x)d^{3}x+\int r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{m}(x)d^{3}x$
Hence
$\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)=\mu d^{2}\leq I(\rho(m))\leq\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+(1-m)R^{2}+mR^{2}=\mu d^{2}+R^{2}$
By definition of $T_{J}$ (16), we have
$\frac{J^{2}}{2\left(\mu d^{2}+R^{2}\right)}\leq T_{J}(\rho(m))=\frac{J^{2}}{2I(\rho(m))}\leq\frac{J^{2}}{2\mu d^{2}}$
Similar inequalities hold true for $G\left(\kappa_{m},\kappa_{1-m}\right)$ and $T_{J}(\kappa(m))$, but we replace $d$ by $\eta$, since $\left|\bar{x}\left(\kappa_{m}\right)-\bar{x}\left(\kappa_{1-m}\right)\right|=r\left(\bar{x}\left(\kappa_{1-m}\right)-\bar{x}\left(\kappa_{m}\right)\right)=\left|y_{m}-y_{1-m}\right|=\eta$. Note $\rho(m)$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, we have $E_{J}(\kappa(m))-E_{J}(\rho(m))=$ $G\left(\kappa_{m},\kappa_{1-m}\right)+T_{J}(\kappa(m))-G\left(\rho_{m},\rho_{1-m}\right)+T_{J}(\rho(m))\geq 0$, just as similar as (40). Using the previous estimates and $J^{2}=\mu^{2}\eta$, we have
$\frac{1}{d-2R}+\frac{\mu\eta}{2\left(\mu d^{2}+R^{2}\right)}\leq-\frac{1}{\eta+2R}+\frac{1}{2\eta}$
Let $x=\frac{d}{\eta}$, $\epsilon=\frac{R}{\eta}$, then
$g_{\epsilon}(x)=-\frac{1}{x-2\epsilon}+\frac{1}{2\left(x^{2}+\epsilon^{\frac{3}{2}}R^{\frac{1}{2}}J^{-1}\right)}+\frac{1}{1+2\epsilon}-\frac{1}{2}\leq 0$ (3.3.1)
$\rho(m)\in W_{m}$ implies $\bar{x}\left(\rho_{m}\right)\in\Omega_{m},\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m}$, thus $\frac{1}{2}\leq x\leq\frac{3}{2}$ by the construction of $\Omega_{m}$ and $\Omega_{1-m}$. Let $g_{0}(z)=-\frac{1}{z}+\frac{1}{2z^{2}}+\frac{1}{2}$, by Lemma 6.1 we know $g_{\epsilon}(z)$ converges uniformly to $g_{0}(z)$ as $\epsilon\rightarrow 0$ on $|z|\geq\frac{1}{2}$. One can see $g_{0}(1)=0$, and $\forall\zeta>0$, $\exists C>0$, such that on $\left[\frac{1}{2},\frac{3}{2}\right]\setminus[1-\zeta,1+\zeta]$, we have $g_{0}(x)>C$. Due to Lemma 6.1, we have $g_{\epsilon}(x)>\frac{C}{2}>0$ on $\left[\frac{1}{2},\frac{3}{2}\right]\setminus[1-\zeta,1+\zeta]$ for $\epsilon$ small enough (i.e. for $m$ small enough since $\epsilon=\frac{R}{\eta}=\frac{Rm^{2}(1-m)^{2}}{J^{2}}$). Thus (3.3.1) implies $x=\frac{d}{\eta}$ has to satisfies $1-\zeta<\frac{d}{\eta}<1+\zeta$, which proves the proposition. ∎
6.2 Existence theorem
After making some minor modifications to McCann’s proof in [46, Section 6], we proceed by translating $\rho$ such that it enjoys a plane of symmetry $z=0$, and then use the results above to show $\rho$ is supported away from the boundary of $\Omega_{1-m}\cup\Omega_{m}$, then we finalize the proof of existence.
Lemma 6.3 (Support Separation from the Boundary).
Fix $J>0,\exists\delta>0,\forall 0<m<\delta$, any constrained minimizer $\rho(m)=\rho_{m}+\rho_{1-m}$ of $E_{J}(\rho)$ on $W_{m}$ will, after a rotation with respect to the $z$-axis and a translation, have support contained in the interior of $\Omega_{1-m}\cup\Omega_{m}$, which means dist(spt $\left.\rho(m),\mathbb{R}^{3}\backslash\left(\Omega_{m}\cup\Omega_{1-m}\right)\right)>0$. Moreover, $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=0$, and $\rho(m)$ will also be symmetric about the plane $z=0$ and a decreasing function of $|z|$.
Proof.
Take $m$ small enough so that minimizers exist. First, we claim that any minimizer $\rho(m)=\rho_{m}+\rho_{1-m}$ for $E_{J}(\rho)$ on $W_{m}$ may be translated so that both $\bar{x}\left(\rho_{m}\right)$ and $\bar{x}\left(\rho_{1-m}\right)$ lie in the plane $z=0$. In fact, since the $\Omega_{m}$ and $\Omega_{1-m}$ are convex and symmetric about $z=0$, it is enough to know that $\rho(m)$ enjoys a plane of symmetry $z=c$. This follows from a strong rearrangement inequality in Lieb [41, Lemma 3] and Fubini’s Theorem: by the definition of $U(\rho)$ and $I(\rho)$, they will not be changed after the symmetric decreasing rearrangement of $\rho$ along lines parallel to the z-axis; however, since $\left(r^{2}+z^{2}\right)^{-\frac{1}{2}}$ is strictly decreasing as a function of $|z|$, the rearrangement increases $G(\rho,\rho)$ unless $\rho$ is already symmetric decreasing about a plane $z=c$. Note $\Omega_{1-m}$ and $\Omega_{m}$ are convex, hence $\rho(m)$’s rearrangement is in $W_{m}$. Since $\rho(m)$ minimizes $E_{J}(\rho)$, then $G(\rho,\rho)$ cannot be increased after rearrangement, which means $\rho(m)$ enjoys a plane of symmetry $z=c$.
Now, similar to the arguments in the proof of Proposition 6.2, we take $m$ small enough so that Corollary 5.24 and Corollary 5.26 provide a bound $R$ such that spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$ if $\rho(m)=\rho_{m}+\rho_{1-m}$ minimizes $E_{J}(\rho)$ on $W_{m}$. Translate $\rho(m)$ so that its symmetry plane is $z=0$ and let $d:=$ $\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|$. Then by construction of $\Omega_{m}$ and $\Omega_{1-m}$ we know that if $d-2R>\frac{\eta}{2}$ and $d+2R<\frac{3\eta}{2}$, a translation and rotation of $\rho(m)$ with respect to $z$-axis yields a minimizer in $W_{m}$ supported away from the boundary of $\Omega_{m}\cup\Omega_{1-m}$. By Proposition 6.2, this is certainly true when $m$ is sufficiently small. ∎
Now we finally come to the proofs of our main results Theorem 2.12 and Theorem 2.13
Theorem 2.12.
Let $\delta$ as in Lemma 6.3 and $\rho(m)$ be the constrained energy minimizer in Theorem 4.14, then Lemma 6.3 shows that spt $\rho({m})$ is compact in the interior of $\Omega_{m}\cup\Omega_{1-m}$, therefore separated from the boundary by a positive distance $\zeta$. Lemma 2.9 (ii) shows that if $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho(m),\kappa)<\eta$, then $\kappa$ in $W_{m}$. Thus $E_{J}(\rho(m))\leq E_{J}(\kappa)$, and $\rho(m)$ is a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$ (Existence of $W^{\infty}$ local energy minimizer). By definition of $W_{m}$, we know all the functions in $W_{m}$ have compact support. Thus after translation we can set $\rho(m)\in{R}_{0}\left(\mathbb{R}^{3}\right)$. Since $\rho(m)$ is a $W^{\infty}$ local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$, in particular on ${R}_{0}\left(\mathbb{R}^{3}\right)$, Theorem 3.1 provides a local minimizer $(\rho(m),v)$ of $E(\rho,v)$ subject to the constraint on $J_{z}=J$, with $v(x)=$ $\frac{J}{I(\rho(m))}\hat{e}_{z}\times x$. [46, Section 2] and [15, Remark 2.31] show that $(\rho(m),v)$ minimizes $E(\rho,v)$ subject to the constraint on the vector angular momentum as well (Part (i)). Part (ii) comes from Lemma 6.3 (Part (ii)). By Theorem 3.1 we know $\rho(m)$ is continuous and satisfies (EP’), and $(\rho(m),v)$ satisfies (EP) (Part (iii)). Parts (iv) and (v) come from Remark 5.16, Remark 5.17, Corollary 5.24 and Corollary 5.26 (Part (iv) and Part (v)). ∎
Theorem 2.13.
The proof of Theorem 2.13 is essentially the same as the proof of Theorem 2.12. Two things we need to note are (1) if we don’t know $\gamma>2$, we may not have the fact that there is a “uniform bound on the radius of $\rho(m)$’s support for all small mm” from Corollary 5.24. However, since the radii of $\Omega_{1-m}$ and $\Omega_{m}$ is $\frac{\eta}{4}\sim m^{-2}$, recall $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$, due to Corollary 5.24, we can still have the fact that $\rho(m)$ is located in the interior of $\Omega_{m}\cup\Omega_{1-m}$ if ${\frac{\gamma-2}{3\gamma-4}}>-2$, that is $\gamma>\frac{10}{7}$; (2) in order to obtain the energy convergence in Proposition 5.8, we hope the term $Cm^{\frac{4\gamma-6}{3\gamma-4}}$ in (39) satisfies $\frac{4\gamma-6}{3\gamma-4}>0$, which means $\gamma>\frac{3}{2}$. Therefore, when $\gamma>\max\{\frac{3}{2},\frac{10}{7}\}=\frac{3}{2}$, we have Theorem 2.13 holds true. ∎
7 Upper Bounds for Distances Between Connected Components of Minimizers
It should be noted that, although we prove the existence of Wasserstein $L^{\infty}$ local minimizer on $R(\mathbb{R}^{3}$), which is also a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ on $W_{m}$ (Theorem 2.12 and Theorem 2.13), in particular, spt $\rho(m)$ has at least two connected components, it is still unclear whether it has exactly two connected components. Neither we in this paper nor McCann [46, Theorem 6.1 and Corollary 6.2] elaborated on or provided proof for the result. If spt $\rho_{1-m}$ or spt $\rho_{m}$ have multiple connected components, the model may correspond to a system with multiple stars and planets, such as the relationship between the Solar System and the asteroid belt. Moreover, one could interpret it as describing a large and a small galaxy; therefore the existence results retain physical relevance and may inspire further study of the Vlasov–Poisson system via Rein’s reduction method (see e.g. [31, 49, 50]).
Although we have not proven simple connectedness in each subdomain, we can still make the following observations and conclusions that the distance between each two convex connected components of spt $\rho_{1-m}$ or spt $\rho_{m}$ should not be too large.
We assume spt $\rho_{m}$ has 2 convex connected components, said $U$ and $V$, and set dist $(U,V)=d$. By Lemma 6.3 we can assume $\rho(m)$ is symmetric about the plane $z=0$. We recall $\rho(m)$ is continuous in the whole space (Theorem 2.12 or Theorem 2.13). Also we notice that $E_{J}$ is invariant with respect to rotation around $z$-axis.Without loss of generality, we can assume there exists $b$, $b^{\prime}$ such that
$\sup\limits_{x\in U}x_{1}<b<\inf\limits_{y\in V}y_{1}\leq\sup\limits_{y\in V}y_{1}<b^{\prime}<\inf\limits_{z\in\operatorname{spt}\rho_{1-m}}z_{1}$ (42)
Then we know $d=\inf\limits_{y\in V}y_{1}-\sup\limits_{x\in U}x_{1}$. We denote the mass in $U$ is $m_{1}\coloneq\int_{U}\rho_{m}$, then the mass in $V$ is $m_{2}\coloneq\int_{V}\rho_{m}=m-m_{1}$. Notice $\rho_{m}(x)=\rho_{U}(x)+\rho_{V}(x)$, where $\rho_{U}(x)=\rho_{m}(x)\cdot\mathbf{1}_{U}(x)$, $\rho_{V}(x)=\rho_{m}(x)\cdot\mathbf{1}_{V}(x)$, $\mathbf{1}_{E}$ is the indicator function of set $E$. We want to move $\rho_{U}(x)$ and $\rho_{V}(x)$ in the following way: given $h>0$, let $h_{1}>0$ and $h_{2}>0$ satisfy $h_{1}+h_{2}=h$, $h_{1}\cdot m_{1}=h_{2}\cdot m_{2}$. With an abuse of notation we denote $h_{1}=h_{1}\cdot\hat{e_{1}}=(h_{1},0,0)^{T}$, similarly $h_{2}=(h_{2},0,0)^{T}$, $h=(h,0,0)^{T}$. Let $\rho_{m}^{[h]}=\rho_{U}(x-h_{1})+\rho_{V}(x+h_{2})$. Easy to see when $h$ is small enough, the connected components of spt $\rho_{m}^{[h]}$ are $U_{h}=U+h_{1}$, and $V_{h}=V-h_{2}$, and we still have $\sup\limits_{x\in U_{h}}x_{1}<b<\inf\limits_{y\in V_{h}}y_{1}$. We can write $\rho_{m}^{[h]}=\rho_{U}^{[h]}+\rho_{V}^{[h]}$ where $\rho_{U}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{U_{h}}$ and $\rho_{V}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{V_{h}}$. We notice the centers of mass of $\rho_{m}$ and $\rho_{m}^{[h]}$ are the same. Let $\rho^{[h]}(m)=\rho_{m}^{[h]}+\rho_{1-m}$, then we will show if $d$ is too large, then $E_{J}(\rho^{[h]}(m))<E_{J}(\rho(m))$.
In fact,
$E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)=T_{J}\left({\rho(m)}\right)-T_{J}\left({\rho^{[h]}(m)}\right)-\frac{1}{2}\left({G\left({\rho(m),\rho(m)}\right)-G\left({\rho^{[h]}(m),\rho^{[h]}(m)}\right)}\right)$ (43)
We apply Lemma 2.4 (Expansion of Moment of Inertia) twice and obtain
$\begin{split}I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)=I\left(\rho_{m}^{[h]}\right)-I\left(\rho_{m}\right)=\frac{m_{1}m_{2}}{m_{1}+m_{2}}r^{2}\left({\bar{x}\left(\rho_{U}^{[h]}\right)-\bar{x}\left(\rho_{V}^{[h]}\right)}\right)-\frac{m_{1}m_{2}}{m_{1}+m_{2}}r^{2}\left({\bar{x}\left(\rho_{U}\right)-\bar{x}\left(\rho_{V}\right)}\right)\end{split}$
Notice ${\bar{x}}\left(\rho_{V}^{[h]}\right)-{\bar{x}}\left(\rho_{U}^{[h]}\right)={\bar{x}}\left(\rho_{V}\right)-{\bar{x}}\left(\rho_{U}\right)-h$, thus we have
$\begin{split}I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)=\frac{m_{1}m_{2}}{m}\left({-2h\left({{\bar{x}}_{1}\left(\rho_{V}\right)-{\bar{x}}_{1}\left(\rho_{U}\right)}\right)+h^{2}}\right)=\frac{1}{m}{\int{-2h\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)d^{3}xd^{3}y}}+\frac{m_{1}m_{2}}{m}h^{2}\end{split}$
Therefore, we have
$\begin{split}&\quad T_{J}\left({\rho(m)}\right)-T_{J}\left({\rho^{[h]}(m)}\right)\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)}-\frac{J^{2}}{2I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}\left({I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)}\right)}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}\cdot{\left({{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{-2h\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)d^{3}xd^{3}y}}+m_{1}m_{2}h^{2}}\right)}\end{split}$ (44)
On the other hand,
$G\left({\rho(m),\rho(m)}\right)-G\left({\rho^{[h]}(m),\rho^{[h]}(m)}\right)=G\left({\rho_{m},\rho_{m}}\right)-G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)+2G\left(\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}\right)$ (45)
Let $h>0$ small enough, such that if $\rho_{U}(x)>0$ and $\rho_{V}(y)>0$ then $y_{1}-x_{1}-h>0$, by Taylor expansion we have $\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h}\right|}=\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)$, therefore,
$\begin{split}G\left({\rho_{m},\rho_{m}}\right)-G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)&=G\left({\rho_{U}+\rho_{V},\rho_{U}+\rho_{V}}\right)-G\left({\rho_{U}^{[h]}+\rho_{V}^{[h]},\rho_{U}^{[h]}+\rho_{V}^{[h]}}\right)\\ &=2G\left({\rho_{U},\rho_{V}}\right)-2G\left({\rho_{U}^{[h]},\rho_{V}^{[h]}}\right)\\ &=2{\int{\frac{\rho_{U}(x)\rho_{V}(y)-\rho_{U}\left({x-h_{1}}\right)\rho_{V}\left({y+h_{2}}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &=2{\int{\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y}\right|}-\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y+h}\right|}d^{3}xd^{3}y}}\\ &=2{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)}\right)d^{3}xd^{3}y}}\\ &=2{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}d^{3}xd^{3}y}}+2m_{1}m_{2}O\left(h^{2}\right)\\ &\leq-\frac{2}{R^{3}(m)}{\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}+2m_{1}m_{2}O\left(h^{2}\right)\end{split}$ (46)
where $R(m)=B(m)\cdot R(J)$ given in Corollary 5.24. Similarly, we can show when $h$ is small enough,
$\begin{split}&G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\\ &={\int{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)+\rho_{V}(y)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}-{\int{\frac{\rho_{1-m}(x)\left({\rho_{U}\left({y-h_{1}}\right)+\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &={\int{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)-\rho_{U}\left({y-h_{1}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}+{\int{\frac{\rho_{1-m}(x)\left({\rho_{V}(y)-\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &={\int{\rho_{1-m}(x)\rho_{U}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y-h_{1}}\right|}}\right)d^{3}xd^{3}y}}+{\int{\rho_{1-m}(x)\rho_{V}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h_{2}}\right|}}\right)d^{3}xd^{3}y}}\\ &\leq{\int{\rho_{1-m}(x)\rho_{U}(y)\left(\frac{h_{1}}{\left|{x-y}\right|\left|{x-y-h_{1}}\right|}\right)d^{3}xd^{3}y}}+{\int{\rho_{1-m}(x)\rho_{V}(y)\left(\frac{h_{2}}{\left|{x-y}\right|\left|{x-y+h_{2}}\right|}\right)d^{3}xd^{3}y}}\end{split}$
By the construction of $\Omega_{1-m}$ and $\Omega_{m}$ (23) we know $\rho_{1-m}(x)>0$, $\rho_{m}(y)>0$ implies $\frac{J^{2}}{2m^{2}\left({1-m}\right)^{2}}\leq\left|{x-y}\right|\leq\frac{3J^{2}}{2m^{2}\left({1-m}\right)^{2}}$, and when $h$ is small enough, $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{1}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$ and $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{2}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$. We also notice that by construction ${\int{\rho_{U}h_{1}}}={\int{\rho_{V}h_{2}}}=\frac{m_{1}m_{2}h}{m_{1}+m_{2}}=\frac{m_{1}m_{2}h}{m}$, then
$G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\leq\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h$
Similarly, we have
$G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)=G\left({\rho_{1-m},\rho_{m}^{[h]}-\rho_{m}}\right)\leq\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h$
Thus,
$\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|\leq\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h$ (47)
Similar to the arguments in Lemma 5.3, we know $\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}\leq Cm^{5}$ when $m$ is small enough. Therefore, similar to Corollary 5.24 and due to Definition 5.1 which gives coefficient $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$, we collect (43) (44) (45) (46) (47) and have (recall we assume $\gamma>\frac{3}{2}$)
$\begin{split}&E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}(m)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}}\right)\cdot{\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}(m)}-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$ (48)
Therefore, when $d>\frac{32m^{3}\left({1-m}\right)^{5}R^{3}(m)}{J^{4}}=\frac{32\left({1-m}\right)^{5}R^{3}(J)}{J^{4}}m^{\frac{12\gamma-18}{3\gamma-4}}$, we know $E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$ for $h$ small enough. That is, if $\rho(m)$ is a constrained minimizer, then $d\leq Cm^{\frac{12\gamma-18}{3\gamma-4}}$ for some $C$ and all small $m$. More precisely, we have the following proposition:
Proposition 7.1.
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13. There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{m}$ has 2 convex connected components, then the distance between these 2 components is less then $\eta$, where $\eta=Cm^{\frac{12\gamma-18}{3\gamma-4}}$.
Proof.
See the arguments above. ∎
Remark 7.2.
We put the condition $\gamma>3/2$ so that we can apply Theorem 2.13, but in fact the estimate (48) actually holds even for smaller values of $\gamma$.
Actually, similar result can be shown for stars’ densities.
Corollary 7.3.
Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13. There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{1-m}$ has 2 convex connected components, then the distance between these 2 components is less then $\eta$, where $\eta=Cm^{5}$.
Proof.
The proof is essentially the same as the proof above. Suppose spt $\rho_{1-m}$ has two convex connected components $U$ and $V$, after moving $U$ and $V$ similarly, we have
$\begin{split}&\quad E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}\left({1-m}\right)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}}\right){\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}\left({1-m}\right)}-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$
We just note that for star density, R(1m)R(1-m) will bounded by $R(J)$ as $m$ goes to 0 ($R(J)$ is given in Corollary 5.26). Thus we know $\frac{1}{R^{3}\left({1-m}\right)}\geq\frac{2J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}$ for $m$ small enough. Moreover, when $d>\frac{32m^{5}\left({1-m}\right)^{3}R^{3}(1-m)}{J^{4}}$, we know $E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$ for $h$ small enough. Thus we can apply the method of contradiction to prove this corollary. ∎
Remark 7.4.
We have estimated distance between (possible) two connected components. For cases with more than two connected components, say $U_{1},U_{1},U_{3},$... (could be uncountable many), if they satisfy certain shape configurations — for example $U_{i}$ and $V_{i}$ satisfy (42) up to rotation, where $V_{i}={\bigcup\limits_{\alpha\neq i}U_{\alpha}}$ — we can then move $U_{i}$ and $V_{i}$ similarly as above and analyze the distance relationships between them. Furthermore, we can also use Theorem 2.13 (ii) to rule out some configurations of connected components. For example, the case where an outer shell encloses an inner solid sphere is impossible for $\rho_{m}$ or $\rho_{1-m}$.
Remark 7.5.
It should be noted that the troublesome term in (48) preventing us from further reducing the distance between connected components $U$ and $V$ is the term $\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}$, which comes from (47). In other words, we might be able to prove simple connectedness in each subdomain by moving and gluing the relevant connected components, if we can show that $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)$ can absorb $\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$. For instance, this could be achieved by proving $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)>10\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$ for all small $m$ (that is $m\in(0,\delta)$ for some $\delta>0$) and for all positive distance $d$ between two connected components $U$ and $V$.
Although we have not proved here that the simple connectedness in each subdomain when $J>0$, we recall when $J=0$ (non-rotating case), the support of any minimizer $\sigma_{m}$ in Theorem 3.6 is truly simply connected (see Remark 3.7). This leads us to the following conjecture:
Conjecture 7.6.
The support of $\rho(m)$ in Theorem 2.13 should have exactly 2 connected components when $m$ small enough.
Remark 7.7.
Results on simple connectedness obtained in other settings may also be consulted; see, for example, [30, 53, 54].
Appendix A Appendix
Here we introduce some preliminary knowledge which are used in the paper.
A.1 Calculus of Variations in Vector Spaces
Since the functional we deal with is not differentiable in any usual sense, we begin with a less restrictive notion of differentiability.
Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$. Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)>0,\text{ such that }u+tv\in U\text{ for }t\in[0,\epsilon(v))\}$. This set is a cone in $X$. If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$. We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $\mu$ on $X(P)$ such that $\forall v\in P$
$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=\mu(v),$
The linearity of $\mu$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $\mu$ by $E_{u}^{\prime}$—with $P$ understood from context.
We say $E$ has a weak local minimum at $u\in U$ if for each $v\in P(u)$, there exists $\{t_{n}\}$ such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$ and $E\left({u+t_{n}v}\right)\geq E(u)$ for $n$ large enough. And $u$ is called weak local minimizer. We discuss more about the motivation of weak local minimizer in [15].
Let $W\subset\left\{u\in U\middle|g(u)=M\right\}$, where $g$ is a linear functional on $X$ and $M$ is a constant. Then similarly, we say $E$ has a weak local minimum at $u$ restricted to $W$ if $u\in W$ and for each $v\in P(u)$, there exists $t_{n}$} such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$, and for for $n$ large enough, we have $u+t_{n}v\in W$, and $E\left({u+t_{n}w}\right)\geq E(u)$. Furthermore, we have a result for weak local minimum restricted to $W$:
Theorem A.1 ((Generalized) Lagrange Multiplier Theorem [4, Section 2] [15, Appendix A]).
Suppose $E$ is $P$-differentiable at $u\in U$ and that its restriction to $W$ has a weak local minimum at $u$. Suppose also that $P$ is convex and that there is a $u_{0}\in P$ such that $u_{0}\in P$ and $g\left(u_{0}\right)\neq 0$. Then there is a unique constant $\lambda$ such that $E_{u}^{\prime}(v)\geq\lambda g(v)$, all $v\in P$.
A.2 Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this paper.
Proposition A.2 (Interpolation Inequality [8, Section 4.2]).
If $f\in L^{p}\cap L^{q}$ with $1\leq p\leq q\leq\infty$, then $f\in L^{r}$ for all $r,p\leq r\leq q$, and the following interpolation inequality holds:
$\|f\|_{r}\leq\|f\|_{p}^{\alpha}\|f\|_{q}^{1-\alpha},\text{ where }\frac{1}{r}=\frac{\alpha}{p}+\frac{1-\alpha}{q},0\leq\alpha\leq 1$
Proposition A.3 (Hardy-Littlewood-Sobolev Inequality [5, Theorem 1.7]).
Let $1<p,r<\infty$ and $0<\alpha<n$ be such that $\frac{1}{p}+\frac{\alpha}{n}=\frac{1}{r}+1$. $\exists C_{p,\alpha,n}>0$, such that
$\left\||\cdot|^{-\alpha}*f\right\|_{L^{r}\left(\mathbb{R}^{n}\right)}\leq C_{p,\alpha,n}\|f\|_{L^{p}\left(\mathbb{R}^{n}\right)}$ (49)
It turns out for potential $V_{\rho}$ given in (9), we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int\frac{y_{j}\rho(x-y)}{|y|^{3}}d^{3}y$ not only in the distribution sense but also in the classical sense:
Proposition A.4 (Differentiability of Potential [4, Proposition 7] [15, Proposition A.13]).
If $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$ for some $p>3$, then $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)$ is continuously differentiable and the weak derivative coincides with the classical one for all $x\in\mathbb{R}^{3}$.
To prove the low semicontinuity of the energy, we hope to connect the weak convergence of densities and strong convergence of potentials.
Proposition A.5 (Compactness of Convolution (Potential) Operator with $q\geq 1$ [16, Proposition A.7]).
Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{3}\right)$. Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{3}\right)$ with $q\geq 1$. Then let $\Omega\subset\mathbb{R}^{3}$ is open, bounded and of class $C^{1}$.
  1. (1)
    If $1\leq q<\frac{3}{2}$, then $\forall 1\leq r<p$, where $\frac{1}{p}=\frac{1}{q}-\frac{2}{3},\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
  2. (2)
    If $q=\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\forall 1\leq r<\infty,\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
  3. (3)
    If $q>\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$.
Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez (master’s thesis advisor) and Dimitri Cobb (second advisor). Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino, Han Cao for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.

### v2

@@
 Abstract
-This paper investigates the existence and properties of stable, uniformly rotating star-planet systems, i.e.
-mass ratio is sufficiently small.
+This paper investigates the existence and properties of stable, uniformly rotating star-planet systems, i.e., mass ratio is sufficiently small.
 It is modeled by the Euler-Poisson equations.
@@
 1 Introduction
-In astrophysical fluid dynamics, gas stars and gas planets can be modeled as isolated self-gravitating fluid masses.
-A fundamental hydrodynamic description for Newtonian stars is provided by the Euler–Poisson system:
+In astrophysical fluid dynamics, gas stars and gas planets can be modeled as isolated self-gravitating fluid masses. A fundamental hydrodynamic description is provided by the Euler–Poisson system:
 $\begin{split}\partial_{t}\rho+\nabla\cdot\left({\rho v}\right)&=0,\\ \rho\partial_{t}v+\rho\left({v\cdot\nabla}\right)v+\nabla P(\rho)&=\rho\nabla V,\\ \Delta V&=-4\pi\rho.\end{split}$ (EP)
@@
 The pressure $P(\rho)$ depends solely on density.
-In this work, we adopt the polytropic equation of state $P(\rho)=K\rho^{\gamma}$.
-The motion of stars and planets has long been a classical subject in celestial mechanics, astrophysics, and mathematics.
-Historical models such as heliocentrism, proposed by Copernicus, and later studies by Newton and others on stellar rotation, form the foundation of rotating star or planet models.
-Early developments by Maclaurin, Jacobi, Poincaré, and Liapounov focused on incompressible, homogeneous or nearly homogeneous bodies like the Earth.
-For gaseous bodies such as Jupiter, or systems like galaxies, density inhomogeneity and compressibility need to be accounted for.
-The first attempts to construct axisymmetric rotating solutions to the compressible Euler–Poisson equations were made by astrophysicists Milne [47] and Zeipel [57] in the early 1920s, followed by Chandrasekhar [11] and Lichtenstein [39] in 1933.
-Further historical accounts can be found in [12, 33].
-Significant progress has been made for the single rotating star problem, where the star rotates about its axis.
-Two main approaches have been employed to prove the existence of such rotating stars.
-The first is variational method, as used by Auchmuty and Beals [3, 4] to establish the existence of axisymmetric equilibria for compressible fluids; see also [2, 4, 9, 13, 38, 16].
-The second is a perturbative method based on the implicit function theorem, applied near non-rotating Lane–Emden stars [29, 30, 28, 53, 54].
-Nonlinear dynamical stability of rotating stars was shown in [43, 44] via variational techniques, while stability theory for non-rotating stars can be found in [18, 32, 49].
-These methods are quite general and apply not only to single-body problems but also to $N$-body systems.
-Variational methods were used by McCann [46] for binary stars and are applied in this paper to star–planet systems.
-Perturbative approaches appear in the work of Alonso-Orán, Kepka, and Velázquez [1] on incompressible Euler–Poisson systems with an external particle, as well as in Lichtenstein’s construction of rotating binary stars [39].
-Each method has its advantages and limitations.
-For example, the variational approach typically requires $\gamma>\frac{4}{3}$ when $P(\rho)=K\rho^{\gamma}$ to prevent gravitational collapse (see [16, Remark 2.11 and Proposition 2.16] [46, Section 2], or [4, Section 1, 6 and 8]).
-In contrast, perturbative methods can handle the range $\frac{6}{5}<\gamma<2$ [30], but they rely on prior knowledge of solutions in special non-rotating cases.
-When considering two stellar objects, such as binary stars or a star–planet system, we begin with the simplest scenario where individual rotations are neglected, and only orbital revolution is considered.
+The existence of solutions under different settings has been established via variational methods (e.g., [3, 4, 2, 7, 8, 18, 24]) and perturbative approaches (e.g., [15, 16, 14, 29, 30, 1, 19]).
+In this work, we adopt the polytropic equation of state $P(\rho)=K\rho^{\gamma}$, unless otherwise stated.
+We mainly consider two stellar objects, under the simplifying assumption that individual rotations are neglected and only orbital revolution is considered.
 We still refer to such configurations as rotating systems.
@@
 $\omega^{2}\tilde{\rho}\left(x\right)P\left(x\right)+\nabla P\left(\tilde{\rho}\left(x\right)\right)-\tilde{\rho}\left(x\right)\left(\nabla V_{\tilde{\rho}}\left(x\right)\right)=0$ (EP’)
+where $\omega>0$ is the angular velocity, $\tilde{\rho}$ is a compactly supported density function, and $V_{\tilde{\rho}}(x)={\int_{\mathbb{R}^{3}}{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}\,dy}}$.
+As shown in [10, Section 2], if $\tilde{\rho}$ satisfies the reduced Euler–Poisson equation (EP’), then $\rho\left({t,y}\right)=\tilde{\rho}\left(R_{-\omega t}y\right)$ and $v\left({t,y}\right)=\omega\left({-y_{2},y_{1},0}\right)^{T}$ satisfy the original system (EP), where $R_{-\omega t}y$ is the rotation map given in Definition 2.1.
+In this case, it is easy to see that $\Delta v=0$, so $(\rho,v)$ also satisfies the Navier-Stokes-Poisson system.
+Moreover, since $\nabla\cdot v=0$, no compression occurs in this uniform rotation scenario.
+Nevertheless, based on physical intuition, we still regard this as a compressible model, corresponding to gaseous stars and planets.
 […]
-A notable feature of McCann’s binary stars is that each star admits a variational characterization as a Hamiltonian or energy minimizer under certain constraints, while also being interpretable as a perturbation of simpler objects—such as non-rotating Lane–Emden stars and point-mass relative equilibria (see subsection 5.3 and Section 6 of this paper or [31, Section 3.2]).
+A notable feature of McCann’s framework is that each star admits a variational characterization as a Hamiltonian or energy minimizer under certain constraints, while also being interpretable as a perturbation of simpler objects—such as non-rotating Lane–Emden stars and point-mass relative equilibria (see subsection 5.3 and Section 6 of this paper or [17, Section 3.2]).
 This characterization is crucial for showing that the support of an energy-constrained minimizer lies in the interior of balls of certain radii.
 Consequently, the minimizer $\rho$ is in fact a local minimizer in the Wasserstein $L^{\infty}$ distance sense.
-This ensures that $\nabla P(\rho)$ is well-defined throughout $\mathbb{R}^{3}$, and that (EP’) holds in the whole space $\mathbb{R}^{3}$ (see Theorem 3.1, [46, Theorem 2.1], [15, Section 2]).
-Beyond existence, many questions remain for two-body system.
-For instance, Jang and Seok [31] study the asymptotic profiles, uniqueness, and orbital stability of McCann’s uniformly rotating binary stars.
-In addition, conjectures concerning the maximal possible number of connected components of the support of solutions are discussed in Section 7 of this paper.
-Researchers have also considered systems with more than two stellar objects, such as galaxies or general $N$-star systems.
-For rotating galaxies, single rotating galaxies were constructed in [52, 53] in the spirit of Lichtenstein and Heilig [28, 39].
-There is an extensive literature on non-rotating galaxies: see [6, 7, 24, 50] and references therein.
-The orbital stability of stationary solutions has seen considerable progress over the past two decades [26, 21, 22, 23, 20, 25, 34, 35, 36, 37, 50, 56], with recent work on linearly oscillating galaxies in [27].
-In [10], Campos, Del Pino, and Dolbeault constructed $N$-body rotating galaxies by connecting to relative equilibria in $N$-body dynamics with small uniform angular velocity, perturbing radial equilibria, and solving for the gravitational potential rather than the density, using a finite-dimensional reduction.
-Although natural, Jang and Seok remark in [31] that the framework in [10] is not ideal for stability analysis.
-They adapt Rein’s reduction method [49, 50] to extend the discussion of binary star solutions to establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [31].
-Furthermore, they note that for $N\geq 3$, uniformly rotating $N$-body stellar configurations lack a variational structure analogous to the binary case and are not generally expected to be stable [31].
+This ensures that $\nabla P(\rho)$ is well-defined throughout $\mathbb{R}^{3}$, and that (EP’) holds in the whole space $\mathbb{R}^{3}$ (see Theorem 3.1, [24, Theorem 2.1], [10, Section 2]).
+Beyond existence, many questions remain for two-body systems.
+For instance, conjectures concerning the maximal possible number of connected components of the support of solution are discussed in Section 7 of this paper.
+Jang and Seok [17] study the asymptotic profiles, uniqueness, and orbital stability of McCann’s uniformly rotating binary stars.
+They also establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [17], a work that extends the discussion of binary star solutions by adapting Rein’s reduction method [26, 27].
 This paper presents the core results extracted from the author’s Master’s thesis [9].
-Building upon the preparatory work in two companion papers [15, 16] originating from the same thesis, it investigates the existence and qualitative properties of stable, uniformly rotating star–planet systems governed by the Euler–Poisson equations.
-This work thereby extends McCann’s framework for binary stars [46] to the setting of small mass ratios, as established in Theorems 2.12 and 2.13.
+It investigates the existence and qualitative properties of stable, uniformly rotating star–planet systems governed by the Euler–Poisson equations, as established in Theorems 2.12 and 2.13.
+Although this paper builds upon the preparatory work in two companion papers [10, 11] originating from the same thesis [9], it is largely self-contained.
+Relevant auxiliary results can also be found in other existing literature cited at the corresponding locations herein, with the author’s papers [10, 11] providing refined and more complete proofs.
 The paper is structured as follows.
@@
 Section 3 collects preliminary results about two-body systems and non-rotating single star cases from [24, 4, 20, 10, 11].
-Section 4, 5 and 6 are the core of this paper, where we will prove existence of star-planet systems.
+Sections 4, 5 and 6 are the core of this paper, where we will prove existence of star-planet systems.
 Sections 7 studies distances between (possible) different connected components of planets or stars, leading to the conjecture that minimizer may consist of at most two connected components, corresponding to exactly one star and exactly one planet respectively.
@@
 2.1 Variational formulation
-We introduce some notations and variational formulation here, which based on McCann’s construction of binary star solutions, see [46] and [15].
+We introduce some notations and variational formulation here, which are based on McCann’s construction of binary star solutions, see [24] and [10].
 Definition 2.1 (Notations).
@@
   1. (i)
-    k{k} dimensional open ball: $B_{R}(a):=\left\{x\in\mathbb{R}^{k}\mid|x-a\mid<R\right\}$ for given $a\in\mathbb{R}^{k},R\in\mathbb{R}$.
+    The projection operator of $x$ to the $x_{1}x_{2}$ plane: $P_{12}(x)=P_{12}\left(x_{1},x_{2},x_{3}\right):=\left(x_{1},x_{2},0\right)$.
   2. (ii)
-    The projection operator of $x$ to the $x_{1}x_{2}$ plane: $P_{12}(x)=P_{12}\left(x_{1},x_{2},x_{3}\right):=\left(x_{1},x_{2},0\right)$.
+    A bilinear form $\langle\cdot,\cdot\rangle_{2}:\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}:\forall x,y\in\mathbb{R}^{3},\langle x,y\rangle_{2}=P_{12}(x)\cdot P_{12}(y)=x_{1}y_{1}+x_{2}y_{2}$.
   3. (iii)
-    A bilinear form $\langle\cdot,\cdot\rangle_{2}:\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}:\forall x,y\in\mathbb{R}^{3},\langle x,y\rangle_{2}=P_{12}(x)\cdot P_{12}(y)=x_{1}y_{1}+x_{2}y_{2}$.
+    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=\sqrt{x_{1}^{2}+x_{2}^{2}}$.
   4. (iv)
-    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=x_{1}^{2}+x_{2}^{2}$.
+    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
   5. (v)
-    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
-  6. (vi)
-    Positive part $\lambda]_{+}:=\max\{\lambda,0\}$.
-  7. (vii)
     $W^{\infty}$ metric: Wasserstein $L^{\infty}$ metric. (See Definition 2.8)
-  8. (viii)
-    Given $\delta>0$ the $\delta$-neighbourhood of $\Omega\subset\mathbb{R}^{3}$ to be the set Ω+B_δ(0):=⋃_y ∈Ω{x ∈R^3∣— x-y —¡δ}
-  9. (ix)
-    $\mu$: Lebesgue measure.
 The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$.
@@
 Moreover, to define internal energy, we need to first consider the equation of state for pressure $P(\rho)$.
-In this paper, we take the polytropic law, i.e.
+In this paper, we take the polytropic law, i.e.,
 $P(\rho)=K\rho^{\gamma}$ (1)
@@
 $A(s):=s\int_{0}^{s}P(\tau)\tau^{-2}d\tau=\frac{K}{\gamma-1}s^{\gamma}$ (2)
-We first give a “potential admissible class” for $\rho$ and $v$ as the following:
-$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=1\right\}$ (3)
+We first give “admissible classes” for $\rho$ and $v$ as the following:
+$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ (3)
 $V\left(\mathbb{R}^{3}\right)$ $:=\left\{v:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}\mid v\text{ is measurable.}\right\}$ (4)
@@
 $E(\rho,v):=U(\rho)-\frac{G(\rho,\rho)}{2}+T(\rho,v)$ (5)
-$U(\rho):=\int A(\rho(x))d^{3}x$ (6)
-$G(\sigma,\rho):=\int V_{\sigma}\rho=\iint\frac{\rho(x)\sigma(y)}{|x-y|}d^{3}yd^{3}x$ (7)
-$T(\rho,v):=\frac{1}{2}\int|v|^{2}\rho$ (8)
-Here $A(\rho)$ is a convex function given in (2), and $U(\rho)$ is internal energy.
+$U(\rho):=\int_{\mathbb{R}^{3}}A(\rho(x))\,dx$ (6)
+$G(\sigma,\rho):=\int_{\mathbb{R}^{3}}V_{\sigma}\rho\,dx=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\rho(x)\sigma(y)}{|x-y|}\,dy\,dx$ (7)
+$T(\rho,v):=\frac{1}{2}\int_{\mathbb{R}^{3}}|v|^{2}\rho\,dx$ (8)
+Here $A(\rho)$ is a convex function given in (2), and $U(\rho)$ is the internal energy.
 We can choose units so that the total mass of fluid is one and the gravitational constant $G=1$, then $V_{\rho}$ represents the gravitational potential of the mass density $\rho(x)$
-$V_{\rho}(x):=\int\frac{\rho(y)}{|y-x|}d^{3}y$ (9)
-Hence $G(\rho,\rho)$ is gravitational interaction energy.
-We can also see $T(\rho,v)$ is kinetic energy.
+$V_{\rho}(x):=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|y-x|}\,dy$ (9)
+Hence $G(\rho,\rho)$ is the gravitational potential energy (also called gravitational interaction energy), and $T(\rho,v)$ is the kinetic energy.
 Remark 2.2.
-Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [15, Proposition 2.10].
-Hence $E(\rho,v)$ is well-defined.
+Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [10, Section 2]. Hence $E(\rho,v)$ is well-defined.
 We can choose a frame of reference in which the center of mass
-$\bar{x}(\rho):=\frac{\int x\rho(x)d^{3}x}{\int\rho(x)d^{3}x}$ (10)
+$\bar{x}(\rho):=\frac{\int_{\mathbb{R}^{3}}x\rho(x)\,dx}{\int_{\mathbb{R}^{3}}\rho(x)\,dx}$ (10)
 is at rest.
-We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$.
+We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and fixed angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$.
 The fluid angular momentum $\boldsymbol{J}(\rho,v)$ is given by:
-$\boldsymbol{J}(\rho,v):=\int(x-\bar{x}(\rho))\times v\rho(x)d^{3}x$ (11)
+$\boldsymbol{J}(\rho,v):=\int_{\mathbb{R}^{3}}(x-\bar{x}(\rho))\times v\rho(x)\,dx$ (11)
 We denote by $J_{z}$ the z-component of $\boldsymbol{J}$, that is, $J_{z}(\rho,v):=\hat{e}_{z}\cdot\boldsymbol{J}(\rho,v)$, where $\hat{e}_{z}=(0,0,1)^{T}$.
@@
 Since the z-component of the angular momentum is specified, the moment of inertia $I(\rho)$ of $\rho$ in the direction of $\hat{e}_{z}$ will be relevant.
-That is, we define $I(\rho)$ to be moment of inertia of $\rho$ in the direction of $\hat{e}_{z}$:
-$I(\rho):=\int\rho r^{2}(x-\bar{x}(\rho))$ (12)
+That is, we define the moment of inertia of $\rho$ in the direction of $\hat{e}_{z}$, denoted by $I(\rho)$, as follows:
+$I(\rho):=\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho))\,dx=\int_{\mathbb{R}^{3}}\rho(x)\left((x_{1}-\bar{x}(\rho)_{1})^{2}+(x_{2}-\bar{x}(\rho)_{2})^{2}\right)\,dx$ (12)
 where $r$ is given in Definition 2.1.
 Remark 2.3.
-When $\rho$ has positive mass, we have $I(\rho)>0$. See [15, Remark 2.9].
+When $\rho$ has positive mass, we have $I(\rho)>0$. See [10, Section 2].
 Since the star and planet are separated, it is convenient to describe the relations between the total moment of inertia $I\left(\rho_{m}+\rho_{1-m}\right)$ and $I\left(\rho_{m}\right),I\left(\rho_{1-m}\right)$.
-Lemma 2.4 (Expansion of Moment of Inertia [15, Lemma 2.15]).
-Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int\rho=m_{1}<\infty,\int\sigma=m_{2}<\infty$.
+Lemma 2.4 (Expansion of Moment of Inertia [10, Section 2]).
+Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int_{\mathbb{R}^{3}}\rho\,dx=m_{1}<\infty$, $\int_{\mathbb{R}^{3}}\sigma\,dx=m_{2}<\infty$.
 $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
@@
     $I(\rho+\sigma)=I(\rho)+I(\sigma)+\frac{m_{1}m_{2}}{m_{1}+m_{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$ (13)
-Let support of $\rho$ be the smallest closed set carrying the full mass of $\rho$ (denoted by spt $\rho$).
+Let the support of $\rho$ be the smallest closed set carrying the full mass of $\rho$ (denoted by spt $\rho$).
 Intuitively we hope spt $\rho$ to be compact as this aligns with the case of a star-planet system and ensures a finite moment of inertia.
@@
 ${R}_{a}\left(\mathbb{R}^{3}\right):=\left\{\rho\in{R}\left(\mathbb{R}^{3}\right)\mid\bar{x}(\rho)=a;\text{ spt }\rho\text{ is bounded.}\right\}$ (14)
-Since the energy is translation-invariant, we may just consider looking for a energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, which means $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$.
+Since the energy is translation-invariant, we may just consider looking for an energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, which means $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$.
 Thanks to Theorem 3.1 below, it turns out the problem of minimizing $E(\rho,v)$ locally in $R_{0}(\mathbb{R}^{3})\times V(\mathbb{R}^{3})$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ locally in $R_{0}(\mathbb{R}^{3})$.
@@
 Remark 2.5.
-Given $\rho\in R(\mathbb{R}^{3})$, if we further know $\rho\in L^{\infty}(\mathbb{R}^{3})$, then we know $U(\rho)$ is finite by interpolation inequality (Proposition A.2).
+Given $\rho\in R(\mathbb{R}^{3})$, if we further know $\rho\in L^{\infty}(\mathbb{R}^{3})$, then we know $U(\rho)$ is finite by interpolation inequality [6, Section 4.2]:
+$\|\rho\|_{L^{r}}\leq\|\rho\|_{L^{p}}^{\theta}\|\rho\|_{L^{q}}^{1-\theta},\quad\text{for }p\leq r\leq q,$ (17)
+where $\frac{1}{r}=\frac{\theta}{p}+\frac{1-\theta}{q}$ for some $\theta\in[0,1]$.
 This observation, together with Remark 2.2 and Remark 2.3, tells us $E_{J}(\rho)$ is well-defined and finite.
-For more general pressure assumptions, see [15, Lemma 5.2].
+Similar results hold for more general pressures; see [10, Section 5].
 We intend to solve (EP’) by minimizing $E_{J}(\rho)$ and using results from calculus of variations, which involves the variational derivative of energy.
 We therefore now proceed to define it.
-We define the perturbation set $P_{\infty}(\rho)$, which depends on energy minimizer $\rho$, as the following:
+We first define the perturbation set $P_{\infty}(\rho)$, which depends on energy minimizer $\rho$, as the following:
 $P_{\infty}(\rho):=\bigcup_{R<\infty}P_{R}(\rho)$ (18)
@@
 […]
-One can see $P_{\infty}(\rho)$ is a convex cone. And we have the following result:
-Lemma 2.6 (Differentiability of Energy $E_{J}(\rho)$ [15, Lemma 5.3, Remark 5.4]).
+One can see $P_{\infty}(\rho)$ is a convex cone. And then we have the following result:
+Lemma 2.6 (Differentiability of Energy $E_{J}(\rho)$ [10, Section 5]).
 Given $\rho\in mR(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, i.e.
 it is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$.
-Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int E_{J}^{\prime}(\rho)\sigma$  [^1] To remain consistent with the notation in [4, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion..
+Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int_{\mathbb{R}^{3}}E_{J}^{\prime}(\rho)\sigma\,dx$  [^1] To remain consistent with the notation in [4, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion..
 The function $E_{J}^{\prime}(\rho)$ on the right‑hand side is given by
@@
 Moreover, $E(\rho,v)$ is local minimum, and $({\rho}(t,x),v(t,x))$ is solution to (EP).
-We will discuss these results in details in Theorem 3.1.
+We will discuss these results in detail in Theorem 3.1.
 For prescribed angular momentum, we also know the energy $E_{J}(\rho)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$.
-However, as in Morgan [48], McCann [46, Section 3] demonstrates that this bound — although approached — will not be attained, see also [45].
+However, as in Morgan [25], McCann [24, Section 3] demonstrates that this bound — although approached — will not be attained, see also [23, 10].
 Therefore, instead of searching for a global energy minimizer in star-planet system, we adopt McCann’s strategy [24]: first establishing the existence of a constrained minimizer of $E_{J}(\rho)$, and then proving that it is also a local minimizer under the appropriate topology induced by Wasserstein $L^{\infty}$ distance.
-Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu^{2}}$, where $\mu=m(1-m)$ is their reduced mass.
-Note that the same symbol $\mu$ is also used as for Lebesgue measure, provided no ambiguity arises.
+Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu_{r}^{2}}$, where $\mu_{r}=m(1-m)$ is their reduced mass.
 For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ is minimized subject to the constraint
-$W_{m}:=\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int\rho_{m}=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (21)
+$W_{m}:=\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int_{\mathbb{R}^{3}}\rho_{m}\,dx=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (22)
 where $\Omega_{m}$ and $\Omega_{1-m}$ are subsets of $\mathbb{R}^{3}$, which are given in the following. Note $W_{m}$, $\Omega_{m}$ and $\Omega_{1-m}$ are actually related to both $m$ and $J$.
 Fix two points $y_{m}$ and $y_{1-m}$ in $\mathbb{R}^{3}$ from the plane $z=0$, which are separated by
-$\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{m^{2}(1-m)^{2}}$ (22)
+$\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{m^{2}(1-m)^{2}}$ (23)
 i.e.
@@
 Remark 2.7.
-The reason we set the separation $\eta=\frac{J^{2}}{\mu^{2}}$ in the definitions above is inspired by the Kepler problem.
-Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass.
-Assume their separation is $d$, the gravitational energy plus kinetic energy is $\frac{\mu}{d}+\frac{J^{2}}{2\mu d^{2}}$, which reaches its minimum at separation $d=\eta$.
-In the star-planet model, when $\eta$ is large, the distance between two objects is large, the gravitational interaction becomes weak.
+The reason we set the separation $\eta=\frac{J^{2}}{\mu_{r}^{2}}$ in the definitions above is inspired by the Kepler problem.
+Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass, if we assume their separation is $d$, then the gravitational energy plus kinetic energy is $\frac{\mu_{r}}{d}+\frac{J^{2}}{2\mu_{r}d^{2}}$, which reaches its minimum at separation $d=\eta$.
+In the star-planet model, when $\eta$ is large, the distance between two objects is large, and the gravitational interaction becomes weak.
 The system thus approximates the ideal case of two isolated, non-rotating, and non-interacting bodies.
-Therefore, it can be proven that the supports of the objects remain bounded, which sizes are hence negligible compared to the distance between them.
+Therefore, it can be proven that the supports of the objects remain bounded, whose sizes are hence negligible compared to the distance between them.
 It implies the distance between their centers of mass asymptotically approaches that of the point mass model.
-We will discuss these facts in more details in Section 5 and Section 6, assuming $m$ is sufficiently small, and then establish their connection to the existence of local energy minimizers under topology induced by $W^{\infty}$ (Wasserstein $L^{\infty}$) distance.
-For the case $J$ is sufficiently large, one can also check McCann’s paper [46, Section 6].
-We note the choice of topology for ${R}\left(\mathbb{R}^{3}\right)$ is quite delicate: for $J>0$, it turns out local energy minimizer for $E_{J}(\rho)$ will not exist if the topology of $R(\mathbb{R}^{3})$ is inherited from a topological vector space ([46, Remark 3.7] and [15, Proposition 5.15]).
+We will discuss these facts in more detail in Section 5 and Section 6, assuming $m$ is sufficiently small.
+We will then establish their connection to the existence of local energy minimizers under the topology induced by $W^{\infty}$ (Wasserstein $L^{\infty}$) distance.
+For the case $J$ is sufficiently large, one can also check McCann’s paper [24, Section 6].
+We note the choice of topology for ${R}\left(\mathbb{R}^{3}\right)$ is quite delicate: for $J>0$, it turns out local energy minimizer for $E_{J}(\rho)$ will not exist if the topology of $R(\mathbb{R}^{3})$ is inherited from a topological vector space ([24, Remark 3.7] and [10, Section 5]).
 Roughly speaking, the reason is that for any reasonable candidate local minimizer $\rho$, one can always split off a portion of it with very small mass and shift the portion arbitrarily far away, thereby decreasing the energy.
-However, the resulting configuration still lies in a small neighborhood of $\rho$ with respect to such a topology, which implies that $\rho$ cannot be a local minimizer.
-See also Remark 3.10.
-Hence we need another topology, i.e. the topology induced by the $W^{\infty}$ distance.
+However, the resulting configuration still lies in a small neighborhood of $\rho$ with respect to such topology, which implies that $\rho$ cannot be a local minimizer.
+Hence we need another topology, i.e., the topology induced by the $W^{\infty}$ distance.
 Definition 2.8 (Wasserstein $L^{\infty}$ distance).
@@
 $f_{\#}:M(S)\rightarrow M(X)$ ($g_{\#}$ is similar) is the corresponding push forward operator defined by
-$f_{\#}\tilde{\mu}(B):=\tilde{\mu}\left(f^{-1}(B)\right)\quad\text{for any Borel set }B\text{ in }X$
+[…]
 We call $f,g$ the transport maps.
 Thanks to Strassen’s Theorem, one can check $W^{\infty}$ is truly a metric, as explained in Givens and Shortt [13].
-We gives two properties of $W^{\infty}$ distance.
-For other properties, one can see [46, Lemma 5.1] and [15, Lemma 4.7].
-Lemma 2.9 (Simple Properties of the Wasserstein $L^{\infty}$ Metric [46, Lemma 5.1]).
+We give two properties of the $W^{\infty}$ distance.
+For other properties, one can see [24, Lemma 5.1] and [10, Section 4].
+Lemma 2.9 (Simple Properties of the Wasserstein $L^{\infty}$ Metric [24, Lemma 5.1]).
 Let $\rho,\kappa$ in $t{R}\left(\mathbb{R}^{3}\right)$, then
   1. (ii)
-    $W^{\infty}(\rho,\kappa)$ does not exceed the diameter of $\operatorname{spt}(\rho-\kappa)$
+    $W^{\infty}(\rho,\kappa)$ does not exceed the diameter of $\operatorname{spt}(\rho-\kappa)$;
   2. (ii)
-    if $W^{\infty}(\rho,\kappa)<\delta$, each connected component of the $\delta$-neighbourhood of spt $\rho$ has the same mass for $\kappa$ as for $\rho$
+    if $W^{\infty}(\rho,\kappa)<\delta$, each connected component of the $\delta$-neighbourhood of spt $\rho$ has the same mass for $\kappa$ as for $\rho$.
+Here the $\delta$-neighbourhood of some $\Omega\subset\mathbb{R}^{3}$ is defined as $\bigcup_{y\in\Omega}\left\{x\in\mathbb{R}^{3}\mid|x-y|<\delta\right\}$.
 Remark 2.10.
-The advantage of the topology induced by Wasserstein $L^{\infty}$ metric is that a neighborhood of $\rho$ does not include such “portion-shifted-to-far-away” configurations mention above due to Lemma 2.9 (ii), while thanks to Lemma 2.9 (i), the neighborhood still contains all perturbations of the form $\rho+t\sigma$ for all $t$, where $\sigma$ is supported on a set which diameter is small enough, with $\int\sigma=0$.
-Hence this setting still allows one to legitimately consider variational derivatives (18) and consequently Euler-Lagrange equations (EL) (see also Remark 3.4.).
+The advantage of the topology induced by Wasserstein $L^{\infty}$ metric is that a neighborhood of $\rho$ does not include such “portion-shifted-to-far-away” configurations mention above due to Lemma 2.9 (ii), while thanks to Lemma 2.9 (i), the neighborhood still contains all perturbations of the form $\rho+t\sigma$ for all $t$, where $\sigma$ is supported on a set whose diameter is small enough, with $\int_{\mathbb{R}^{3}}\sigma\,dx=0$.
+Hence this setting still allows one to legitimately consider variational derivatives (19) and consequently Euler-Lagrange equations (EL) and Euler-Poisson equations (EP’) (see also Remark 3.4).
 Remark 2.11.
-According to Lemma 2.9, to prove $\rho(m)$ is a $W^{\infty}$ local minimizer (and hence solution to (EP’) due to Theorem 3.1), it suffices to show $\rho(m)$ is a constraint minimizer on $W_{m}$ such that its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$.
-In fact, once we know $\rho(m)$’s support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$, one can also show $\rho(m)$ is a solution to (EP’), as well as most results in Theorem 3.1 (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) without specifying $\rho(m)$ is a $W^{\infty}$ local minimizer, see [15, Remark 3.10].
+According to Lemma 2.9, to prove $\rho(m)$ is a $W^{\infty}$ local minimizer (and hence a solution to (EP’) due to Theorem 3.1), it suffices to show $\rho(m)$ is a constraint minimizer on $W_{m}$ such that its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$.
+In fact, once we know $\rho(m)$’s support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$, one can also show $\rho(m)$ is a solution to (EP’), as well as most results in Theorem 3.1 (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) without specifying $\rho(m)$ is a $W^{\infty}$ local minimizer, see [10, Section 3].
 It motivates our choice to define $\Omega_{m}$ and $\Omega_{1-m}$ with their radii that increase as $\eta$ increases, and then show the size of stars or planets will not expand too much as $\eta$ increases.
-We will show those results for small $m$ later, inspired by McCann’s arguments for large $J$ [46, Section 6].
+We will show those results for small $m$ later, inspired by McCann’s arguments for large $J$ [24, Section 6].
 2.2 Statement of Main Results
@@
   • $(iii)$
-    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int{\frac{\rho(m)(y)}{\left|{y-x}\right|}d^{3}y}}$.
+    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int_{\mathbb{R}^{3}}{\frac{\rho(m)(y)}{\left|{y-x}\right|}\,dy}}$.
 Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
 The support for the star density spt $\rho_{1-m}$ is contained in a ball of radius $R(J)$ for all small $m$.
-Actually, the results can be true in a larger range of index $\gamma$ if we don’t require the support of the planet to shrink as the mass goes to 0.
+Actually, the results can be true in a larger range of index $\gamma$ if we do not require the support of the planet to shrink as the mass goes to 0.
 Theorem 2.13 (Existence of Star-Planet Systems for $\gamma>\frac{3}{2}$).
@@
   • $(iii)$
-    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int{\frac{\rho(m)(y)}{\left|{y-x}\right|}d^{3}y}}$.
+    $\rho(m)$ is continuous and satisfies (EP’) with $V_{\rho(m)}(x)={\int_{\mathbb{R}^{3}}{\frac{\rho(m)(y)}{\left|{y-x}\right|}\,dy}}$.
 Moreover, the uniformly rotating fluid $(\widetilde{\rho(m)},v)$ satisfies (EP) with $V(t,x)=V_{{\rho(m)}}\left(R_{-\omega t}x\right)$, here $({\widetilde{\rho(m)}}(t,x),v(t,x))=\left({\rho(m)}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
@@
 To achieve this, we first prove the existence of constrained minimizers $\rho(m)$ of $E_{J}(\rho)$ on $W_{m}$ when the mass ratio is sufficiently small in Section 4.
-In Section 5, we prove convergence of scaling densities in $L^{p}$, uniform $L^{\infty}$ bound, and then derives quantitative bounds on the supports of minimizers, .
+In Section 5, we prove convergence of scaling densities in $L^{p}$, uniform $L^{\infty}$ bound, and then derive quantitative bounds on the supports of minimizers.
 In Section 6 we estimate the center of mass separation and show that $\rho(m)$ can be situated within the interior of the considered domains we consider, that is, $\text{dist }\left(\text{spt}~\rho(m),~\mathbb{R}^{3}\backslash\left(\Omega_{m}\cup\Omega_{1-m}\right)\right)>0$, and then we verify $\rho(m)$ satisfies results in Theorem 2.12 and Theorem 2.13.
@@
 Recall that in the calculus of variations, if an energy functional admits a minimizer, then this minimizer satisfies Euler–Lagrange equation.
-When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, Theorem A.1 or Theorem 3.1.
+When the minimization is carried out over a restricted class, the Euler–Lagrange equation may take the form of an inequality or involve taking the positive part; see, for example, Theorem 3.1 or [4, Section 2].
 In this section, we review some variational‑based results and their extensions.
 In subsection 3.1, we discuss McCann’s work on Wasserstein $L^{\infty}$ local minimizers and existence of binary star system results.
-In subsection 3.2, we review some results for single-star system.
+In subsection 3.2, we review some results for single-star systems.
 3.1 Results for 2-body systems
-In this subsection, we review McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution with slight modification.
+In this subsection, we review McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution with slight modifications.
 We also make some remarks on them.
-The following are McCann’s results on $W^{\infty}$ local minimizers.
-Theorem 3.1 (Properties of $W^{\infty}$-Local Energy Minimizers [46, Section 2][15, Section 2]).
+The following are McCann’s results for $W^{\infty}$ local minimizers.
+Theorem 3.1 (Properties of $W^{\infty}$-Local Energy Minimizers [24, Section 2][10, Section 2]).
 Let $J>0$.
@@
 It describes uniformly rotating fluid, in the sense that after defining $({\widetilde{\rho}}(t,x),v(t,x))=\left({\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
-Notice that $E(\rho,v)=E(\widetilde{\rho},v)=E_{J}(\widetilde{\rho})=E_{J}(\rho)$.
+In particular, $E(\rho,v)=E(\widetilde{\rho},v)=E_{J}(\widetilde{\rho})=E_{J}(\rho)$.
   3. (iii)
@@
     $A^{\prime}(\rho(x))=\left[\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+V_{\rho}(x)+\lambda_{i}\right]_{+}$ (EL)
-    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component;
+    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component. Here $\cdot$$_{+}$ is the nonnegative (positive) part function defined as $\lambda]_{+}:=\max\{\lambda,0\}$;
   5. (v)
@@
 Remark 3.2.
-In Theorem 3.1 we only mention the case where $P(\rho)$ satisfies polytropic law (1), but the results hold true for a more general form of $P(\rho)$, see for instance [15, Theorem 2.25].
+In Theorem 3.1 we only mention the case where $P(\rho)$ satisfies polytropic law (1), but the results hold true for a more general form of $P(\rho)$, see for instance [10, Section 2].
 Remark 3.3.
 Notice in Theorem 3.1, since the local minimizer $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\bar{x}(\rho)=(0,0,0)^{T}$ in (EL).
-However, one can generalize the case where the center of mass is not at the origin and $\rho$ is constrained minimizer instead of $W^{\infty}$ local minimizer, hence we retain the term $\bar{x}(\rho)$ in (EL) in Theorem 3.1 (iv) and (EL) holds on $\Omega_{i}$ instead of $\delta$-neighborhood of $\Omega_{i}$.
-This is consistent with the form in McCann’s paper [46, Section 6].
+However, one can generalize to the case where the center of mass is not at the origin, hence we retain the term $\bar{x}(\rho)$ in (EL) in Theorem 3.1 (iv).
+This is consistent with the form in McCann’s paper [24, Section 6].
+Furthermore, if $\rho$ is a constrained minimizer instead of a $W^{\infty}$ local minimizer, (EL) still holds on $\Omega_{i}$.
 Remark 3.4.
 McCann initially introduced and proved Theorem 3.1 in [24].
-We later complemented this work in [15] by showing why (EP) also holds on the boundary of support of minimizer.
+We later complemented this work in [10] by showing why (EP) also holds on the boundary of the support of minimizer.
 Additionally, in [10] we verified the local minimizer $\rho$ has finite internal energy $U(\rho)<\infty$, hence $E_{J}$ is indeed differentiable at $\rho$ due to Lemma 2.6.
@@
 And we know the following theorem:
-Theorem 3.5 (Existence of Binary Stars [46, Theorem 6.1, Corollary 6.2]).
+Theorem 3.5 (Existence of Binary Stars [24, Theorem 6.1, Corollary 6.2]).
 […]
@@
 3.2 Results for Non-rotating Bodies
-In this subsection, we introduce the existence and uniqueness of non-rotating solutions of Euler-Poisson equations, corresponding the single-star case.
+In this subsection, we introduce the existence and uniqueness of non-rotating solutions of Euler-Poisson equations, corresponding to the single-star case.
 We also indicate some useful properties about the solutions.
@@
     $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave.
-In particular, for $m>0$, $e_{0}(m)<0$, $e_{0}^{\prime}\left(m^{+}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{+}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ and $e_{0}^{\prime}\left(m^{-}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{-}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ exists, and $e_{0}^{\prime}(m^{+})<e_{0}^{\prime}(m^{-})<0$;
+In particular, for $m>0$, $e_{0}(m)<0$, $e_{0}^{\prime}\left(m^{+}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{+}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ and $e_{0}^{\prime}\left(m^{-}\right)\coloneq\lim\limits_{\widetilde{m}\rightarrow m^{-}}{e_{0}^{\prime}\left(\widetilde{m}\right)}$ exists, and we have $e_{0}^{\prime}(m^{+})\leq e_{0}^{\prime}(m^{-})<0$;
   3. (iii)
@@
   4. (iv)
-    $\left\|\sigma_{m}\right\|_{L^{\infty}}\leq C(m)$;
+    $\left\|\sigma_{m}\right\|_{L^{\infty}}\leq C(m)$, and $U(\sigma_{m})<\infty$ due to Remark 2.5;
   5. (v)
@@
 Remark 3.8.
-Since $\sigma$ has mass 1 and satisfies Theorem 3.6 (iii), $\sigma\in L^{1}\cap L^{\infty}$, then $U(\sigma)$ is finite by interpolation inequality Proposition A.2.
-Hence the equation for $\lambda$ in Theorem 3.6 (vii) makes sense.
-Remark 3.9.
 In Theorem 3.6, we only mention the case where $P(\rho)$ satisfies polytropic law (1), but some results hold true for a more general form of $P(\rho)$, see for instance [4, 20, 24, 11].
 Auchmuty and Beals initially concluded that (EP’) is satisfied in the region where $\rho>0$ [4, Theorem A].
-We later improved the result in [16, Theorem 2.7] by showing the equation holds over the entire $\mathbb{R}^{3}$ space.
-Theorem 3.6 (ii) was stated without proof by McCann [46, Section 3], who mentioned it could be demonstrated using a method analogous to that in [40]—which is in the quantum mechanical framework.
+We later improved the result in [11, Section 2] by showing the equation holds over the entire $\mathbb{R}^{3}$ space.
+Theorem 3.6 (ii) was stated without proof by McCann [24, Section 3], who mentioned it could be demonstrated using a method analogous to that in [20]—which is in the quantum mechanical framework.
 A proof for the polytropic case was later given in [11].
-Note that without polytropic law assumption, the uniqueness of minimizer, the existence of $e^{\prime}_{0}(m)$ and the equality relations between $\lambda_{m}$, $U(\sigma$) and $e^{\prime}_{0}(m)$ in (vii) and (viii) are not guaranteed.
+Note that without polytropic law assumption, the uniqueness of minimizer, the existence of $e^{\prime}_{0}(m)$, and the equality relations between $\lambda_{m}$, $U(\sigma$) and $e^{\prime}_{0}(m)$ in (vii) and (viii) are not guaranteed.
 Instead, we have $e_{0}^{\prime}\left(m^{+}\right)\leq\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$.
 Note $e_{0}^{\prime}\left(m^{-}\right)<0$ if $m>0$ due to Theorem 3.6 (ii).
-Although uniqueness result in general case is unknown, uniqueness can still be established under assumptions slightly weaker than the polytropic law; see [46, 16].
-Remark 3.10.
-Intuitively, when the distance between two celestial bodies is sufficiently large, the kinetic energy and the graviational interaction energy becomes negligible.
-Consequently, $E_{J}$ will approach $e_{0}$ without equaling it.
-This intuition is indeed correct [16, Remark 2.43 and Remark 2.44].
-In Section 5, we will also compare $e_{0}$ and the energy $E_{0}(\widetilde{\rho}_{m})$ for scaling density $\widetilde{\rho_{m}}$.
-This comparison allows us to apply concentration-compactness lemma in [42] and will also provide a clearer mathematical understanding of the relation between $e_{0}$ and $E_{J}$.
+Although uniqueness result in the general case is unknown, uniqueness can still be established under assumptions slightly weaker than the polytropic law; see [24, 11].
 It turns out under polytropic law (1) assumption, minimizers with different masses satisfy scaling relation, which helps to strengthen Theorem 3.6 (ii), as shown in the following results:
-Theorem 3.11 (Relations between Minimizers and Minimal Energies with Different Mass [16, Theorem 3.2]).
+Theorem 3.9 (Relations between Minimizers and Minimal Energies with Different Mass [11, Section 3]).
 Let $\sigma$ be the minimizer with mass 1 and the corresponding minimal energy is $e_{0}=E_{0}(\sigma)$, then any other minimizer with mass $m$ can be represented as $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$, where $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, and the corresponding minimal energy is $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$.
 Sketch of Proof.
-Recall (19) that $E_{0}(\sigma)=U(\sigma)-\frac{G(\sigma,\sigma)}{2}=\int A(\sigma)-\frac{1}{2}\iint\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y$, where $A(\sigma)=\frac{K}{\gamma-1}\sigma^{\gamma}$.
+Recall (20) that $E_{0}(\sigma)=U(\sigma)-\frac{G(\sigma,\sigma)}{2}=\int_{\mathbb{R}^{3}}A(\sigma)\,dx-\frac{1}{2}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy$, where $A(\sigma)=\frac{K}{\gamma-1}\sigma^{\gamma}$.
 Then let $\sigma_{\widetilde{A},B}=\widetilde{A}\sigma(Bx)$, we can compute
-$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int A\left(\sigma_{\widetilde{A},B}\right)$ (27)
-$=\int\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}$
-$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}d^{3}(Bx)$
+$U\left(\sigma_{\widetilde{A},B}\right)$ $=\int_{\mathbb{R}^{3}}A\left(\sigma_{\widetilde{A},B}\right)\,dx$ (28)
+$=\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\widetilde{A}\sigma(Bx))^{\gamma}\,dx$
+$=\frac{\widetilde{A}^{\gamma}}{B^{3}}\int_{\mathbb{R}^{3}}\frac{K}{\gamma-1}(\sigma(Bx))^{\gamma}\,d(Bx)$
 $=\widetilde{A}^{\gamma}B^{-3}U(\sigma)$
-$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}d^{3}xd^{3}y$ (28)
-$=\widetilde{A}^{2}B^{-5}\int\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}d^{3}(Bx)d^{3}(By)$
+$G\left(\sigma_{\widetilde{A},B},\sigma_{\widetilde{A},B}\right)$ $=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma_{\widetilde{A},B}(x)\sigma_{\widetilde{A},B}(y)}{|x-y|}\,dx\,dy$ (29)
+$=\widetilde{A}^{2}B^{-5}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(Bx)\sigma(By)}{|Bx-By|}\,d(Bx)\,d(By)$
 $=\widetilde{A}^{2}B^{-5}G(\sigma,\sigma)$
-$\int\sigma_{\widetilde{A},B}$ $=\widetilde{A}B^{-3}\int\sigma(Bx)d^{3}(Bx)$ (29)
-$=\widetilde{A}B^{-3}\int\sigma$
+$\int_{\mathbb{R}^{3}}\sigma_{\widetilde{A},B}\,dx$ $=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma(Bx)\,d(Bx)$ (30)
+$=\widetilde{A}B^{-3}\int_{\mathbb{R}^{3}}\sigma\,dx$
 From the scaling relations (28) (29) (30), together with the uniqueness result Theorem 3.6 (i), if we take $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, then we conclude the proof.
-For a complete discussion, one can see [16, Theorem 3.2].
+For a complete discussion, one can see [11, Section 3].
 ∎
@@
 When $\gamma<2$, we still have $\lim\limits_{m\rightarrow 0}A=+\infty$, but $\lim\limits_{m\rightarrow 0}B=\infty$, and the size of $\sigma_{m}$’s support can go to $\infty$ when $m\rightarrow 0$ (flatten out or dispread).
-It coincides with a result mentioned in Lieb and Yau’s paper [40, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$.
+It coincides with a result mentioned in Lieb and Yau’s paper [20, Theorem 5], which says that the radius $R_{m}\rightarrow\infty$ as $m\rightarrow 0$.
 In their paper quantum mechanics (fermions case) is discussed.
@@
 Actually the converse also holds true: all radial solutions of (EL) are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$.
-Furthermore, the minimizer (or radial solution to (EL)) can be uniquely parameterized by its central density, with each central density corresponding to a specific mass, see [16, Proposition 2.48].
+Furthermore, the minimizer (or radial solution to (EL)) can be uniquely parameterized by its central density, with each central density corresponding to a specific mass, see [11, Section 2].
 Thanks to Theorem 3.6, $\left\|\sigma_{m}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is actually the central density of $\sigma_{m}$, thus Theorem 3.9 shows as the central density decreases, the mass also decreases.
 Moreover, it gives the decay rate.
-For more general pressure $P(\rho)$, this decreasing relationship is also established in [16, Corollary 2.41].
+For more general pressure $P(\rho)$, this decreasing relationship is also established in [11, Section 2].
 Let $\sigma_{\widetilde{A},B}=\widetilde{A}\sigma(Bx)$, with $A=m^{-\frac{2}{3\gamma-4}}$, $B=m^{\frac{\gamma-2}{3\gamma-4}}$ as above, actually we also have the following scaling relations:
@@
 $E_{0}^{\prime}\left(\sigma_{\widetilde{A},B}\right)(x)=\widetilde{A}^{\gamma-1}E_{0}^{\prime}(\sigma)\left(Bx\right)=m^{\frac{2-2\gamma}{3\gamma-4}}E_{0}^{\prime}(\sigma)\left({B}x\right)$ (33)
-Therefore, we can further discuss the relationship of variational derivatives (see Lemma 2.6) between densities with different mass (not necessary to be minimizers).
-Proposition 3.14 (Relations of Variational Derivatives Between Densities with Different Mass [16, Proposition 3.4]).
-Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies $E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}$ $E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$, where $\rho(x):=\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$, $\rho$ has mass 1.
+Therefore, we can further discuss the relationship between variational derivatives (see Lemma 2.6) of densities with different mass (not necessarily to be minimizers).
+Proposition 3.12 (Relations Between Variational Derivatives of Densities with Different Mass [11, Section 3]).
+Given $\rho_{m}\in mR(\mathbb{R}^{3})$ with $U(\rho_{m})<\infty$, then $E_{0}\left(\rho_{m}\right)$ is $P_{\infty}\left(\rho_{m}\right)$ differentiable at $\rho_{m}$ and the derivative at $\rho_{m}$ satisfies
+$E_{0}^{\prime}\left(\rho_{m}\right)(x)=A^{-\gamma-1}E_{0}^{\prime}(\rho)\left(\frac{x}{B}\right)=m^{\frac{2\gamma-2}{3\gamma-4}}E_{0}^{\prime}(\rho)\left(\frac{1}{B}x\right)$
+where $\rho(x):=\left(\rho_{m}\right)_{A,B}(x)=A\rho_{m}(Bx)\in R(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $A=m^{-\frac{2}{3\gamma-4}},B=m^{\frac{\gamma-2}{3\gamma-4}}$.
+In particular, $\rho$ has mass 1.
 4 Existence of Constrained Minimizers of $E_{J}(\rho)$ on ${W}_{m}$
@@
 and $\rho_{1-m}$ similarly, then check $\rho(m)=$ $\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$.
-We want to prove the existence of “double constrained” minimizers via direct method [17, Theorem 1.15].
+We want to prove the existence of “double constrained” minimizers via direct method [12, Theorem 1.15].
 Here, “double constrained” means $\rho(m)\in W_{m}$ and $\rho(m)\leq R$ a.e..
-To do that we first show the lower semicontinuity of $E_{J}$ by reformulating arguments in [4] where the authors only consider non-rotating one body case.
-With an abuse of notation, we denote sequences $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and their limit $\rho_{R}$. Note their total mass $\int\rho_{n,R}$} is always one.
+To do that, we first show the lower semicontinuity of $E_{J}$ by reformulating arguments in [4] where the authors only consider non-rotating one-body case.
+With an abuse of notation, we denote sequences in $W_{m,R}$ by $\left\{\rho_{n,R}\right\}$ and their limit by $\rho_{R}$. Note their total mass ${\int_{\mathbb{R}^{3}}\rho_{n,R}}\,dx$ is always one.
 Lemma 4.3 (Lower Semicontinuity of $E_{J}$ in $W_{m,R}$ w.r.t the Weak Topology on $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$).
@@
 Proof.
-If $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$, then for almost every $x\in\mathbb{R}^{3},0\leq\rho_{n,R}\leq$ $R$ and up to subsequence $\rho_{n,R}\rightarrow\rho_{R}$.
+If $\left\{\rho_{n,R}\right\}\subset W_{m,R}$ and $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$, then for almost every $x\in\mathbb{R}^{3},0\leq\rho_{n,R}\leq$ $R$ and up to a subsequence $\rho_{n,R}\rightarrow\rho_{R}$.
 Then $0\leq\rho_{R}\leq R$ almost everywhere.
@@
 Therefore, $W_{m,R}$ is weakly closed.
-To proof the lower semicontinuity, we recall $E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)=\int A(\rho)$- $\frac{\int V_{\rho}\cdot\rho}{2}+T_{J}(\rho)=\int\frac{K(\rho(x))^{\gamma}}{\gamma-1}d^{3}x-\frac{1}{2}\iint\frac{\rho(x)\cdot\rho(y)}{|x-y|}d^{3}xd^{3}y+\frac{J^{2}}{2(I(\rho))}$.
-We first notice $A(\rho)$ is a convex function, so $\int A(\rho)$ is a convex functional.
+To prove the lower semicontinuity, we recall
+$E_{J}(\rho)$ $=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$
+$=\int_{\mathbb{R}^{3}}A(\rho)\,dx-\frac{\int_{\mathbb{R}^{3}}V_{\rho}\cdot\rho\,dx}{2}+T_{J}(\rho)$
+$=\int_{\mathbb{R}^{3}}\frac{K(\rho(x))^{\gamma}}{\gamma-1}\,dx-\frac{1}{2}\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\rho(x)\cdot\rho(y)}{|x-y|}\,dx\,dy+\frac{J^{2}}{2(I(\rho))}$
+We first notice $A(\rho)$ is a convex function, so $\int_{\mathbb{R}^{3}}A(\rho)\,dx$ is a convex functional.
 Then for each constant $k$, $\left\{\rho_{R}\in\right$.
-$\left.W_{m,R}\mid\int A\left(\rho_{R}\right)\leq k\right\}$ is convex.
+$\left.W_{m,R}\mid\int_{\mathbb{R}^{3}}A\left(\rho_{R}\right)\,dx\leq k\right\}$ is convex.
 $\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}$ implies almost everywhere convergence up to a subsequence, and by construction we know $\rho_{n,R}\leq R\cdot 1_{\Omega}$, here $1_{\Omega}$ is the indicator function of compact set $\Omega$.
 By (2) we know $A(0)=0$, $A(\cdot)$ is non-decreasing, and thus $A(\rho_{n,R})\leq A(R)\cdot 1_{\Omega}\in L^{1}$.
-Therefore, if $\int A\left(\rho_{n,R}\right)\leq k$ then $\int A\left(\rho_{R}\right)\leq k$ by dominated convergence theorem or Fatou’s lemma, which means $\left\{\rho_{R}\in W_{m,R}\mid\int A\left(\rho_{R}\right)\leq k\right\}$ is not only convex but also closed.
-Thus, it is also weakly closed and then $\int A(\rho)$ is a weakly lower semicontinuous functional.
-Next, thanks to Proposition A.5, we know $\mathcal{V}:\rho\rightarrow V_{\rho}$ is compact from $L^{\frac{4}{3}}(\Omega)$ to $L^{r}(\Omega)$ for any $1\leq r<12$, thus $\mathcal{V}$ is completely continuous.
+Therefore, if $\int_{\mathbb{R}^{3}}A\left(\rho_{n,R}\right)\,dx\leq k$ then $\int_{\mathbb{R}^{3}}A\left(\rho_{R}\right)\,dx\leq k$ by dominated convergence theorem or Fatou’s lemma, which means $\left\{\rho_{R}\in W_{m,R}\mid\int_{\mathbb{R}^{3}}A\left(\rho_{R}\right)\,dx\leq k\right\}$ is not only convex but also closed.
+Thus, it is also weakly closed and then $\int_{\mathbb{R}^{3}}A(\rho)\,dx$ is a weakly lower semicontinuous functional.
+Next, we know $\mathcal{V}:\rho\rightarrow V_{\rho}$ is compact from $L^{\frac{4}{3}}(\Omega)$ to $L^{r}(\Omega)$ for any $1\leq r<12$ (see, e.g., [11, Appendix A]), thus $\mathcal{V}$ is completely continuous.
 Take $r=11$, then $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}(\Omega)$ implies $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{11}(\Omega)$.
@@
 Since $\rho_{R}$ and $\rho_{n,R}$ vanish outside $\Omega$, we have
-$\left|G\left(\rho_{R},\rho_{R}\right)-G\left(\rho_{n,R},\rho_{n,R}\right)\right|=$ $\left|\int_{\mathbb{R}^{3}}V_{\rho_{R}}\cdot\rho_{R}-\int_{\mathbb{R}^{3}}V_{\rho_{n,R}}\cdot\rho_{n,R}\right|$
-$=$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|+\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}-\int_{\Omega}V_{\rho_{n,R}}\cdot\rho_{n,R}\right|$
-$\leq$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|+\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}$
-The last inequality comes from the Hölder’s inequality.
-Note $V_{\rho_{R}}\in L^{11}(\Omega)\subset L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$, together with $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ we know $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\right|\rightarrow 0$.
+$\left|G\left(\rho_{R},\rho_{R}\right)-G\left(\rho_{n,R},\rho_{n,R}\right)\right|=$ $\left|\int_{\mathbb{R}^{3}}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\mathbb{R}^{3}}V_{\rho_{n,R}}\cdot\rho_{n,R}\,dx\right|$
+$=$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx\right|+\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx-\int_{\Omega}V_{\rho_{n,R}}\cdot\rho_{n,R}\,dx\right|$
+$\leq$ $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx\right|+\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}$
+The last inequality comes from Hölder’s inequality.
+Note $V_{\rho_{R}}\in L^{11}(\Omega)\subset L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$, together with $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ we know $\left|\int_{\Omega}V_{\rho_{R}}\cdot\rho_{R}\,dx-\int_{\Omega}V_{\rho_{R}}\cdot\rho_{n,R}\,dx\right|\rightarrow 0$.
 Note $\rho_{n,R}\rightharpoonup\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ also implies $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq C$, thus $\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|V_{\rho_{R}}-V_{\rho_{n,R}}\right\|_{L^{4}(\Omega)}\rightarrow 0$ due to $V_{\rho_{n,R}}\rightarrow V_{\rho_{R}}$ in $L^{4}(\Omega)$.
 Hence we have $G\left(\rho_{n,R},\rho_{n,R}\right)\rightarrow G\left(\rho_{R},\rho_{R}\right)$.
-Finally, if $\rho\in W_{m,R}$, then $\bar{x}(\rho)=\frac{\int_{\mathbb{R}^{3}}\rho(x)x}{\int_{\mathbb{R}^{3}}\rho}=\int_{\Omega}\rho(x)x$, and $\rho_{n,R}\rightharpoonup\rho_{R}$ implies $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ since $f(x)=x\in L^{\infty}\left(\Omega;\mathbb{R}^{3}\right)\subset\left(L^{\frac{4}{3}}\left(\Omega;\mathbb{R}^{3}\right)\right)^{*}$.
+Finally, if $\rho\in W_{m,R}$, then $\bar{x}(\rho)=\frac{\int_{\mathbb{R}^{3}}\rho(x)x\,dx}{\int_{\mathbb{R}^{3}}\rho\,dx}=\int_{\Omega}\rho(x)x\,dx$, and $\rho_{n,R}\rightharpoonup\rho_{R}$ implies $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ since $f(x)=x\in L^{\infty}\left(\Omega;\mathbb{R}^{3}\right)\subset\left(L^{\frac{4}{3}}\left(\Omega;\mathbb{R}^{3}\right)\right)^{*}$.
 Moreover,
 $\qquad\left|I\left(\rho_{n,R}\right)-I\left(\rho_{R}\right)\right|$
-$=\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
-$\leq\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|+\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
-$\leq\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right\|_{L^{4}(\Omega)}+\left|\int_{\Omega}\left(\rho_{n,R}(x)-\rho_{R}(x)\right)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right|$
+$=\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\,dx-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
+$\leq\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\,dx-\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
+$+\left|\int_{\Omega}\rho_{n,R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx-\int_{\Omega}\rho_{R}(x)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
+$\leq\left\|\rho_{n,R}\right\|_{L^{\frac{4}{3}}(\Omega)}\cdot\left\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\right\|_{L^{4}(\Omega)}+\left|\int_{\Omega}\left(\rho_{n,R}(x)-\rho_{R}(x)\right)r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\,dx\right|$
 By an argument similar to the above, we know $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)$ and $r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ are in $L^{\infty}(\Omega)\subset$ $L^{4}(\Omega)=\left(L^{\frac{4}{3}}(\Omega)\right)^{*}$.
 $\bar{x}\left(\rho_{n,R}\right)\rightarrow\bar{x}\left(\rho_{R}\right)$ means $r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)\rightarrow r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)$ everywhere, thus by dominated convergence theorem we know $\|r^{2}\left(x-\bar{x}\left(\rho_{n,R}\right)\right)-r^{2}\left(x-\bar{x}\left(\rho_{R}\right)\right)\|_{L^{4}(\Omega)}\rightarrow 0$.
-$\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies weak convergence in $L^{\frac{4}{3}}(\Omega)$, $\left\|\rho_{n,R}\right\|_{L^{3}(\Omega)}\leq C$, then we have $I\left(\rho_{n,R}\right)\rightarrow I\left(\rho_{R}\right),\frac{J^{2}}{2\left(I\left(\rho_{n,R}\right)\right)}\rightarrow\frac{J^{2}}{2\left(I\left(\rho_{R}\right)\right)}$.
-Collect the results above we obtain $E_{J}\left(\rho_{R}\right)\leq\liminf\limits_{n\rightarrow\infty}E_{J}\left(\rho_{n,R}\right)$.
-∎
+$\rho_{n,R}\rightarrow\rho_{R}$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies weak convergence in $L^{\frac{4}{3}}(\Omega)$, $\left\|\rho_{n,R}\right\|_{L^{3}(\Omega)}\leq C$, then we have $I\left(\rho_{n,R}\right)\rightarrow I\left(\rho_{R}\right)$, and then $\frac{J^{2}}{2\left(I\left(\rho_{n,R}\right)\right)}\rightarrow\frac{J^{2}}{2\left(I\left(\rho_{R}\right)\right)}$.
+Collecting the results above, we obtain $E_{J}\left(\rho_{R}\right)\leq\liminf\limits_{n\rightarrow\infty}E_{J}\left(\rho_{n,R}\right)$. ∎
 Remark 4.4.
@@
 Theorem 4.5 (Existence of “Double Constrained” Minimizers).
-Fix $J>0$, give $m>0,R>\frac{384}{\pi\eta^{3}}$, there is a $\rho_{R}\in W_{m,R}$ which minimizes $E_{J}(\rho)$ on $W_{m,R}$.
+Fix $J>0$, given $m>0,R>\frac{384}{\pi\eta^{3}}$, there is a $\rho_{R}\in W_{m,R}$ which minimizes $E_{J}(\rho)$ on $W_{m,R}$.
 Any such $\rho_{R}$ is continuous on $\Omega=\Omega_{m}\cup\Omega_{1-m}$.
@@
 Given a minimizing sequence $\left\{\rho_{n,R}\right\}$ with $\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$.
-By Interpolation Inequality Proposition A.2 we know $W_{m,R}$ is bounded in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$.
-Since $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ is reflexive, $\exists\rho_{R}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ such that $\rho_{n,R}\rightharpoonup\rho_{R}$ up to subsequence [8, Theorem 3.18].
-By Lemma 4.3 $W_{m,R}$ is weakly compact in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right),\rho_{R}\in W_{m,R}$, and $\inf\limits_{\rho\in W_{m,R}}E(\rho)\leq E\left(\rho_{R}\right)\leq\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$, which means $\rho_{R}$ is a minimizer.
-A slight modification of the proof of [4, Lemma 2], [15, Proposition 5.8] and [16, Lemma 2.21], which means we need to consider the perturbations have compact supports in $\Omega$, shows that (33) and (34) hold for almost all points which are mentioned in (33) (34) with suitable $\lambda_{m,R}$ or $\lambda_{1-m,R}$.
-The continuity of $\rho_{R}$ on $\Omega$ then follows as in the proof of Theorem 3.6 (see e.g.
-[15, Theorem 2.25] and [16, Remark 2.15]), where the idea is making use of (33) (34) (35) and improving the regularity of potential $V_{\rho_{R}}$ and density $\rho_{R}$ in turns (bootstrap).
+By Interpolation Inequality [6, Section 4.2] we know $W_{m,R}$ is bounded in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$.
+Since $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ is reflexive, $\exists\rho_{R}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ such that $\rho_{n,R}\rightharpoonup\rho_{R}$ up to subsequence [6, Theorem 3.18].
+By Lemma 4.3 $W_{m,R}$ is weakly compact in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right),\rho_{R}\in W_{m,R}$, and
+$\inf\limits_{\rho\in W_{m,R}}E(\rho)\leq E\left(\rho_{R}\right)\leq\lim\limits_{n\rightarrow\infty}E\left(\rho_{n,R}\right)=\inf\limits_{\rho\in W_{m,R}}E(\rho)$
+It means $\rho_{R}$ is a minimizer.
+A slight modification of the proof of [4, Lemma 2], [10, Section 5] and [11, Section 2], which means we need to consider the perturbations have compact supports in $\Omega$, shows that (34) and (35) hold for almost all points which are mentioned in (34) (35) with suitable $\lambda_{m,R}$ or $\lambda_{1-m,R}$.
+The continuity of $\rho_{R}$ on $\Omega$ then follows as in the proof of Theorem 3.6 (see, e.g., subsection 5.2 of this paper, [4], [10, Section 2] and [11, Section 2]), where the idea is making use of (34) (35) (36) and improving the regularity of potential $V_{\rho_{R}}$ and density $\rho_{R}$ in turns (bootstrap method).
 Then we know (34) and (35) hold indeed for all points we considered.
@@
 Remark 4.6.
-Another type of bootstrap, commonly used in fluid dynamics and general relativity, refers to bootstrapping with respect to time. This differs from the regularity bootstrap discussed in our paper.
+Since we consider potential and density alternatively, we call such method the bootstrap method.
+However, another type of bootstrap method, commonly used in fluid dynamics and general relativity, refers to bootstrapping with respect to time.
+It differs from the regularity bootstrap discussed in this paper.
 We modify the arguments in [4, Section 6] and apply again bootstrap method to show that there is a $\widetilde{R}$ such that the “double constrained” (“DC” in short) minimizer $\rho_{R}$ remains in $W_{m,\widetilde{R}}$ for all $R\geq\widetilde{R}$.
@@
 We have $\forall R\in\left[R_{0},\infty\right),\rho_{R_{0}}\in W_{m,R}$, then $E\left(\rho_{R}\right)\leq E\left(\rho_{R_{0}}\right)<\infty$.
-Then, one can adapt the arguments from the non-rotating case—see, for instance, [16, Lemma 2.21] or [4, Section 6]—to show that the $L^{\frac{4}{3}}$ norm of $\rho_{R}$ is actually bounded by a constant depends only on $E\left(\rho_{R_{0}}\right)$.
+Then, one can adapt the arguments from the non-rotating case—see, for instance, [11, Section 2] or [4, Section 6]—to show that the $L^{\frac{4}{3}}$ norm of $\rho_{R}$ is actually bounded by a constant that depends only on $E\left(\rho_{R_{0}}\right)$.
 ∎
 Deducing the uniform boundedness in $L^{\infty}$ from that in $L^{\frac{4}{3}}$ can be accomplished in several ways.
-For instance, one may refer to the methods of Li in [38, Proposition 1.4], as McCann mentioned in [46, Section 6].
-Li’s method [38] can be applied to obtain uniform boundedness in $L^{\infty}$ for $\gamma>\frac{4}{3}$ by choosing an appropriate test function, although it does not first verify whether Lagrange multiplier is negative — Li verified this after the $L^{\infty}$ bound was obtained in [38].
+For instance, one may refer to the methods of Li in [18, Proposition 1.4], as McCann mentioned in [24, Section 6].
+Li’s method [18] can be applied to obtain uniform boundedness in $L^{\infty}$ for $\gamma>\frac{4}{3}$ by choosing an appropriate test function, although it is not first verified that Lagrange multiplier is negative — Li verified this after the $L^{\infty}$ bound was obtained in [18].
 Here, we present an alternative approach, which is inspired by the method in [4].
 Both methods rely on the result that velocity $v^{2}(x)=\frac{J^{2}r^{2}(x-\bar{x}(\rho))}{I^{2}(\rho)}$ is sufficiently small.
-Under assumption $\gamma>\frac{3}{2}$, we first prove that Lagrange multipliers are negative, using the idea from [4, Lemma 5] and [38, Proposition 2.5].
+Under assumption $\gamma>\frac{3}{2}$, we first prove that Lagrange multipliers are negative, using the idea from [4, Lemma 5] and [18, Proposition 2.5].
 Lemma 4.8 (Negavie Lagrange-multipliers for DC Minimizers).
@@
 Proof.
-For $\lambda_{m,R}$, since the volume (measure) of $\Omega_{m}$ is ${Vol}:=\frac{4}{3}\pi\left(\frac{\eta}{4}\right)^{3}=$ $\frac{\pi J^{6}}{48m^{6}(1-m)^{6}}$, while the mass in $\Omega_{m}$ is $\int_{\Omega_{m}}\rho_{R}=m$.
-Claim: $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$.
-In fact, if not, then $\rho_{R}(x)-\frac{m}{Vol}>0$ almost everywhere in $\Omega_{m}$, while $\int_{\Omega_{m}}\left(\rho_{R}(x)-\frac{m}{Vol}\right)=0$ implies $\rho_{R}(x)-\frac{m}{Vol}=0$ almost everywhere in $\Omega_{m}$ [51, Section 4.3].
+For $\lambda_{m,R}$, since the volume (measure) of $\Omega_{m}$ is ${Vol}:=\frac{4}{3}\pi\left(\frac{\eta}{4}\right)^{3}=$ $\frac{\pi J^{6}}{48m^{6}(1-m)^{6}}$, while the mass in $\Omega_{m}$ is $\int_{\Omega_{m}}\rho_{R}\,dx=m$.
+Claim: $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$, where $\mu$ denotes the Lebesgue measure.
+In fact, if not, then $\rho_{R}(x)-\frac{m}{Vol}>0$ almost everywhere in $\Omega_{m}$, while $\int_{\Omega_{m}}\left(\rho_{R}(x)-\frac{m}{Vol}\right)\,dx=0$ implies $\rho_{R}(x)-\frac{m}{Vol}=0$ almost everywhere in $\Omega_{m}$ [28, Section 4.3].
 Since the measure of $\Omega_{m}$ is not 0, it makes a contradiction.
 Now we have $\mu\left(\left\{x\in\Omega_{m}\left\lvert\,\rho_{R}(x)\leq\frac{m}{Vol}\right.\right\}\right)>0$, which means there is at least a point $x_{0}$ in $\Omega_{m}$ such that $\rho_{R}\left(x_{0}\right)\leq\frac{m}{Vol}$ and (34) holds true.
-Notice ${Vol}\sim m^{-6}$ as $m\ll 1$, i.e.
-$\exists C_{1}>C_{2}>0$, such that if $m$ is sufficiently small, then $C_{2}m^{-6}\leq{Vol}\leq C_{1}m^{-6}$.
+Notice ${Vol}\sim m^{-6}$ as $m\ll 1$, i.e., $\exists C_{1}>C_{2}>0$, such that if $m$ is sufficiently small, then $C_{2}m^{-6}\leq{Vol}\leq C_{1}m^{-6}$.
 Therefore, $\frac{m}{Vol}\sim$ $m^{7},A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\left(x_{0}\right)\stackrel{{\scriptstyle\gamma>\frac{3}{2}}}{{=}}O\left(m^{\frac{7}{3}}\right)$.
-By construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\forall y\in\Omega_{1-m},\left|x_{0}-y\right|\leq\operatorname{diam}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}},V_{\rho_{R}}\left(x_{0}\right)\geq$ $\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{\left|x_{0}-y\right|}d^{3}y\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)$ when $0<m<m_{1}$ for some $m_{1}$.
-Let $R>\frac{m_{1}}{{Vol}\left(m_{1}\right)}$, then $\rho_{R}(x_{0})$ that satisfies the above relation exists, and (33) holds true thus
+By construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\forall y\in\Omega_{1-m}$,
+$\left|x_{0}-y\right|\leq\operatorname{diam}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}}$
+$V_{\rho_{R}}\left(x_{0}\right)\geq\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{\left|x_{0}-y\right|}\,dy\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)$
+when $0<m<m_{1}$ for some $m_{1}$. Let $R>\frac{m_{1}}{{Vol}\left(m_{1}\right)}$, then $\rho_{R}(x_{0})$ that satisfies the above relation exists, and (34) holds true thus
 $\lambda_{m,R}$ $\leq E_{J}^{\prime}\left(\rho_{R}\right)\left(x_{0}\right)$
@@
 $\leq A^{\prime}\left(\rho_{R}\left(x_{0}\right)\right)-V_{\rho_{R}}\left(x_{0}\right)<0$
+Similarly, $\exists m_{2}>0,\forall 0<m<m_{2},\exists\widetilde{x_{0}}\in\Omega_{1-m}$, such that $\rho_{R}\left(\widetilde{x_{0}}\right)\leq\frac{1-m}{Vol}$, and since $\gamma>\frac{3}{2}$,
 […]
@@
 Lemma 4.9 (Moment of Inertia Increasing Rate).
-Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall 0<m<M$ and $\forall R>0$, $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $I\left(\rho_{1-m}+\rho_{m}\right)>{Cm}^{-3}$.
+Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall 0<m<M$ and $\forall R>0$, $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $I\left(\rho_{1-m}+\rho_{m}\right)>{Cm}^{-3}$.
 Proof.
 $\rho(m)=\rho_{1-m}+\rho_{m}\in{W}_{m,R}$.
-The centers of mass $\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m},\bar{x}\left(\rho_{m}\right)\in\Omega_{m}$, as well as their projections onto $z=0$, are separated by at least $\frac{\eta}{2},\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
+The centers of mass $\bar{x}\left(\rho_{1-m}\right)\in\Omega_{1-m}$, $\bar{x}\left(\rho_{m}\right)\in\Omega_{m}$, as well as their projections onto $z=0$, are separated by at least $\frac{\eta}{2}$, where $\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
 Due to Lemma 2.4, we have
-$I(\rho(m))=\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+I\left(\rho_{1-m}\right)+I\left(\rho_{m}\right)\geq\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)\geq\frac{J^{4}}{4\mu^{3}}$
-Since $\mu=m(1-m)$, we know $\exists C>0$, such that $I(\rho(m))>Cm^{-3}$ when $m\rightarrow 0$. ∎
+$I(\rho(m))=\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+I\left(\rho_{1-m}\right)+I\left(\rho_{m}\right)\geq\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)\geq\frac{J^{4}}{4\mu_{r}^{3}}$
+Since $\mu_{r}=m(1-m)$, we know $\exists C>0$, such that $I(\rho(m))>Cm^{-3}$ when $m\rightarrow 0$. ∎
 Corollary 4.10 (Velocity Decreasing Rate).
-Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $0<m<M$, $R>0$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $v(x,m)=\frac{Jr(x-\bar{x}(\rho(m)))}{I(\rho(m))}<Cm$.
+Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $0<m<M$, $R>0$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is in $W_{m,R}$, we have $v(x,m)=\frac{Jr(x-\bar{x}(\rho(m)))}{I(\rho(m))}<Cm$.
 Proof.
@@
 Thanks to Corollary 4.10, when $m$ is sufficiently small, $\exists C_{1}>0$, such that $\forall x\in\Omega=\Omega_{m}\cup\Omega_{1-m}$, $v^{2}(x)\leq C_{1}m^{2}$.
-Moreover, similarly as above if $x\in\Omega$, we have $\forall y\in\Omega_{1-m},|x-y|\leq\operatorname{diam}\left(\Omega_{m}\cup\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}},V_{\rho_{R}}(x)\geq$ $\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{|x-y|}d^{3}y\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>C_{2}m^{2}$ for some $C_{2}>0$ when $m$ is small enough.
+Moreover, similarly as above if $x\in\Omega$, we have $\forall y\in\Omega_{1-m}$, $|x-y|\leq\operatorname{diam}\left(\Omega_{m}\cup\Omega_{1-m}\right)=\frac{3\eta}{2}=\frac{3J^{2}}{2m^{2}(1-m)^{2}}$, and
+$V_{\rho_{R}}(x)\geq\int_{\Omega_{1-m}}\frac{\rho_{R}(y)}{|x-y|}\,dy\geq\frac{2m^{2}(1-m)^{3}}{3J^{2}}>C_{2}m^{2}$
+for some $C_{2}>0$ when $m$ is small enough.
 Therefore, we know $\exists\widetilde{C}>0$, such that $\forall x\in\Omega$, $v^{2}(x)\leq\widetilde{C}V_{\rho_{R}}(x)$.
-By (34), and Lemma 4.8 (Lagrange-multipliers are less then 0), when $R>R_{1}$, we have if $\rho_{R}(x)>0$, then $A^{\prime}\left(\rho_{R}(x)\right)\leq V_{\rho}(x)+v^{2}(x)\leq CV_{\rho}(x)$.
-If $\rho_{R}(x)=0$, the $0=$ $A^{\prime}\left(\rho_{R}(x)\right)\leq CV_{\rho_{R}}(x)$ is obvious.
+By (35), and Lemma 4.8 (Lagrange-multipliers are less than 0), when $R>R_{1}$, we have if $\rho_{R}(x)>0$, then $A^{\prime}\left(\rho_{R}(x)\right)\leq V_{\rho}(x)+v^{2}(x)\leq CV_{\rho}(x)$.
+If $\rho_{R}(x)=0$, then $0=$ $A^{\prime}\left(\rho_{R}(x)\right)\leq CV_{\rho_{R}}(x)$ is obvious.
 Thanks to Lemma 4.7, we know if $R>R_{0}+R_{1},\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}}\leq k_{0}$.
 By HardyLittlewood-Sobolev Inequality (Proposition A.1), with an abuse of notation $C$, we have $\left\|V_{\rho_{R}}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}\leq$ $C\left\|\rho_{R}\right\|_{L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\leq C$.
-$A^{\prime}\left(\rho_{R}\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\leq CV_{\rho}$ implies $\left\|\rho_{R}\right\|_{L^{12(\gamma-1)}}=$ $\left(\int\rho_{R}{}^{12(\gamma-1)}(x)d^{3}x\right)^{\frac{1}{12(\gamma-1)}}\leq C\left(\int V_{\rho_{R}}^{12}(x)d^{3}x\right)^{\frac{1}{12(\gamma-1)}}=C\left\|V_{\rho_{R}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C$.
-Since we know $\gamma>\frac{3}{2}>\frac{4}{3}$ and $\Omega$ is bounded, by interpolation inequality (Proposition A.2) we have $\left\|\rho_{R}\right\|_{L^{6}}\leq C$.
+And then $A^{\prime}\left(\rho_{R}\right)=\frac{K\gamma}{\gamma-1}\rho_{R}^{\gamma-1}\leq CV_{\rho}$ implies
+$\left\|\rho_{R}\right\|_{L^{12(\gamma-1)}}=\left(\int_{\mathbb{R}^{3}}\rho_{R}{}^{12(\gamma-1)}(x)\,dx\right)^{\frac{1}{12(\gamma-1)}}\leq C\left(\int_{\mathbb{R}^{3}}V_{\rho_{R}}^{12}(x)\,dx\right)^{\frac{1}{12(\gamma-1)}}=C\left\|V_{\rho_{R}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C$
+Since we know $\gamma>\frac{3}{2}>\frac{4}{3}$ and $\Omega$ is bounded, by interpolation inequality [6, Section 4.2] we have $\left\|\rho_{R}\right\|_{L^{6}}\leq C$.
 Since we also know $\left\|\rho_{R}\right\|_{L^{1}\left(\mathbb{R}^{3}\right)}=1$ are bounded, thanks to Proposition A.2, we have $V_{\rho_{R}}$ is continuous and $\left\|V_{\rho_{R}}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is bounded uniformly.
@@
 Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, the $L^{\infty}$ norm of $\sigma$ may be unbounded, hence in order to show the “double constrained minimizer” $\rho_{R}$ is indeed a constrained minimizer for $R$ large enough, we need to prove that energy of $\sigma$ still exceeds energy of $\rho_{R}$, which was not addressed in McCann’s work [24].
-We provides a supplementary proof for it by adapting the arguments in the single-star case [4, Section 6].
+We provide a supplementary proof for it by adapting the arguments in the single-star case [4, Section 6].
 Theorem 4.14 (Existence of the Constrained Minimizers).
@@
 The facts that $\rho$ is continuous on $\Omega$ and $\rho<R(m)$ can be handled similarly to the case for “double constrained” minimizer, see Theorem 4.5, Theorem 4.11, and Remark 4.13.
-To prove the existence of constrained minimizer, let $k_{1}$ is given in Theorem 4.11, we have $E_{J}\left(\rho_{R}\right)=E_{J}\left(\rho_{k_{1}}\right)$ for $R\geq k_{1}$ due to Theorem 4.11.
+To prove the existence of constrained minimizer, let $k_{1}$ be given in Theorem 4.11, we have $E_{J}\left(\rho_{R}\right)=E_{J}\left(\rho_{k_{1}}\right)$ for $R\geq k_{1}$ due to Theorem 4.11.
 We will prove $\rho_{k_{1}}$ is a constrained minimizer.
@@
 Given any $\sigma\in W_{m}\subset{L^{1}(\mathbb{R}^{3})\cap L^{\frac{4}{3}}(\mathbb{R}^{3})}$, due to Remark 2.2 we know $G(\sigma,\sigma)<\infty$.
-For $\sigma$ with $U(\sigma)=\int A(\sigma(x))d^{3}x=\infty$, we know $E_{J}(\rho_{k_{1}})<E_{J}(\sigma)=\infty$.
+For $\sigma$ with $U(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma(x))\,dx=\infty$, we know $E_{J}(\rho_{k_{1}})<E_{J}(\sigma)=\infty$.
 For $\sigma$ with $U(\sigma)<\infty$, let
 […]
-When $R$ is large enough, we can find a $\tau\in\widetilde{P_{R}}(\sigma)$, with $\tau\geq 0$ and $\int_{\Omega_{m}}\tau=\int_{\Omega_{1-m}}\tau=1$.
+When $R$ is large enough, we can find a $\tau\in\widetilde{P_{R}}(\sigma)$, with $\tau\geq 0$ and $\int_{\Omega_{m}}\tau\,dx=\int_{\Omega_{1-m}}\tau\,dx=1$.
 (See e.g.
-[15, Remark 4.10].) Let $\sigma_{s}^{[1]}(x),\sigma_{s}^{[2]}(x)$ and $\sigma_{s}$ be
+[10, Section 4].) Let $\sigma_{s}^{[1]}(x),\sigma_{s}^{[2]}(x)$ and $\sigma_{s}$ be
 $\sigma_{s}^{[1]}(x)=\left\{\begin{array}[]{ll}\sigma(x),&\text{ if }x\in\Omega\text{ and }\sigma(x)<\frac{1}{2}s\\ 0,&\text{ otherwise }\end{array}\right$.
-$\sigma_{s}^{[2]}(x)=\left\{\begin{array}[]{ll}\left(m-\int_{\Omega_{m}}\sigma_{s}^{[1]}\right)\tau,&\text{ if }x\in\Omega_{m}\\ \left((1-m)-\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\right)\tau,&\text{ if }x\in\Omega_{1-m}\\ 0,&\text{ otherwise }\end{array}\right$.
+$\sigma_{s}^{[2]}(x)=\left\{\begin{array}[]{ll}\left(m-\int_{\Omega_{m}}\sigma_{s}^{[1]}\,dx\right)\tau,&\text{ if }x\in\Omega_{m}\\ \left((1-m)-\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\,dx\right)\tau,&\text{ if }x\in\Omega_{1-m}\\ 0,&\text{ otherwise }\end{array}\right$.
 $\sigma_{s}=\sigma_{s}^{1}+\sigma_{s}^{2}$
-By construction we have $\int_{\Omega_{m}}\sigma_{s}=m$ and $\int_{\Omega_{1-m}}\sigma_{s}=1-m$, then $\sigma_{s}\in W_{m}$.
-Similar again to [15, Remark 4.10], we have $\int_{\Omega_{m}}\sigma_{s}^{[1]}\rightarrow\int_{\Omega_{m}}\sigma=m$, $\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\rightarrow\int_{\Omega_{1-m}}\sigma=1-m$, therefore $\sigma_{s}^{[2]}\rightarrow 0$ almost everywhere in $\mathbb{R}^{3}$ as $s\rightarrow\infty$ due to $\tau\in L^{\infty}$.
+By construction we have $\int_{\Omega_{m}}\sigma_{s}\,dx=m$ and $\int_{\Omega_{1-m}}\sigma_{s}\,dx=1-m$, then $\sigma_{s}\in W_{m}$.
+Similar again to [10, Section 4], we have $\int_{\Omega_{m}}\sigma_{s}^{[1]}\,dx\rightarrow\int_{\Omega_{m}}\sigma\,dx=m$, $\int_{\Omega_{1-m}}\sigma_{s}^{[1]}\,dx\rightarrow\int_{\Omega_{1-m}}\sigma\,dx=1-m$, therefore $\sigma_{s}^{[2]}\rightarrow 0$ almost everywhere in $\mathbb{R}^{3}$ as $s\rightarrow\infty$ due to $\tau\in L^{\infty}$.
 $\sigma\in W_{m}$ implies $\sigma$ is finite almost everywhere, thus $\sigma_{s}^{[1]}$ increases to $\sigma$ almost everywhere as $s\rightarrow\infty$.
@@
 $A(\sigma+\|\tau\|_{L^{\infty}})=A(\frac{2\sigma+2\|\tau\|_{L^{\infty}}}{2})\leq\frac{A(2\sigma)+A(2\|\tau\|_{L^{\infty}})}{2}$
-Note $U(\sigma)=\int A(\sigma(x))d^{3}x<\infty$ implies $A(2\sigma)\in L^{1}(\mathbb{R}^{3})$, and $\Omega_{>0}$ is bounded implies $A(2\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$.
+Note $U(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma(x))\,dx<\infty$ implies $A(2\sigma)\in L^{1}(\mathbb{R}^{3})$, and $\Omega_{>0}$ is bounded implies $A(2\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$.
 Therefore, we know $A(\sigma+\|\tau\|_{L^{\infty}})\in L^{1}(\Omega_{>0})$.
-We can then split $\mathbb{R}^{3}$ into $\mathbb{R}^{3}\setminus\Omega_{>0}$ and $\Omega_{>0}$, and show $\int_{\mathbb{R}^{3}}A\left(\sigma_{s}\right)\rightarrow\int_{\mathbb{R}^{3}}A(\sigma)$ as $s\rightarrow\infty$ by the monotone convergence theorem and dominated convergence theorem.
-Since $\Omega$ is bounded, by Hölder’s inequality we know $\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq C\|\tau\|_{L^{\infty}(\Omega)}$, then $\left\|\sigma_{s}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\left\|\sigma_{s}^{[1]}\right\|_{L^{\frac{4}{3}}(\Omega)}+\left\|\sigma_{s}^{[2]}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq$ $\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+C\|\tau\|_{L^{\infty}(\Omega)}$.
+We can then split $\mathbb{R}^{3}$ into $\mathbb{R}^{3}\setminus\Omega_{>0}$ and $\Omega_{>0}$, and show $\int_{\mathbb{R}^{3}}A\left(\sigma_{s}\right)\,dx\rightarrow\int_{\mathbb{R}^{3}}A(\sigma)\,dx$ as $s\rightarrow\infty$ by the monotone convergence theorem and dominated convergence theorem.
+Since $\Omega$ is bounded, by Hölder’s inequality we know $\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq C\|\tau\|_{L^{\infty}(\Omega)}$, then
+$\left\|\sigma_{s}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\left\|\sigma_{s}^{[1]}\right\|_{L^{\frac{4}{3}}(\Omega)}+\left\|\sigma_{s}^{[2]}\right\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+\|\tau\|_{L^{\frac{4}{3}}(\Omega)}\leq\|\sigma\|_{L^{\frac{4}{3}}(\Omega)}+C\|\tau\|_{L^{\infty}(\Omega)}$
 It means $\left\{\sigma_{s}\right\}$ is a bounded set in $L^{\frac{4}{3}}(\Omega)$.
-Since $L^{\frac{4}{3}}(\Omega)$ is reflexive, thanks to [8, Theorem 3.18], there is a sequence $\left\{\sigma_{s_{n}}\right\}$ which converges weakly to a weak limit $\widetilde{\sigma}$.
+Since $L^{\frac{4}{3}}(\Omega)$ is reflexive, thanks to [6, Theorem 3.18], there is a sequence $\left\{\sigma_{s_{n}}\right\}$ which converges weakly to a weak limit $\widetilde{\sigma}$.
 Since we know $\sigma_{s_{n}}\rightarrow\sigma$ a.e., by [31], we have $\widetilde{\sigma}=\sigma$.
 That is, $\sigma_{s_{n}}\rightharpoonup\sigma$ in $L^{\frac{4}{3}}(\Omega)$, therefore, we can apply what we did in the proof of Lemma 4.3, and obtain $G\left(\sigma_{s_{n}},\sigma_{s_{n}}\right)\rightarrow G(\sigma,\sigma)$, $T_{J}\left(\sigma_{s_{n}}\right)\rightarrow T_{J}(\sigma)$.
-Collect the results above, we have $E_{J}(\sigma)=\lim\limits_{n\rightarrow\infty}E_{J}\left(\sigma_{s_{n}}\right)$.
+Collecting the results above, we have $E_{J}(\sigma)=\lim\limits_{n\rightarrow\infty}E_{J}\left(\sigma_{s_{n}}\right)$.
 By the same arguments, we know for any sequence of $\left\{\sigma_{s_{m}}\right\}$, we can always pick a subsequence $\left\{\sigma_{s_{m_{k}}}\right\}$ which also satisfies $E_{J}(\sigma)=\lim\limits_{k\rightarrow\infty}E_{J}\left(\sigma_{s_{m_{k}}}\right)$.
@@
 Consequently, the same argument is applicable to more general pressure.
-For the polytropic case studied in this paper, one may also use the following estimate: $A(\sigma+\|\tau\|_{L^{\infty}})=\frac{K}{\gamma-1}(\sigma+\|\tau\|_{L^{\infty}})^{\gamma}\leq\frac{K}{\gamma-1}(2\max\{\sigma,\|\tau\|_{L^{\infty}}\})^{\gamma}=\frac{K}{\gamma-1}(2^{\gamma}\max\{\sigma^{\gamma},\|\tau\|_{L^{\infty}}^{\gamma}\})\leq\frac{K2^{\gamma}}{\gamma-1}(\sigma^{\gamma}+\|\tau\|_{L^{\infty}}^{\gamma})=2^{\gamma}(A(\sigma)+A(\|\tau\|_{L^{\infty}}))$.
+For the polytropic case studied in this paper, one may also use the following estimate:
+$A(\sigma+\|\tau\|_{L^{\infty}})$ $=\frac{K}{\gamma-1}(\sigma+\|\tau\|_{L^{\infty}})^{\gamma}$
+$\leq\frac{K}{\gamma-1}(2\max\{\sigma,\|\tau\|_{L^{\infty}}\})^{\gamma}$
+$=\frac{K2^{\gamma}}{\gamma-1}\max\{\sigma^{\gamma},\|\tau\|_{L^{\infty}}^{\gamma}\}$
+$\leq\frac{K2^{\gamma}}{\gamma-1}(\sigma^{\gamma}+\|\tau\|_{L^{\infty}}^{\gamma})$
+$=2^{\gamma}A(\sigma)+2^{\gamma}A(\|\tau\|_{L^{\infty}})$
 Remark 4.16.
 Following the statement in Section 5, one can show that the constant $R(m)$ is actually a uniform constant for all $m\in(0,m_{0})$ for some small $m_{0}$.
-Remark 4.17.
-Note that the continuity of $\rho$ has so far been established only on $\Omega$, not on the entire $\mathbb{R}^{3}$.
-This would suggest that there might be a sense of discontinuous density transition around the boundary for stars or planets.
-However, by additionally showing that the support lies in the interior, we will rule out this scenario and further extend the result to continuity on $\mathbb{R}^{3}$.
-Moreover, we have established the existence of a constrained minimizer, one can verify that such a minimizer satisfies the Euler–Lagrange equations (EL) on $\Omega$.
-However, to conclude that the reduced Euler–Poisson equations (EP’) hold in $\mathbb{R}^{3}$, we need to show the existence of $\nabla P(\rho)$ in $\mathbb{R}^{3}$, particularly on $\partial\{\rho>0\}$.
-It turns out $W^{\infty}$ local energy minimizers can do this job as we mention in Remark 3.4, and we have Theorem 3.1.
-Thanks to Remark 2.11, it suffices to show that the minimizer can be placed inside the interior of $\Omega$, which will be discussed in the following sections.
 5 Bound of the Supports of Density Functions
-In this section we estimate a uniform bound for the supports of the density functions, under small mass ratio assumption which is desirable physically in star-planet systems.
+In this section we estimate a uniform bound for the supports of the density functions, under small mass ratio assumption which is physically desirable in star-planet systems.
 The analysis focuses on the planet’s density.
-For stars densities, the estimates are similar or even simpler, and can be adapted directly from the arguments in [46] and does not require the scaling method in subsection 5.1 and subsection 5.2.
+For the star’s density, the estimates are similar or even simpler, and can be adapted directly from the arguments in [24] and do not require the scaling method in subsection 5.1 and subsection 5.2.
 With these uniform bounds we will show in Section 6 that the minimizers can be located in the interior of the domains we define.
@@
 $\quad E_{J}\left({\rho_{1-m}+\rho_{m}}\right)$ (38)
-$={\int{\frac{K{\left({\rho_{1-m}(x)+\rho_{m}(x)}\right)}^{\gamma}}{\gamma-1}d^{3}x}}-\frac{1}{2}{\iint{\frac{\rho(m)(x)\cdot\rho(m)(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}+\frac{J^{2}}{2\left({I\left({\rho_{1-m}+\rho_{m}}\right)}\right)}$
+$={\int_{\mathbb{R}^{3}}{\frac{K{\left({\rho_{1-m}(x)+\rho_{m}(x)}\right)}^{\gamma}}{\gamma-1}\,dx}}-\frac{1}{2}{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho(m)(x)\cdot\rho(m)(y)}{\left|{x-y}\right|}\,dx\,dy}}+\frac{J^{2}}{2\left({I\left({\rho_{1-m}+\rho_{m}}\right)}\right)}$
 $=E_{0}\left(\rho_{1-m}\right)+E_{0}\left(\rho_{m}\right)-G\left({\rho_{1-m},\rho_{m}}\right)+T_{J}\left({\rho_{1-m}+\rho_{m}}\right)$
@@
 $>E_{0}\left(\rho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)-G\left({\rho_{1-m},\rho_{m}}\right)$
-Notice since $\rho_{1-m}\in\Omega_{1-m}$, $\rho_{m}\in\Omega_{m}$, which distance is at least $\frac{\eta}{2}=\frac{J^{2}}{2m^{2}(1-m)^{2}}$ (24), we obtain $G\left({\rho_{1-m},\rho_{m}}\right)\leq\frac{2m\left({1-m}\right)}{\eta}<Cm^{3}$ for some $C>0$.
+Notice since $\rho_{1-m}\in\Omega_{1-m}$, $\rho_{m}\in\Omega_{m}$, whose distance is at least $\frac{\eta}{2}=\frac{J^{2}}{2m^{2}(1-m)^{2}}$ (25), we obtain $G\left({\rho_{1-m},\rho_{m}}\right)\leq\frac{2m\left({1-m}\right)}{\eta}<Cm^{3}$ for some $C>0$.
 […]
 Lemma 5.3 (Moment of Inertia Increasing Rate as Mass Decreases).
-Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $I\left(\rho_{1-m}+\rho_{m}\right)>Cm^{-3}$.
+Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $I\left(\rho_{1-m}+\rho_{m}\right)>Cm^{-3}$.
 Proof.
-The proof is similar as in the proof of Lemma 4.9. ∎
+The proof is similar to the proof of Lemma 4.9. ∎
 Corollary 5.4 (Kinetic Energy Decreasing Rate as Mass Decreases).
-Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $T_{J}(\rho(m))<Cm^{3}$.
+Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $T_{J}(\rho(m))<Cm^{3}$.
 Proof.
@@
 Corollary 5.5 (Velocity Decreasing Rate as Mass Decreases).
-Fix $J>0$, there exist a $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $v(x,m)<Cm$.
+Fix $J>0$, there exist an $M>0$ and a $C>0$, such that $\forall x\in\Omega_{m}\cup\Omega_{1-m}$, $m<M$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer, we have $v(x,m)<Cm$.
 Proof.
-The proof is similar as in the proof of Lemma 4.10. ∎
+The proof is similar to the proof of Lemma 4.10. ∎
 Remark 5.6.
@@
 Similarly, $\sigma_{1-m}=\frac{1}{\widetilde{A}}\sigma\left(\frac{1}{\widetilde{B}}x\right)$, where $\widetilde{A}=(1-m)^{-\frac{2}{3\gamma-4}}$, $\widetilde{B}=(1-m)^{\frac{\gamma-2}{3\gamma-4}}$.
-Moreover, $\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\varrho_{m}\right)$, $\frac{{\widetilde{B}}^{3}}{{\widetilde{A}}^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{1-m}\right)=E_{0}\left(\varrho_{1-m}\right)$.
+Moreover, $\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\varrho_{m}\right)$, and $\frac{{\widetilde{B}}^{3}}{{\widetilde{A}}^{\gamma}}E_{0}(\sigma)=E_{0}\left(\sigma_{1-m}\right)=E_{0}\left(\varrho_{1-m}\right)$.
 Proposition 5.8 (Energy Converges to Non-Rotating Minimum).
@@
 $E_{J}\left({\varrho_{m}+\varrho_{1-m}}\right)-E_{0}\left(\varrho_{1-m}\right)-E_{0}\left(\varrho_{m}\right)=-G\left({\varrho_{1-m},\varrho_{m}}\right)+T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)$ (41)
-Again, by Theorem 3.6, $\varrho_{1-m}$ and $\varrho_{m}$ have some nice properties, in particular they are spherically symmetry and radially decreasing after translations.
-Therefore, by comparison with the point masses, we get $G\left({\varrho_{1-m},\varrho_{m}}\right)=\mu\eta^{-1}=\frac{\mu^{3}}{J^{2}}$ (one can verify this by using Legendre polynomials or refer to [16, Corollary 2.31]), and $T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)\leq\frac{\mu^{3}}{2J^{2}}$ similarly as in Corollary 5.4.
+Again, by Theorem 3.6, $\varrho_{1-m}$ and $\varrho_{m}$ have some nice properties, in particular they are spherically symmetric and radially decreasing after translations.
+Therefore, by comparison with the point masses, we get $G\left({\varrho_{1-m},\varrho_{m}}\right)=\mu_{r}\eta^{-1}=\frac{\mu_{r}^{3}}{J^{2}}$ (one can verify this by using Legendre polynomials or refer to [24] [11, Section 2]), and $T_{J}\left({\varrho_{1-m}+\varrho_{m}}\right)\leq\frac{\mu_{r}^{3}}{2J^{2}}$ similar to that in Corollary 5.4.
 Thus, the right side of (41) is less than 0.
@@
 $\geq E_{0}\left(\varrho_{1-m}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)-G\left({\rho_{1-m},\rho_{m}}\right)$
-Therefore, we get $\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)<G\left({\rho_{1-m},\rho_{m}}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)$, or $E_{0}\left(\widetilde{\rho_{m}}\right)<\frac{A^{\gamma}}{B^{3}}G\left({\rho_{1-m},\rho_{m}}\right)+E_{0}(\sigma)$.
+Therefore, we get $\frac{B^{3}}{A^{\gamma}}E_{0}\left(\widetilde{\rho_{m}}\right)<G\left({\rho_{1-m},\rho_{m}}\right)+\frac{B^{3}}{A^{\gamma}}E_{0}(\sigma)$ (i.e.
+$E_{0}\left(\widetilde{\rho_{m}}\right)<\frac{A^{\gamma}}{B^{3}}G\left({\rho_{1-m},\rho_{m}}\right)+E_{0}(\sigma)$).
 Since $\frac{A^{\gamma}}{B^{3}}=m^{\frac{6-5\gamma}{3\gamma-4}}$, $G\left({\rho_{1-m},\rho_{m}}\right)\leq Cm^{3}$ due to Remark 5.2, we have
@@
 Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14.
-Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A\rho_{m}\left({Bx}\right)$ as in Definition 5.1, we have, up to translation, $\{\widetilde{\rho_{m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
+Let $0<m<m_{0}$, and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A\rho_{m}\left({Bx}\right)$ as in Definition 5.1, we have, up to translation, $\{\widetilde{\rho_{m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
 Here $\sigma$ is the non-rotating minimizer with mass 1.
@@
 Due to Proposition 5.8, we know ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{m}}\right)}}=e_{0}<0$.
-Moreover, we can apply the concentration-compactness lemma of Lions [42, Theorem II.2 and Corollary II.1] to get that $\{\widetilde{\rho_{m}}\}$ is relatively compact up to translations in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$.
+Moreover, we can apply the concentration-compactness lemma of Lions [22, Theorem II.2 and Corollary II.1] to get that $\{\widetilde{\rho_{m}}\}$ is relatively compact up to translations in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$.
 Thus, there is a subsequence $\{\widetilde{\rho_{m_{k}}}\}$ that converges, up to translations, to some function $f$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$, $L^{1}\left(\mathbb{R}^{3}\right)$ and almost everywhere in $\mathbb{R}^{3}$.
@@
 In particular, $\left\|f\right\|_{L^{1}}=1$.
-And we claim that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}$.
-In fact, we recall $E_{0}(\rho)={\int{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}-\frac{1}{2}{\iint{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}$.
-For the inertial energy $\int{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}$, by $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$.
+And we claim that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}$. In fact, we recall
+$E_{0}(\rho)={\int_{\mathbb{R}^{3}}{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}\,dx}}-\frac{1}{2}{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}\,dx\,dy}}$
+For the inertial energy $\int_{\mathbb{R}^{3}}{\frac{K\left({\rho(x)}\right)^{\gamma}}{\gamma-1}\,dx}$, by $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$.
 a.e.
-and Fatou’s Lemma we have ${\int{\frac{K\left({f(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}\leq\lim\limits_{k\rightarrow\infty}\inf{\int{\frac{K\left({\widetilde{\rho_{m_{k}}}(x)}\right)^{\gamma}}{\gamma-1}d^{3}x}}$.
-For the potential energy $\frac{1}{2}{\iint{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}d^{3}xd^{3}y}}=\frac{1}{2}{\int{\rho(x)V_{\rho}(x)d^{3}x}}$, we have
-$\begin{split}&\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &=\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}+{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &\leq\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}}\right|+\left|{{\int{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\\ &\leq\left\|{\widetilde{\rho_{m_{k}}}-f}\right\|_{L^{\frac{4}{3}}}\cdot\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}+\left\|f\right\|_{L^{\frac{12}{11}}}\cdot\left\|{V_{f}-V_{\widetilde{\rho_{m_{k}}}}}\right\|_{L^{12}}\end{split}$
-The last inequality above comes from the Hölder’s inequality.
-Thanks to Hardy-Littlewood-Sobolev Inequality Proposition A.3, we have $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$, and $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$.
+and Fatou’s Lemma we have ${\int_{\mathbb{R}^{3}}{\frac{K\left({f(x)}\right)^{\gamma}}{\gamma-1}\,dx}}\leq\lim\limits_{k\rightarrow\infty}\inf{\int_{\mathbb{R}^{3}}{\frac{K\left({\widetilde{\rho_{m_{k}}}(x)}\right)^{\gamma}}{\gamma-1}\,dx}}$.
+For the potential energy $\frac{1}{2}{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho(x)\cdot\rho(y)}{\left|{x-y}\right|}\,dx\,dy}}=\frac{1}{2}{\int_{\mathbb{R}^{3}}{\rho(x)V_{\rho}(x)\,dx}}$, we have
+$\begin{split}&\quad\,\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\\ &=\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}+{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\\ &\leq\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}}\right|+\left|{{\int_{\mathbb{R}^{3}}{f(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\\ &\leq\left\|{\widetilde{\rho_{m_{k}}}-f}\right\|_{L^{\frac{4}{3}}}\cdot\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}+\left\|f\right\|_{L^{\frac{12}{11}}}\cdot\left\|{V_{f}-V_{\widetilde{\rho_{m_{k}}}}}\right\|_{L^{12}}\end{split}$
+The last inequality above comes from Hölder’s inequality.
+Thanks to Hardy-Littlewood-Sobolev Inequality (Proposition A.1), we have $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$, and $\left.\widetilde{\rho_{m_{k}}}\rightarrow f\right$.
 in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ implies $\left.V_{\widetilde{\rho_{m_{k}}}}\rightarrow V_{f}(x)\right$.
 in $L^{12}\left(\mathbb{R}^{3}\right)$.
-Thanks to the Interpolation Inequality, we have $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}\leq\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{1}}\cdot\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{4}{3}}}$, $\left\|f\right\|_{L^{\frac{12}{11}}}\leq\left\|f\right\|_{L^{1}}\cdot\left\|f\right\|_{L^{\frac{4}{3}}}$.
+Due to the Interpolation Inequality [6, Section 4.2], we have $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}\leq\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{1}}\cdot\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{4}{3}}}$, and $\left\|f\right\|_{L^{\frac{12}{11}}}\leq\left\|f\right\|_{L^{1}}\cdot\left\|f\right\|_{L^{\frac{4}{3}}}$.
 Thus $\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$ and $\left\|f\right\|_{L^{\frac{12}{11}}}$ are bounded, which—together with the inequality $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}\leq C\left\|\widetilde{\rho_{m_{k}}}\right\|_{L^{\frac{12}{11}}}$—implies the boundedness of $\left\|V_{\widetilde{\rho_{m_{k}}}}\right\|_{L^{4}}$.
 Then
-$\left.\left|{{\int{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)d^{3}x}}-{\int{f(x)V_{f}(x)d^{3}x}}}\right|\rightarrow 0\right$.
+$\left.\left|{{\int_{\mathbb{R}^{3}}{\widetilde{\rho_{m_{k}}}(x)V_{\widetilde{\rho_{m_{k}}}}(x)\,dx}}-{\int_{\mathbb{R}^{3}}{f(x)V_{f}(x)\,dx}}}\right|\rightarrow 0\right$.
 Thus, we know the claim above holds that $E_{0}(f)\leq{\lim\limits_{k\rightarrow\infty}{E_{0}\left(\widetilde{\rho_{m_{k}}}\right)}}=e_{0}=E_{0}(\sigma)$.
-In fact, the convergence up to a subsequence can be strengthened to convergence of the whole sequence.
+In fact, the convergence up to a subsequence can be strengthened to the convergence of the whole sequence.
 Since the non-rotating minimizer is unique up to translation due to Theorem 3.6, we get $\left\{\widetilde{\rho_{m_{k}}}\right\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ up to translations.
@@
 Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14.
-Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{1-m}}=A(1-m)\rho_{1-m}\left({B(1-m)x}\right)$ as in Definition 5.1, we have, up to translations, $\{\widetilde{\rho_{1-m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
+Let $0<m<m_{0}$, and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{1-m}}=A(1-m)\rho_{1-m}\left({B(1-m)x}\right)$ as in Definition 5.1, we have, up to translations, $\{\widetilde{\rho_{1-m}}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
 Proof.
@@
 $E_{0}\left(\widetilde{\rho_{1-m}}\right)<C(1-m)^{\frac{6-5\gamma}{3\gamma-4}}\cdot m^{3}+E_{0}(\sigma)$
-Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow the similar proof as in Theorem 5.9 above to get the result. ∎
+Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\widetilde{\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow a proof similar to that in Theorem 5.9 above to get the result. ∎
 Corollary 5.11 (Convergence of Densities of Stars in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$).
 Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14.
-Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, we have, up to translations, $\{\rho_{1-m}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
+Let $0<m<m_{0}$, and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, we have, up to translations, $\{\rho_{1-m}\}$ converges strongly to $\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ and $L^{1}\left(\mathbb{R}^{3}\right)$ as $m\rightarrow 0$.
 Proof.
@@
 $E_{0}\left(\rho_{1-m}\right)<Cm^{3}+(1-m)^{\frac{5\gamma-6}{3\gamma-4}}E_{0}(\sigma)$
-Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left({\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow the similar proof as in Theorem 5.9 above to get the result. ∎
+Thus ${\lim\limits_{m\rightarrow 0}{E_{0}\left({\rho_{1-m}}\right)}}=e_{0}<0$, and we can follow a proof similar to that in Theorem 5.9 above to get the result. ∎
 Remark 5.12.
@@
 $E_{0}\left(\rho_{m}\right)<Cm^{3}+m^{\frac{5\gamma-6}{3\gamma-4}}\cdot E_{0}(\sigma)$
-Then we have ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\rho_{m}\right)}}=0$, which violates the condition of concentration-compactness lemma [42, Theorem II.2 and Corollary II.1].
+Then we have ${\lim\limits_{m\rightarrow 0}{E_{0}\left(\rho_{m}\right)}}=0$, which violates the condition of concentration-compactness lemma [22, Theorem II.2 and Corollary II.1].
 In this case, we can actually show that $\rho_{m}$ converges to 0: from the scaling relation (37) and the fact $\lim\limits_{m\rightarrow 0}{A(m)}$$\infty$ and ${\lim\limits_{m\rightarrow 0}{B(m)}}=0$, we can show $\left.\left\|\rho_{m}\right\|_{L^{\frac{4}{3}}}=\left({A(m)}\right)^{-\frac{4}{3}}\left({B(m)}\right)^{3}\left\|\widetilde{\rho_{m}}\right\|_{L^{\frac{4}{3}}}\rightarrow 0\right$.
@@
 However, this vanishing limit result is not enough.
-We seek a stronger estimate that will allow us to estimate the Lagrange multiplier and then the support size of $\rho_{m}$ using the arguments in the following.
+We seek a stronger estimate that will allow us to estimate the Lagrange multiplier and then the support size of $\rho_{m}$, as we will see later.
 This is the reason we introduce the scaling density.
 Remark 5.13.
-Note the convergence holds true for the whole sequence instead of up to subsequence.
+Note the convergence holds true for the whole sequence instead of up to a subsequence.
 5.2 Uniform bound of scaling densities in $L^{\infty}\left(\mathbb{R}^{3}\right)$
 Thanks to the convergence results in subsection 5.1, now we can estimate the uniform $L^{\infty}$ bound of scaling densities $\left\{\widetilde{\rho_{m}}\right\}$ when $m$ is sufficiently small.
-Since $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ (Theorem 5.9), thus when $m$ is sufficiently small, $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$.
-The next steps are the following: (1) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{12}\left(\mathbb{R}^{3}\right)$; (2) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{6}\left(\mathbb{R}^{3}\right)$; (3) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$; (4) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$.
-Since we consider $\left\{\widetilde{\rho_{m}}\right\}$ and $\left\{V_{\widetilde{\rho_{m}}}\right\}$ alternatively, we call such method the bootstrap method.
+Since $\widetilde{\rho_{m}}\rightarrow\sigma$ in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ (Theorem 5.9), we know $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)$ when $m$ is sufficiently small.
+The next steps are to apply bootstrap method as follows: (1) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{12}\left(\mathbb{R}^{3}\right)$; (2) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{6}\left(\mathbb{R}^{3}\right)$; (3) $\left\{V_{\widetilde{\rho_{m}}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$; (4) $\left\{\widetilde{\rho_{m}}\right\}$ are bounded uniformly in $L^{\infty}\left(\mathbb{R}^{3}\right)$.
 Theorem 5.14 (Uniform bound of Scaling Densities of Planets and Potentials in $L^{\infty}\left(\mathbb{R}^{3}\right)$).
 Fix $J>0$, $m_{0}>0$ which is given in Theorem 4.14.
-Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A(m)\rho_{m}\left({B(m)x}\right)$ as in Definition 5.1, then $\exists C>0$, $\delta>0$, $\forall 0<m<\delta$, we have $\widetilde{\rho_{m}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, $V_{\widetilde{\rho_{m}}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, where $V_{\widetilde{\rho_{m}}}(x)={\int{\frac{\widetilde{\rho_{m}}(y)}{\left|{x-y}\right|}d^{3}y}}$ is the potential from $\widetilde{\rho_{m}}$.
+Let $0<m<m_{0}$ and $\rho(m)=\rho_{1-m}+\rho_{m}$ is the constrained minimizer on $W_{m}$, $\widetilde{\rho_{m}}=A(m)\rho_{m}\left({B(m)x}\right)$ as in Definition 5.1, then $\exists C>0$, $\delta>0$, $\forall 0<m<\delta$, we have $\widetilde{\rho_{m}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, $V_{\widetilde{\rho_{m}}}\in L^{\infty}\left(\mathbb{R}^{3}\right)$, and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}(\mathbb{R}^{3})}\leq C$, where $V_{\widetilde{\rho_{m}}}(x)={\int_{\mathbb{R}^{3}}{\frac{\widetilde{\rho_{m}}(y)}{\left|{x-y}\right|}\,dy}}$ is the potential from $\widetilde{\rho_{m}}$.
 Moreover, $V_{\widetilde{\rho_{m}}}(x)$ is continuous.
@@
 […]
-$\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}=\left({\int{{\widetilde{\rho_{m}}}^{12{({\gamma-1})}}(x)d^{3}x}}\right)^{\frac{1}{12{({\gamma-1})}}}\leq{\widetilde{C}\left({\int{V_{\widetilde{\rho_{m}}}^{12}(x)d^{3}x}}\right)}^{\frac{1}{12{({\gamma-1})}}}=\widetilde{C}\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C_{2}$
-By assumption $\gamma>\frac{3}{2}>\frac{4}{3}$, together with Interpolation Inequality Proposition A.2 we have $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}}\leq\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}^{\alpha}\cdot\left\|\widetilde{\rho_{m}}\right\|_{L^{1}}^{1-\alpha}\leq C$ for some $C>0$.
-Since we already know $\left\|\widetilde{\rho_{m}}\right\|_{L^{1}{(\mathbb{R}^{3})}}=1$ and $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}{(\mathbb{R}^{3})}}$ are bounded uniformly, thanks to Proposition A.4, we have $V_{\widetilde{\rho_{m}}}$ is continuous and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}{(\mathbb{R}^{3})}}$ is bounded uniformly.
+$\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}=\left({\int_{\mathbb{R}^{3}}{{\widetilde{\rho_{m}}}^{12{({\gamma-1})}}(x)\,dx}}\right)^{\frac{1}{12{({\gamma-1})}}}\leq{\widetilde{C}\left({\int_{\mathbb{R}^{3}}{V_{\widetilde{\rho_{m}}}^{12}(x)\,dx}}\right)}^{\frac{1}{12{({\gamma-1})}}}=\widetilde{C}\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{12}}^{\frac{1}{\gamma-1}}\leq C_{2}$
+By assumption $\gamma>\frac{3}{2}>\frac{4}{3}$, together with Interpolation Inequality [6, Section 4.2] we have $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}}\leq\left\|\widetilde{\rho_{m}}\right\|_{L^{12{({\gamma-1})}}}^{\alpha}\cdot\left\|\widetilde{\rho_{m}}\right\|_{L^{1}}^{1-\alpha}\leq C$ for some $C>0$.
+Since we already know $\left\|\widetilde{\rho_{m}}\right\|_{L^{1}{(\mathbb{R}^{3})}}=1$ and $\left\|\widetilde{\rho_{m}}\right\|_{L^{12}{(\mathbb{R}^{3})}}$ are bounded uniformly, due to Proposition A.2, we have $V_{\widetilde{\rho_{m}}}$ is continuous and $\left\|V_{\widetilde{\rho_{m}}}\right\|_{L^{\infty}{(\mathbb{R}^{3})}}$ is bounded uniformly.
 Again by ${\widetilde{\rho_{m}}}^{\gamma-1}\leq CV_{\widetilde{\rho_{m}}}$ we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is also bounded uniformly.
@@
 Remark 5.15.
-The uniform bound in $L^{\infty}(\mathbb{R}^{3})$ can be comparable with Li’s result [38, Proposition 1.4], where the rotating stars (corresponding to single-star case) are considered and the bound $C(m)$ however depends on $m$.
-Although one might conjecture $C(m)$ is continuous when $m>0$, they may still need to consider if $\lim\limits_{m\rightarrow 0}{\sup{C(m)}}$$\infty$ in order to obtain a result analogous to Theorem 5.14.
+The uniform bound in $L^{\infty}(\mathbb{R}^{3})$ can be comparable with Li’s result [18, Proposition 1.4], where the rotating single stars are considered and the bound $C(m)$ however depends on $m$.
+Hence, if it can be shown that $\lim\limits_{m\rightarrow 0}{\sup{C(m)}}$$\infty$, then an analogous result to Theorem 5.14 should also hold.
 Remark 5.16.
-By the similar arguments, we can obtain that when $m$ is sufficient small, $\left\{\widetilde{\rho_{1-m}}\right\}$ and $\left\{\rho_{1-m}\right\}$ are also bounded uniformly in $L^{\infty}(\mathbb{R}^{3})$.
+By the similar arguments, we can obtain that when $m$ is sufficiently small, $\left\{\widetilde{\rho_{1-m}}\right\}$ and $\left\{\rho_{1-m}\right\}$ are also bounded uniformly in $L^{\infty}(\mathbb{R}^{3})$.
 Remark 5.17.
-Similarly as Remark 3.12, once we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly, by definition we have $\rho_{m}=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)$, $A=m^{-\frac{2}{3\gamma-4}}$, thus ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$.
+Similar to the discussion in Remark 3.10, once we know $\left\|\widetilde{\rho_{m}}\right\|_{L^{\infty}(\mathbb{R}^{3})}$ is bounded uniformly, by definition we have $\rho_{m}=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)$, $A=m^{-\frac{2}{3\gamma-4}}$, thus ${\lim\limits_{m\rightarrow 0}\left\|\rho_{m}\right\|_{L^{\infty}(\mathbb{R}^{3})}}=0$.
 5.3 Estimate for the supports of densities
@@
 We first seek an upper bound for the Lagrange multiplier.
-To avoid confusion with the Lagrange multiplier $\lambda_{m}$ corresponding to star planet system, we will use $\kappa_{m}$ to represent the corresponding non-rotating Lagrange multiplier with mass $m$ in this subsection.
+To avoid confusion with the Lagrange multiplier $\lambda_{m}$ corresponding to star-planet system, we will use $\kappa_{m}$ to represent the corresponding non-rotating Lagrange multiplier with mass $m$ in this subsection.
 Since both the function $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ in (2) and the scaling coefficient $A$ in (37) will appear frequently in this subsection, to avoid confusion, we will use $\mathcal{A}(\rho)$ to denote the function $A(\rho)$, while keeping $A$ to represent the coefficient $A$ in this subsection.
-Similar to the arguments in the proof of Theorem 4.5 (see also [4, Lemma 2] [15, Proposition 5.8] [16, Lemma 2.21]), it can be shown that given a constrained minimizer $\rho(m)$, then $\rho(m)$ satisfies Euler-Lagrange equation in the following sense:
+Similar to the arguments in the proof of Theorem 4.5 (see also [4, Lemma 2] [10, Section 5] [11, Section 2]), it can be shown that given a constrained minimizer $\rho(m)$, $\rho(m)$ satisfies Euler-Lagrange equation in the following sense:
 $\mathcal{A}^{\prime}(\rho(m)(x))=\left[\frac{J^{2}}{2I^{2}(\rho(m))}r^{2}(x-\bar{x}(\rho(m)))+V_{\rho(m)}(x)+\lambda_{\alpha}\right]_{+}\text{on }\Omega_{\alpha}$ (ELm)
@@
 $\lambda_{m}\leq\mathcal{A}^{\prime}\left(\rho_{m}\right)-V_{\rho_{m}}\text{ a.e. on }\Omega_{m}$
-where $\mathcal{A}(\rho)(x)=\frac{K\left(\rho^{\gamma}(x)\right)}{\gamma-1},\mathcal{A}^{\prime}(\rho)(x)=\frac{K\gamma\left(\rho^{\gamma-1}(x)\right)}{\gamma-1}$.
-Due to (30) (31), we have $\mathcal{A}^{\prime}\left(\rho_{m}\right)(x)=\mathcal{A}^{\prime}\left(\frac{1}{A}\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)$, $V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$.
+where $\mathcal{A}(\rho)(x)=\frac{K\left(\rho^{\gamma}(x)\right)}{\gamma-1}$, $\mathcal{A}^{\prime}(\rho)(x)=\frac{K\gamma\left(\rho^{\gamma-1}(x)\right)}{\gamma-1}$. Due to (31) (32), we have
+$\mathcal{A}^{\prime}\left(\rho_{m}\right)(x)=\mathcal{A}^{\prime}\left(\frac{1}{A}\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)$
+$V_{\rho_{m}}(x)=\frac{B^{2}}{A}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)=\frac{1}{A^{\gamma-1}}V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$
 Therefore, we have
@@
 on $\mathbb{R}^{3}$, let $E_{n}=\left\{x\in\mathbb{R}^{3}\mid\frac{C}{n}\leq\sigma(x)<C\right\}$, then $\mathbb{R}^{3}=\cup_{n=1}^{\infty}E_{n}$.
-And then $1=\int_{\mathbb{R}^{3}}\sigma=\sum_{n=1}^{\infty}\int_{E_{n}}\sigma$.
-Then $\exists N>0$, such that $\frac{1}{2}<\sum_{n=1}^{N}\int_{E_{n}}\sigma<\sum_{n=1}^{N}\int_{E_{n}}C=C\mu(\{\{x\in\mathbb{R}^{3}\mid\frac{C}{N}\leq\sigma(x)<C\})$.
-Therefore, let $M:=\mu\left(\left\{x\in\mathbb{R}^{3}\mid\frac{C}{N}<\sigma(x)<C\right\}\right)$, then $0<\frac{1}{2C}<M<\infty$.
-Thus $\exists\delta>0,\forall 0<m<\delta$, due to the claim we have
+And then $1=\int_{\mathbb{R}^{3}}\sigma\,dx=\sum_{n=1}^{\infty}\int_{E_{n}}\sigma\,dx$.
+Then $\exists N>0$, such that
+$\frac{1}{2}<\sum_{n=1}^{N}\int_{E_{n}}\sigma\,dx<\sum_{n=1}^{N}\int_{E_{n}}C\,dx=C\mu(\{\{x\in\mathbb{R}^{3}\mid\frac{C}{N}\leq\sigma(x)<C\})$
+Therefore, let $M:=\mu\left(\left\{x\in\mathbb{R}^{3}\mid\frac{C}{N}<\sigma(x)<C\right\}\right)$, then $0<\frac{1}{2C}<M<\infty$. Thus $\exists\delta>0$, $\forall 0<m<\delta$, due to the claim we have
 $\mu\left(\left\{x\in\mathbb{R}^{3}\mid\left|\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{x}{B}\right)-\mathcal{A}^{\prime}(\sigma)\left(\frac{x}{B}\right)\right|\geq\frac{\epsilon}{3}\right\}\right)$
@@
 […]
+$\widetilde{\rho_{m}}\left(\frac{x}{B}\right)\geq\sigma\left(\frac{x}{B}\right)-\left|\widetilde{\rho_{m}}\left(\frac{x}{B}\right)-\sigma\left(\frac{x}{B}\right)\right|>\frac{C}{N}-\frac{C}{3N}>0$
+Hence $\rho_{m}(x)=\frac{1}{A}\widetilde{\rho_{m}}\left(\frac{x}{B}\right)>0$. Thus $x\in\Omega_{m}$ (up to translations),
 $A^{\gamma-1}\lambda_{m}$ $\leq\mathcal{A}^{\prime}\left(\widetilde{\rho_{m}}\right)\left(\frac{1}{B}x\right)-V_{\widetilde{\rho_{m}}}\left(\frac{1}{B}x\right)$ (42)
@@
 Note that the above proof relies on convergence in measure.
-In McCann’s paper [46, Lemma 6.5], only almost everywhere convergence up to subsequence is used.
-Nonetheless, since in this paper we assume $m$ is very small, relying solely on almost everywhere convergence up to subsequence will lead to the following two issues.
-One is that it is up to a subsequence, i.e.
-$\widetilde{\rho_{m_{k}}}(x)\rightarrow\sigma(x)$ a.e.
-on $\mathbb{R}^{3}$ instead of $\left\{\widetilde{\rho_{m}}\right\}$ itself, while we hope the bound for the Lagrange multiplier should be valid for all sufficient small mass.
+In McCann’s paper [24, Lemma 6.5], only almost everywhere convergence up to a subsequence is used.
+Nonetheless, since in this paper we assume $m$ is very small, relying solely on almost everywhere convergence up to a subsequence will lead to the following two issues.
+One is that it is up to a subsequence, i.e., $\widetilde{\rho_{m_{k}}}(x)\rightarrow\sigma(x)$ a.e.
+on $\mathbb{R}^{3}$ instead of $\left\{\widetilde{\rho_{m}}\right\}$ itself, while we hope the bound for the Lagrange multiplier should be valid for all sufficiently small mass.
 Another one is when $m\rightarrow 0$, not only $\widetilde{\rho_{m}}(x)\rightarrow\sigma(x)$ a.e.
@@
 Therefore, we make more use of $L^{p}$ convergence—namely, that it implies convergence in measure, to solve the issues.
-Before we finally estimate the sizes of supports of scaling densities, we supplement the proof of a proposition about the $L^{\infty}$ bound of potential mentioned in [46, Proposition 6.6], which can be viewed as a complement to [4, Proposition 5], in which the indices $\frac{2}{3}$ and $\frac{1}{3}$ can be approximated but cannot be reached.
-Proposition 5.20 (Bound of Potential in $L^{\infty}\left(\mathbb{R}^{3}\right)$ [46, Section 6]).
-$\exists k>0,\forall\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, one has $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$, where $V_{\rho}(x)=\int\frac{\rho(y)}{|x-y|}d^{3}y$ is the potential from $\rho$.
+Before we finally estimate the sizes of supports of scaling densities, we supplement the proof of a proposition about the $L^{\infty}$ bound of potential mentioned in [24, Proposition 6.6], which can be viewed as a complement to [4, Proposition 5], in which the indices $\frac{2}{3}$ and $\frac{1}{3}$ can be approximated but cannot be reached.
+Proposition 5.20 (Bound of Potential in $L^{\infty}\left(\mathbb{R}^{3}\right)$ [24, Section 6]).
+$\exists k>0,\forall\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, one has $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq k\|\rho\|_{L^{1}\left(\mathbb{R}^{3}\right)}^{\frac{2}{3}}\|\rho\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}^{\frac{1}{3}}$, where $V_{\rho}(x)=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$ is the potential from $\rho$.
 Proof.
-$V_{\rho}(x)$ $=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}d^{3}y$
+Let $B_{R}(\cdot)$ be the open ball defined by $B_{R}(\cdot)\coloneq\left\{x\in\mathbb{R}^{3}\mid|x-\cdot|<R\right\}$.
+$V_{\rho}(x)$ $=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy$
 […]
@@
 Take $m$ and $\epsilon$ small enough so that constrained minimizers $\rho(m)$ exist by Theorem 4.14 and $A(m)^{\gamma-1}\cdot\lambda_{m}\leq\kappa_{1}+\epsilon<$ 0 by Lemma 5.18, and the velocity $v(x,m)$ satisfies $v(x,m)^{2}\leq Cm^{2}$ by Corollary 5.5.
-By the construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\operatorname{dist}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{\eta}{2}$, $\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$, thus $\exists C>0,\forall x\in\Omega_{m}$, $V_{\rho_{1-m}}(x)<Cm^{2}$.
+By the construction of $\Omega_{m}$ and $\Omega_{1-m}$, we have $\operatorname{dist}\left(\Omega_{m},\Omega_{1-m}\right)=\frac{\eta}{2}$, where $\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$, thus $\exists C>0$, $\forall x\in\Omega_{m}$, $V_{\rho_{1-m}}(x)<Cm^{2}$.
 In the Euler-Lagrange equation (ELm), we know for almost every $x\in\Omega_{m}$,
@@
 Then $\widetilde{\rho_{m}}=\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$, $V_{\widetilde{\rho_{m}}}=V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}+V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}$.
-Given $y$ outside the larger ball of radius $\tilde{R}(J)$ centered on 0, $V_{\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}}(y)\leq-\frac{\kappa_{1}+\epsilon}{6}$ due to $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$.
+Given $y$ outside the larger ball of radius $\tilde{R}(J)$ centered on 0, due to $\frac{1}{\tilde{R}(J)-R_{0}}\leq-\frac{\kappa_{1}+\epsilon}{6}$, we have
+$V_{\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|<R_{0}\right\}}}(y)\leq-\frac{\kappa_{1}+\epsilon}{6}$
 For $V_{\widetilde{\rho_{m}}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}(y)$, by Theorem 5.14 we can show $\widetilde{\rho_{m}}\cdot\mathbf{1}_{\left\{|x|\geq R_{0}\right\}}\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$ with mass $\delta$.
@@
 For example, instead of applying Proposition 5.20, one can apply [4, Proposition 5] directly.
-Remark 5.23.
-In Lemma 5.18 and Theorem 5.21 we consider the the polytropic law (1) $P(\rho)=K\rho^{\gamma}$.
-McCann actually considered more general form of pressure $P(\rho)$ [46, Lemma 6.5 and Theorem 6.6].
-In [46], $\kappa_{1}$ is replaced by $e_{0}^{\prime}(1^{-})$, where $e_{0}^{\prime}(1^{-})$ is negative and is a upper bound of Lagrange multipliers corresponding to non-rotating minimizers with fixed mass $1$, even in cases where multiple non-rotating minimizers exist (see Theorem 3.6 and Remark 3.9).
-Nevertheless, since we are dealing with scaling density here, we focus on this polytropic law.
 Corollary 5.23 (Bound on the sizes of supports of original densities $\rho_{m}$).
 There exists a radius $R(J)$ and $\delta>0$ independent of $m,\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then spt $\rho_{m}$ is contained in a ball of radius $B\cdot R(J)$, where $B$ is defined in Definition 5.1.
-In particular, spt $\rho_{m}$ is contained in a ball of a radius which goes to 0 with rate $B$ since $\lim\limits_{m\rightarrow 0}B=0$ if $\gamma>2$.
+In particular, if $\gamma>2$, spt $\rho_{m}$ is contained in a ball whose radius goes to 0 with rate $B$ since $\lim\limits_{m\rightarrow 0}B=0$.
 Proof.
@@
 Our next step is to show the centers of mass can be chosen to be not too far away from the centers of $\Omega_{1-m}$ and $\Omega_{m}$.
-Thanks to those results, we then can show the constrained minimizers are actually $W^{\infty}$ local energy minimizers, which represent solutions to the reduced Euler-Poisson system (EP’) up to translation due to Theorem 3.1.
+Thanks to those results, we can then show the constrained minimizers are actually $W^{\infty}$ local energy minimizers, which represent solutions to the reduced Euler-Poisson system (EP’) up to translation due to Theorem 3.1.
 6.1 Estimate for the Center of Mass Separation
-In this subsection, we estimate for the centers of mass separation by modifying the arguments in [46, Lemma 6.7 and Proposition 6.8] such that they can be applied in our case.
+In this subsection, we estimate the centers of mass separation by modifying the arguments in [24, Lemma 6.7 and Proposition 6.8] such that they can be applied in our case.
 Lemma 6.1.
@@
 Proof.
-One can observe when $\epsilon$ is small enough, the function $g_{\epsilon}$ and $g_{\epsilon}^{\prime}$ are analytic and uniformly bounded on $\left\{z\in\mathbb{C}\mid|z|>\frac{1}{4}\right\}$.
-The result can be check by direct computation and estimation.
+One can observe when $\epsilon$ is small enough, the functions $g_{\epsilon}$ and $g_{\epsilon}^{\prime}$ are analytic and uniformly bounded on $\left\{z\in\mathbb{C}\mid|z|>\frac{1}{4}\right\}$.
+The result can be checked by direct computation and estimation.
 ∎
 Proposition 6.2 (Estimate for the Center of Mass Separation).
-Fix $J>0$, let $0<\zeta<\frac{1}{2},\exists\delta>0$, such that $\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then $1-\zeta<\frac{\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|}{\left|y_{m}-y_{1-m}\right|}<1+\zeta$, where $y_{m}$ and $y_{1-m}$ are the centers of $\Omega_{m}$ and $\Omega_{1-m},\left|y_{m}-y_{1-m}\right|=\eta=\frac{J^{2}}{\mu^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
+Fix $J>0$, let $0<\zeta<\frac{1}{2},\exists\delta>0$, such that $\forall 0<m<\delta$, if $\rho(m)=\rho_{1-m}+\rho_{m}$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, then $1-\zeta<\frac{\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|}{\left|y_{m}-y_{1-m}\right|}<1+\zeta$, where $y_{m}$ and $y_{1-m}$ are the centers of $\Omega_{m}$ and $\Omega_{1-m}$, $\left|y_{m}-y_{1-m}\right|=\eta=\frac{J^{2}}{\mu_{r}^{2}}=\frac{J^{2}}{(m(1-m))^{2}}$.
 Proof.
@@
 That is, $d:=\left|\bar{x}\left(\rho_{m}\right)-\bar{x}\left(\rho_{1-m}\right)\right|=r\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)$, where $r$ is given in Definition 2.1.
-Since spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$, we have (recall $\mu=m(1-m)$)
-$\frac{\mu}{d+2R}\leq G\left(\rho_{m},\rho_{1-m}\right)=\iint\frac{\rho_{m}(x)\cdot\rho_{1-m}(y)}{|x-y|}d^{3}xd^{3}y\leq\frac{\mu}{d-2R}$
+Since spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$, we have (recall $\mu_{r}=m(1-m)$)
+$\frac{\mu_{r}}{d+2R}\leq G\left(\rho_{m},\rho_{1-m}\right)=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\rho_{m}(x)\cdot\rho_{1-m}(y)}{|x-y|}\,dx\,dy\leq\frac{\mu_{r}}{d-2R}$
 By Lemma 2.4, we know
-$I(\rho(m))=\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+\int r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{1-m}(x)d^{3}x+\int r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{m}(x)d^{3}x$
+$I(\rho(m))=\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+\int_{\mathbb{R}^{3}}r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{1-m}(x)\,dx+\int_{\mathbb{R}^{3}}r^{2}\left(x-\bar{x}\left(\rho_{m}\right)\right)\rho_{m}(x)\,dx$
 Hence
-$\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)=\mu d^{2}\leq I(\rho(m))\leq\mu r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+(1-m)R^{2}+mR^{2}=\mu d^{2}+R^{2}$
+$\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)=\mu_{r}d^{2}\leq I(\rho(m))\leq\mu_{r}r^{2}\left(\bar{x}\left(\rho_{1-m}\right)-\bar{x}\left(\rho_{m}\right)\right)+(1-m)R^{2}+mR^{2}=\mu_{r}d^{2}+R^{2}$
 By definition of $T_{J}$ (16), we have
-$\frac{J^{2}}{2\left(\mu d^{2}+R^{2}\right)}\leq T_{J}(\rho(m))=\frac{J^{2}}{2I(\rho(m))}\leq\frac{J^{2}}{2\mu d^{2}}$
+$\frac{J^{2}}{2\left(\mu_{r}d^{2}+R^{2}\right)}\leq T_{J}(\rho(m))=\frac{J^{2}}{2I(\rho(m))}\leq\frac{J^{2}}{2\mu_{r}d^{2}}$
 Similar inequalities hold true for $G\left(\kappa_{m},\kappa_{1-m}\right)$ and $T_{J}(\kappa(m))$, but we replace $d$ by $\eta$, since $\left|\bar{x}\left(\kappa_{m}\right)-\bar{x}\left(\kappa_{1-m}\right)\right|=r\left(\bar{x}\left(\kappa_{1-m}\right)-\bar{x}\left(\kappa_{m}\right)\right)=\left|y_{m}-y_{1-m}\right|=\eta$.
-Note $\rho(m)$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, we have $E_{J}(\kappa(m))-E_{J}(\rho(m))=$ $G\left(\kappa_{m},\kappa_{1-m}\right)+T_{J}(\kappa(m))-G\left(\rho_{m},\rho_{1-m}\right)+T_{J}(\rho(m))\geq 0$, just as similar as (40).
-Using the previous estimates and $J^{2}=\mu^{2}\eta$, we have
-$\frac{1}{d-2R}+\frac{\mu\eta}{2\left(\mu d^{2}+R^{2}\right)}\leq-\frac{1}{\eta+2R}+\frac{1}{2\eta}$
+Note $\rho(m)$ minimizes $E_{J}(\rho)$ on ${W}_{m}$, based on the energy decomposition similar to (41), we have
+$E_{J}(\kappa(m))-E_{J}(\rho(m))=-G\left(\kappa_{m},\kappa_{1-m}\right)+T_{J}(\kappa(m))-G\left(\rho_{m},\rho_{1-m}\right)+T_{J}(\rho(m))\geq 0$
+Using the previous estimates and $J^{2}=\mu_{r}^{2}\eta$, we have
+$\frac{1}{d-2R}+\frac{\mu_{r}\eta}{2\left(\mu_{r}d^{2}+R^{2}\right)}\leq-\frac{1}{\eta+2R}+\frac{1}{2\eta}$
 Let $x=\frac{d}{\eta}$, $\epsilon=\frac{R}{\eta}$, then
@@
 for $m$ small enough since $\epsilon=\frac{R}{\eta}=\frac{Rm^{2}(1-m)^{2}}{J^{2}}$).
-Thus (3.3.1) implies $x=\frac{d}{\eta}$ has to satisfies $1-\zeta<\frac{d}{\eta}<1+\zeta$, which proves the proposition.
+Thus (3.3.1) implies $x=\frac{d}{\eta}$ has to satisfy $1-\zeta<\frac{d}{\eta}<1+\zeta$, which proves the proposition.
 ∎
 6.2 Existence theorem
-After making some minor modifications to McCann’s proof in [46, Section 6], we proceed by translating $\rho$ such that it enjoys a plane of symmetry $z=0$, and then use the results above to show $\rho$ is supported away from the boundary of $\Omega_{1-m}\cup\Omega_{m}$, then we finalize the proof of existence.
+After making some modifications to McCann’s proof in [24, Section 6], we proceed by translating $\rho$ such that it enjoys a plane of symmetry $z=0$, and then use the results above to show $\rho$ is supported away from the boundary of $\Omega_{1-m}\cup\Omega_{m}$, then we finalize the proof of existence.
 Lemma 6.3 (Support Separation from the Boundary).
@@
 In fact, since the $\Omega_{m}$ and $\Omega_{1-m}$ are convex and symmetric about $z=0$, it is enough to know that $\rho(m)$ enjoys a plane of symmetry $z=c$.
-This follows from a strong rearrangement inequality in Lieb [41, Lemma 3] and Fubini’s Theorem: by the definition of $U(\rho)$ and $I(\rho)$, they will not be changed after the symmetric decreasing rearrangement of $\rho$ along lines parallel to the z-axis; however, since $\left(r^{2}+z^{2}\right)^{-\frac{1}{2}}$ is strictly decreasing as a function of $|z|$, the rearrangement increases $G(\rho,\rho)$ unless $\rho$ is already symmetric decreasing about a plane $z=c$.
+This follows from a strong rearrangement inequality in Lieb [21, Lemma 3] and Fubini’s Theorem: by the definition of $U(\rho)$ and $I(\rho)$, they will not be changed after the symmetric decreasing rearrangement of $\rho$ along lines parallel to the z-axis; however, since $\left(r^{2}+z^{2}\right)^{-\frac{1}{2}}$ is strictly decreasing as a function of $|z|$, the rearrangement increases $G(\rho,\rho)$ unless $\rho$ is already symmetric decreasing about a plane $z=c$.
 Note $\Omega_{1-m}$ and $\Omega_{m}$ are convex, hence $\rho(m)$’s rearrangement is in $W_{m}$.
-Since $\rho(m)$ minimizes $E_{J}(\rho)$, then $G(\rho,\rho)$ cannot be increased after rearrangement, which means $\rho(m)$ enjoys a plane of symmetry $z=c$.
+Since $\rho(m)$ minimizes $E_{J}(\rho)$, $G(\rho,\rho)$ cannot be increased after rearrangement, which means $\rho(m)$ enjoys a plane of symmetry $z=c$.
 Now, similar to the arguments in the proof of Proposition 6.2, we take $m$ small enough so that Corollary 5.23 and Corollary 5.25 provide a bound $R$ such that spt $\rho_{m}\subset B_{R}\left(\bar{x}\left(\rho_{m}\right)\right)$ and spt $\rho_{1-m}\subset B_{R}\left(\bar{x}\left(\rho_{1-m}\right)\right)$ if $\rho(m)=\rho_{m}+\rho_{1-m}$ minimizes $E_{J}(\rho)$ on $W_{m}$.
@@
 Thus $E_{J}(\rho(m))\leq E_{J}(\kappa)$, and $\rho(m)$ is a Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$ (Existence of $W^{\infty}$ local energy minimizer).
-By definition of $W_{m}$, we know all the functions in $W_{m}$ have compact support.
+By the definition of $W_{m}$ (22), we know all the functions in $W_{m}$ have compact support.
 Thus after translation we can set $\rho(m)\in{R}_{0}\left(\mathbb{R}^{3}\right)$.
 Since $\rho(m)$ is a $W^{\infty}$ local energy minimizer on ${R}\left(\mathbb{R}^{3}\right)$, in particular on ${R}_{0}\left(\mathbb{R}^{3}\right)$, Theorem 3.1 provides a local minimizer $(\rho(m),v)$ of $E(\rho,v)$ subject to the constraint on $J_{z}=J$, with $v(x)=$ $\frac{J}{I(\rho(m))}\hat{e}_{z}\times x$.
-[46, Section 2] and [15, Remark 2.31] show that $(\rho(m),v)$ minimizes $E(\rho,v)$ subject to the constraint on the vector angular momentum as well (Part (i)).
+[24, Section 2] and [10, Section 2] show that $(\rho(m),v)$ minimizes $E(\rho,v)$ subject to the constraint on the vector angular momentum as well (Part (i)).
 Part (ii) comes from Lemma 6.3 (Part (ii)).
@@
 The proof of Theorem 2.13 is essentially the same as the proof of Theorem 2.12.
-Two things we need to note are (1) if we don’t know $\gamma>2$, we may not have the fact that there is a “uniform bound on the radius of $\rho(m)$’s support for all small mm” from Corollary 5.24.
+Two things we need to note are (1) if we do not know $\gamma>2$, we may not have the fact that there is a “uniform bound on the radius of $\rho(m)$’s support for all small mm” from Corollary 5.23.
 However, since the radii of $\Omega_{1-m}$ and $\Omega_{m}$ is $\frac{\eta}{4}\sim m^{-2}$, recall $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$, due to Corollary 5.23, we can still have the fact that $\rho(m)$ is located in the interior of $\Omega_{m}\cup\Omega_{1-m}$ if ${\frac{\gamma-2}{3\gamma-4}}>-2$, that is $\gamma>\frac{10}{7}$; (2) in order to obtain the energy convergence in Proposition 5.8, we hope the term $Cm^{\frac{4\gamma-6}{3\gamma-4}}$ in (40) satisfies $\frac{4\gamma-6}{3\gamma-4}>0$, which means $\gamma>\frac{3}{2}$.
-Therefore, when $\gamma>\max\{\frac{3}{2},\frac{10}{7}\}=\frac{3}{2}$, we have Theorem 2.13 holds true.
+Therefore, when $\gamma>\max\{\frac{3}{2},\frac{10}{7}\}=\frac{3}{2}$, Theorem 2.13 holds.
 ∎
@@
 It should be noted that, although we prove the existence of Wasserstein $L^{\infty}$ local minimizer on $R(\mathbb{R}^{3}$), which is also a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ on $W_{m}$ (Theorem 2.12 and Theorem 2.13), in particular, spt $\rho(m)$ has at least two connected components, it is still unclear whether it has exactly two connected components.
-Neither we in this paper nor McCann [46, Theorem 6.1 and Corollary 6.2] elaborated on or provided proof for the result.
+Neither we in this paper nor McCann [24, Theorem 6.1 and Corollary 6.2] elaborated on or provided proof for the result.
 If spt $\rho_{1-m}$ or spt $\rho_{m}$ have multiple connected components, the model may correspond to a system with multiple stars and planets, such as the relationship between the Solar System and the asteroid belt.
@@
 Then we know $d=\inf\limits_{y\in V}y_{1}-\sup\limits_{x\in U}x_{1}$.
-We denote the mass in $U$ is $m_{1}\coloneq\int_{U}\rho_{m}$, then the mass in $V$ is $m_{2}\coloneq\int_{V}\rho_{m}=m-m_{1}$.
+We denote the mass in $U$ by $m_{1}\coloneq\int_{U}\rho_{m}\,dx$, then the mass in $V$ is $m_{2}\coloneq\int_{V}\rho_{m}\,dx=m-m_{1}$.
 Notice $\rho_{m}(x)=\rho_{U}(x)+\rho_{V}(x)$, where $\rho_{U}(x)=\rho_{m}(x)\cdot\mathbf{1}_{U}(x)$, $\rho_{V}(x)=\rho_{m}(x)\cdot\mathbf{1}_{V}(x)$, $\mathbf{1}_{E}$ is the indicator function of set $E$.
-We want to move $\rho_{U}(x)$ and $\rho_{V}(x)$ in the following way: given $h>0$, let $h_{1}>0$ and $h_{2}>0$ satisfy $h_{1}+h_{2}=h$, $h_{1}\cdot m_{1}=h_{2}\cdot m_{2}$.
+We want to move $\rho_{U}(x)$ and $\rho_{V}(x)$ in the following way: given $h>0$, let $h_{1}>0$ and $h_{2}>0$ satisfy $h_{1}+h_{2}=h$ and $h_{1}\cdot m_{1}=h_{2}\cdot m_{2}$.
 With an abuse of notation we denote $h_{1}=h_{1}\cdot\hat{e_{1}}=(h_{1},0,0)^{T}$, similarly $h_{2}=(h_{2},0,0)^{T}$, $h=(h,0,0)^{T}$.
-Let $\rho_{m}^{[h]}=\rho_{U}(x-h_{1})+\rho_{V}(x+h_{2})$.
-Easy to see when $h$ is small enough, the connected components of spt $\rho_{m}^{[h]}$ are $U_{h}=U+h_{1}$, and $V_{h}=V-h_{2}$, and we still have $\sup\limits_{x\in U_{h}}x_{1}<b<\inf\limits_{y\in V_{h}}y_{1}$.
+Then we define $\rho_{m}^{[h]}=\rho_{U}(x-h_{1})+\rho_{V}(x+h_{2})$.
+Easy to see when $h$ is small enough, the connected components of spt $\rho_{m}^{[h]}$ are $U_{h}=U+h_{1}$ and $V_{h}=V-h_{2}$, and we still have $\sup\limits_{x\in U_{h}}x_{1}<b<\inf\limits_{y\in V_{h}}y_{1}$.
 We can write $\rho_{m}^{[h]}=\rho_{U}^{[h]}+\rho_{V}^{[h]}$ where $\rho_{U}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{U_{h}}$ and $\rho_{V}^{[h]}=\rho_{m}^{[h]}\cdot\mathbf{1}_{V_{h}}$.
@@
 Notice ${\bar{x}}\left(\rho_{V}^{[h]}\right)-{\bar{x}}\left(\rho_{U}^{[h]}\right)={\bar{x}}\left(\rho_{V}\right)-{\bar{x}}\left(\rho_{U}\right)-h$, thus we have
-$\begin{split}I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)=\frac{m_{1}m_{2}}{m}\left({-2h\left({{\bar{x}}_{1}\left(\rho_{V}\right)-{\bar{x}}_{1}\left(\rho_{U}\right)}\right)+h^{2}}\right)=\frac{1}{m}{\int{-2h\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)d^{3}xd^{3}y}}+\frac{m_{1}m_{2}}{m}h^{2}\end{split}$
+$\begin{split}I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)&=\frac{m_{1}m_{2}}{m}\left({-2h\left({{\bar{x}}_{1}\left(\rho_{V}\right)-{\bar{x}}_{1}\left(\rho_{U}\right)}\right)+h^{2}}\right)\\ &=\frac{1}{m}{(-2h)\iint{\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)\,dx\,dy}}+\frac{m_{1}m_{2}}{m}h^{2}\end{split}$
 Therefore, we have
-$\begin{split}&\quad T_{J}\left({\rho(m)}\right)-T_{J}\left({\rho^{[h]}(m)}\right)\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)}-\frac{J^{2}}{2I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}\left({I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)}\right)}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}\cdot{\left({{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{-2h\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)d^{3}xd^{3}y}}+m_{1}m_{2}h^{2}}\right)}\end{split}$ (44)
+$\begin{split}&\quad T_{J}\left({\rho(m)}\right)-T_{J}\left({\rho^{[h]}(m)}\right)\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)}-\frac{J^{2}}{2I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}\left({I\left({\rho^{[h]}(m)}\right)-I\left({\rho(m)}\right)}\right)}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\\ &=\frac{J^{2}}{2I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}\cdot{\left({{(-2h)\iint_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\left({y_{1}-x_{1}}\right)\rho_{U}(x)\rho_{V}(y)\,dx\,dy}}+m_{1}m_{2}h^{2}}\right)}\end{split}$ (45)
 On the other hand,
@@
 Let $h>0$ small enough, such that if $\rho_{U}(x)>0$ and $\rho_{V}(y)>0$ then $y_{1}-x_{1}-h>0$, by Taylor expansion we have $\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h}\right|}=\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)$, therefore,
-$\begin{split}G\left({\rho_{m},\rho_{m}}\right)-G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)&=G\left({\rho_{U}+\rho_{V},\rho_{U}+\rho_{V}}\right)-G\left({\rho_{U}^{[h]}+\rho_{V}^{[h]},\rho_{U}^{[h]}+\rho_{V}^{[h]}}\right)\\ &=2G\left({\rho_{U},\rho_{V}}\right)-2G\left({\rho_{U}^{[h]},\rho_{V}^{[h]}}\right)\\ &=2{\int{\frac{\rho_{U}(x)\rho_{V}(y)-\rho_{U}\left({x-h_{1}}\right)\rho_{V}\left({y+h_{2}}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &=2{\int{\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y}\right|}-\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y+h}\right|}d^{3}xd^{3}y}}\\ &=2{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)}\right)d^{3}xd^{3}y}}\\ &=2{\int_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}d^{3}xd^{3}y}}+2m_{1}m_{2}O\left(h^{2}\right)\\ &\leq-\frac{2}{R^{3}(m)}{\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}+2m_{1}m_{2}O\left(h^{2}\right)\end{split}$ (46)
+$\begin{split}G\left({\rho_{m},\rho_{m}}\right)-G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)&=G\left({\rho_{U}+\rho_{V},\rho_{U}+\rho_{V}}\right)-G\left({\rho_{U}^{[h]}+\rho_{V}^{[h]},\rho_{U}^{[h]}+\rho_{V}^{[h]}}\right)\\ &=2G\left({\rho_{U},\rho_{V}}\right)-2G\left({\rho_{U}^{[h]},\rho_{V}^{[h]}}\right)\\ &=2{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{U}(x)\rho_{V}(y)-\rho_{U}\left({x-h_{1}}\right)\rho_{V}\left({y+h_{2}}\right)}{\left|{x-y}\right|}\,dx\,dy}}\\ &=2{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y}\right|}-\frac{\rho_{U}(x)\rho_{V}(y)}{\left|{x-y+h}\right|}\,dx\,dy}}\\ &=2{\iint_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({\frac{-\left({y_{1}-x_{1}}\right)h}{\left|{x-y}\right|^{3}}+O\left(h^{2}\right)}\right)\,dx\,dy}}\\ &=2h{\iint_{\{{\rho_{U}(x)>0,\rho_{V}(y)>0}\}}{\rho_{U}(x)\rho_{V}(y)\frac{\left(-\left({y_{1}-x_{1}}\right)\right)}{\left|{x-y}\right|^{3}}\,dx\,dy}}+2m_{1}m_{2}O\left(h^{2}\right)\\ &\leq-\frac{2h}{R^{3}(m)}{\iint_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)\,dx\,dy}}+2m_{1}m_{2}O\left(h^{2}\right)\end{split}$ (47)
 where $R(m)=B(m)\cdot R(J)$ given in Corollary 5.23. Similarly, we can show when $h$ is small enough,
-$\begin{split}&G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\\ &={\int{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)+\rho_{V}(y)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}-{\int{\frac{\rho_{1-m}(x)\left({\rho_{U}\left({y-h_{1}}\right)+\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &={\int{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)-\rho_{U}\left({y-h_{1}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}+{\int{\frac{\rho_{1-m}(x)\left({\rho_{V}(y)-\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}d^{3}xd^{3}y}}\\ &={\int{\rho_{1-m}(x)\rho_{U}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y-h_{1}}\right|}}\right)d^{3}xd^{3}y}}+{\int{\rho_{1-m}(x)\rho_{V}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h_{2}}\right|}}\right)d^{3}xd^{3}y}}\\ &\leq{\int{\rho_{1-m}(x)\rho_{U}(y)\left(\frac{h_{1}}{\left|{x-y}\right|\left|{x-y-h_{1}}\right|}\right)d^{3}xd^{3}y}}+{\int{\rho_{1-m}(x)\rho_{V}(y)\left(\frac{h_{2}}{\left|{x-y}\right|\left|{x-y+h_{2}}\right|}\right)d^{3}xd^{3}y}}\end{split}$
+$\begin{split}&G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\\ &={\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)+\rho_{V}(y)}\right)}{\left|{x-y}\right|}\,dx\,dy}}-{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{U}\left({y-h_{1}}\right)+\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}\,dx\,dy}}\\ &={\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{U}(y)-\rho_{U}\left({y-h_{1}}\right)}\right)}{\left|{x-y}\right|}\,dx\,dy}}+{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\frac{\rho_{1-m}(x)\left({\rho_{V}(y)-\rho_{V}\left({y+h_{2}}\right)}\right)}{\left|{x-y}\right|}\,dx\,dy}}\\ &={\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{U}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y-h_{1}}\right|}}\right)\,dx\,dy}}+{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{V}(y)\left({\frac{1}{\left|{x-y}\right|}-\frac{1}{\left|{x-y+h_{2}}\right|}}\right)\,dx\,dy}}\\ &\leq{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{U}(y)\left(\frac{h_{1}}{\left|{x-y}\right|\left|{x-y-h_{1}}\right|}\right)\,dx\,dy}}+{\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}{\rho_{1-m}(x)\rho_{V}(y)\left(\frac{h_{2}}{\left|{x-y}\right|\left|{x-y+h_{2}}\right|}\right)\,dx\,dy}}\end{split}$
 By the construction of $\Omega_{1-m}$ and $\Omega_{m}$ (24) we know $\rho_{1-m}(x)>0$, $\rho_{m}(y)>0$ implies $\frac{J^{2}}{2m^{2}\left({1-m}\right)^{2}}\leq\left|{x-y}\right|\leq\frac{3J^{2}}{2m^{2}\left({1-m}\right)^{2}}$, and when $h$ is small enough, $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{1}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$ and $\frac{J^{2}}{4m^{2}\left({1-m}\right)^{2}}<\left|{x-y-h_{2}}\right|<\frac{2J^{2}}{m^{2}\left({1-m}\right)^{2}}$.
-We also notice that by construction ${\int{\rho_{U}h_{1}}}={\int{\rho_{V}h_{2}}}=\frac{m_{1}m_{2}h}{m_{1}+m_{2}}=\frac{m_{1}m_{2}h}{m}$, then
+We also notice that by construction
+${\int_{\mathbb{R}^{3}}{\rho_{U}h_{1}}\,dx}={\int_{\mathbb{R}^{3}}{\rho_{V}h_{2}}\,dx}=\frac{m_{1}m_{2}h}{m_{1}+m_{2}}=\frac{m_{1}m_{2}h}{m}$
+Then
 $G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)\leq\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h$
@@
 Therefore, similar to Corollary 5.23 and due to Definition 5.1 which gives coefficient $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$, we collect (44) (45) (46) (47) (48) and have (recall we assume $\gamma>\frac{3}{2}$)
-$\begin{split}&E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}(m)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}}\right)\cdot{\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}(m)}-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$ (48)
-Therefore, when $d>\frac{32m^{3}\left({1-m}\right)^{5}R^{3}(m)}{J^{4}}=\frac{32\left({1-m}\right)^{5}R^{3}(J)}{J^{4}}m^{\frac{12\gamma-18}{3\gamma-4}}$, we know $E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$ for $h$ small enough.
-That is, if $\rho(m)$ is a constrained minimizer, then $d\leq Cm^{\frac{12\gamma-18}{3\gamma-4}}$ for some $C$ and all small $m$.
-More precisely, we have the following proposition:
+$\begin{split}&E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}(m)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{m}}\right)\cdot h\cdot{\iint_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)\,dx\,dy}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}(m)}-\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$ (49)
+Therefore, when $d>\frac{32m^{3}\left({1-m}\right)^{5}R^{3}(m)}{J^{4}}=\frac{32\left({1-m}\right)^{5}R^{3}(J)}{J^{4}}m^{\frac{12\gamma-18}{3\gamma-4}}$, we know for $h$ small enough,
+$E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)>0$
+That is, if $\rho(m)$ is a constrained minimizer, then $d\leq Cm^{\frac{12\gamma-18}{3\gamma-4}}$ for some $C$ and all small $m$. More precisely, we have the following proposition:
 Proposition 7.1.
 Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13.
-There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{m}$ has 2 convex connected components, then the distance between these 2 components is less then $\eta$, where $\eta=Cm^{\frac{12\gamma-18}{3\gamma-4}}$.
+There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{m}$ has 2 convex connected components, then the distance between these 2 components is less than $\eta$, where $\eta=Cm^{\frac{12\gamma-18}{3\gamma-4}}$.
 Proof.
@@
 Remark 7.2.
-We put the condition $\gamma>3/2$ so that we can apply Theorem 2.13, but in fact the estimate (48) actually holds even for smaller values of $\gamma$.
-Actually, similar result can be shown for stars’ densities.
+We put the condition $\gamma>\frac{3}{2}$ so that we can apply Theorem 2.13, but in fact the estimate (49) actually holds even for smaller values of $\gamma$.
+Actually, a similar result can be shown for stars’ densities.
 Corollary 7.3.
 Given polytropic law $P(\rho)=K\rho^{\gamma}$ indexed by $\gamma>\frac{3}{2}$ and a constrained minimizer $\rho(m)=\rho_{1-m}+\rho_{m}$ mentioned in Theorem 2.13.
-There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{1-m}$ has 2 convex connected components, then the distance between these 2 components is less then $\eta$, where $\eta=Cm^{5}$.
+There is a $C>0$ and a $\delta>0$, such that for all $m\in(0,\delta)$, if spt $\rho_{1-m}$ has 2 convex connected components, then the distance between these 2 components is less than $\eta$, where $\eta=Cm^{5}$.
 Proof.
 The proof is essentially the same as the proof above. Suppose spt $\rho_{1-m}$ has two convex connected components $U$ and $V$, after moving $U$ and $V$ similarly, we have
-$\begin{split}&\quad E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}\left({1-m}\right)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}}\right){\int_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)hd^{3}xd^{3}y}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}\left({1-m}\right)}-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$
-We just note that for star density, R(1m)R(1-m) will bounded by $R(J)$ as $m$ goes to 0 ($R(J)$ is given in Corollary 5.26).
+$\begin{split}&\quad E_{J}\left({\rho(m)}\right)-E_{J}\left({\rho^{[h]}(m)}\right)\\ &\geq\left({\frac{1}{R^{3}\left({1-m}\right)}-\frac{J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}}\right){\iint_{\{{\rho_{U}{(x)}>0,\rho_{V}{(y)}>0}\}}{\rho_{U}(x)\rho_{V}(y)\left({y_{1}-x_{1}}\right)h\,dx\,dy}}\\ &\quad+O\left(h^{2}\right)-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}h\\ &\geq\left({\frac{dm_{1}m_{2}}{2R^{3}\left({1-m}\right)}-\frac{16m_{1}m_{2}m^{5}\left({1-m}\right)^{3}}{J^{4}}}\right)h+O\left(h^{2}\right)\end{split}$
+We just note that for star density, R(1m)R(1-m) will be bounded by $R(J)$ as $m$ goes to 0 ($R(J)$ is given in Corollary 5.25).
 Thus we know $\frac{1}{R^{3}\left({1-m}\right)}\geq\frac{2J^{2}}{I\left({\rho(m)}\right)I\left({\rho^{[h]}(m)}\right)}\frac{1}{1-m}$ for $m$ small enough.
@@
 Furthermore, we can also use Theorem 2.13 (ii) to rule out some configurations of connected components.
-For example, the case where an outer shell encloses an inner solid sphere is impossible for $\rho_{m}$ or $\rho_{1-m}$.
+For example, the case in which an outer shell encloses an inner solid sphere is impossible for $\rho_{m}$ or $\rho_{1-m}$.
 Remark 7.5.
 It should be noted that the troublesome term in (49) preventing us from further reducing the distance between connected components $U$ and $V$ is the term $\frac{16m_{1}m_{2}m^{3}\left({1-m}\right)^{5}}{J^{4}}$, which comes from (48).
-In other words, we might be able to prove simple connectedness in each subdomain by moving and gluing the relevant connected components, if we can show that $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)$ can absorb $\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$.
+In other words, it may become possible to prove simple connectedness in each subdomain by moving and gluing the relevant connected components, if one can show that $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)$ can absorb $\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$.
 For instance, this could be achieved by proving $G\left({\rho_{m}^{[h]},\rho_{m}^{[h]}}\right)-G\left({\rho_{m},\rho_{m}}\right)>10\left|{G\left({\rho_{1-m},\rho_{m}-\rho_{m}^{[h]}}\right)}\right|$ for all small $m$ (that is $m\in(0,\delta)$ for some $\delta>0$) and for all positive distance $d$ between two connected components $U$ and $V$.
@@
 Conjecture 7.6.
-The support of $\rho(m)$ in Theorem 2.13 should have exactly 2 connected components when $m$ small enough.
+The support of $\rho(m)$ in Theorem 2.13 should have exactly 2 connected components when $m$ is small enough.
 Remark 7.7.
 Results on simple connectedness obtained in other settings may also be consulted; see, for example, [16, 29, 30].
-Appendix A Appendix
-Here we introduce some preliminary knowledge which are used in the paper.
-A.1 Calculus of Variations in Vector Spaces
-Since the functional we deal with is not differentiable in any usual sense, we begin with a less restrictive notion of differentiability.
-Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$.
-Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)>0,\text{ such that }u+tv\in U\text{ for }t\in[0,\epsilon(v))\}$.
-This set is a cone in $X$.
-If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$.
-We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $\mu$ on $X(P)$ such that $\forall v\in P$
-$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=\mu(v),$
-The linearity of $\mu$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $\mu$ by $E_{u}^{\prime}$—with $P$ understood from context.
-We say $E$ has a weak local minimum at $u\in U$ if for each $v\in P(u)$, there exists $\{t_{n}\}$ such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$ and $E\left({u+t_{n}v}\right)\geq E(u)$ for $n$ large enough.
-And $u$ is called weak local minimizer.
-We discuss more about the motivation of weak local minimizer in [15].
-Let $W\subset\left\{u\in U\middle|g(u)=M\right\}$, where $g$ is a linear functional on $X$ and $M$ is a constant.
-Then similarly, we say $E$ has a weak local minimum at $u$ restricted to $W$ if $u\in W$ and for each $v\in P(u)$, there exists $t_{n}$} such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$, and for for $n$ large enough, we have $u+t_{n}v\in W$, and $E\left({u+t_{n}w}\right)\geq E(u)$.
-Furthermore, we have a result for weak local minimum restricted to $W$:
-Theorem A.1 ((Generalized) Lagrange Multiplier Theorem [4, Section 2] [15, Appendix A]).
-Suppose $E$ is $P$-differentiable at $u\in U$ and that its restriction to $W$ has a weak local minimum at $u$.
-Suppose also that $P$ is convex and that there is a $u_{0}\in P$ such that $u_{0}\in P$ and $g\left(u_{0}\right)\neq 0$.
-Then there is a unique constant $\lambda$ such that $E_{u}^{\prime}(v)\geq\lambda g(v)$, all $v\in P$.
-A.2 Properties of Sobolev Spaces
-We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this paper.
-Proposition A.2 (Interpolation Inequality [8, Section 4.2]).
-If $f\in L^{p}\cap L^{q}$ with $1\leq p\leq q\leq\infty$, then $f\in L^{r}$ for all $r,p\leq r\leq q$, and the following interpolation inequality holds:
-$\|f\|_{r}\leq\|f\|_{p}^{\alpha}\|f\|_{q}^{1-\alpha},\text{ where }\frac{1}{r}=\frac{\alpha}{p}+\frac{1-\alpha}{q},0\leq\alpha\leq 1$
+Appendix
+Appendix A Properties of Sobolev Spaces
+Here we recall two propositions about the potential $V_{\rho}(x)=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|x-y|}\,dy=\left(\frac{1}{|\cdot|}*\rho\right)(x)$.
 Proposition A.1 (Hardy-Littlewood-Sobolev Inequality [5, Theorem 1.7]).
@@
 $\left\||\cdot|^{-\alpha}*f\right\|_{L^{r}\left(\mathbb{R}^{n}\right)}\leq C_{p,\alpha,n}\|f\|_{L^{p}\left(\mathbb{R}^{n}\right)}$ (50)
-It turns out for potential $V_{\rho}$ given in (9), we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int\frac{y_{j}\rho(x-y)}{|y|^{3}}d^{3}y$ not only in the distribution sense but also in the classical sense:
-Proposition A.4 (Differentiability of Potential [4, Proposition 7] [15, Proposition A.13]).
+Proposition A.2 (Differentiability of Potential [4, Section 3] [10, Appendix B]).
 If $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$ for some $p>3$, then $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)$ is continuously differentiable and the weak derivative coincides with the classical one for all $x\in\mathbb{R}^{3}$.
-To prove the low semicontinuity of the energy, we hope to connect the weak convergence of densities and strong convergence of potentials.
-Proposition A.5 (Compactness of Convolution (Potential) Operator with $q\geq 1$ [16, Proposition A.7]).
-Let $G(x)=\frac{1}{|x|}$, then $G\in L_{w}^{3}\left(\mathbb{R}^{3}\right)$.
-Let F=GBF=G*B, where $B$ is a bounded set in $L^{q}\left(\mathbb{R}^{3}\right)$ with $q\geq 1$.
-Then let $\Omega\subset\mathbb{R}^{3}$ is open, bounded and of class $C^{1}$.
-  1. (1)
-    If $1\leq q<\frac{3}{2}$, then $\forall 1\leq r<p$, where $\frac{1}{p}=\frac{1}{q}-\frac{2}{3},\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
-  2. (2)
-    If $q=\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\forall 1\leq r<\infty,\left.F\right|_{\Omega}$ has compact closure in $L^{r}(\Omega)$.
-  3. (3)
-    If $q>\frac{3}{2}$, we assume in addition $B$ is bounded in $L^{1}\left(\mathbb{R}^{3}\right)$, then $\left.F\right|_{\Omega}$ has compact closure in $C(\bar{\Omega})$.
 Acknowledgments
@@
 This work was primarily carried out during the author’s Master’s studies at the University of Bonn.
-The author thanks Juan Velázquez (master’s thesis advisor) and Dimitri Cobb (second advisor).
+The author thanks Juan Velázquez and Dimitri Cobb for their continued advice and support since the author’s time in Bonn, as well as Christof Sparber and Mimi Dai for their comments and support during the author’s Ph.D.
+studies at the University of Illinois Chicago.
 Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino, Han Cao for discussions, and to Théophile Dolmaire and other instructors.